### v1

# Transverse Instability of Stokes Waves at Finite Depth

Abstract.
A Stokes wave is a traveling free-surface periodic water wave that is constant in the direction transverse to the direction of propagation. In 1981 McLean discovered via numerical methods that Stokes waves are unstable with respect to transverse perturbations. In [17] for the case of infinite depth we proved rigorously that the spectrum of the water wave system linearized at small Stokes waves, with respect to transverse perturbations, contains unstable eigenvalues lying approximately on an ellipse. In this paper we consider the case of finite depth and prove that the same spectral instability result holds for all but finitely many values of the depth. The computations and some aspects of the theory are considerably more complicated in the finite depth case.
Key words and phrases:
Stokes waves, gravity waves, transverse instabilities, unstable eigenvalues, finite depth
MSC Classification: 76B07, 35Q35, 35R35, 35C07, 35B35.
1. Introduction
We consider classical water waves that are irrotational, incompressible, and inviscid. The water lies below an unknown free surface $S$. Such waves have been studied for over two centuries, notably by Stokes [36]. A Stokes wave is a two-dimensional steady wave traveling in a fixed horizontal direction at a fixed speed $c$. It has been known for a century that a curve of small-amplitude Stokes waves exists [31, 27, 37]. Several decades ago it was proven that the curve extends to large amplitudes as well [24].
A brief history of longitudinal (non-transverse) instabilities was given in [17], to which the reader is referred. That literature left open the question of whether a small Stokes wave could be unstable when perturbed in both horizontal directions but keeping the longitudinal period unperturbed. This transverse instability problem was studied numerically first by Bryant [12] and then by much more detailed work of McLean et al [30, 28, 29]. While these remarkable papers did detect transverse instabilities, a mathematical proof of these three-dimensional instabilities has been missing ever since. In [17] we considered the case of infinite depth and proved rigorously that the spectrum of the water wave system linearized at small Stokes waves contains unstable eigenvalues lying approximately on an ellipse. The purpose of the present paper is to prove such a spectral instability result for the case of finite depth.
For background and references on longitudinal (modulational and high frequency) instabilities, we refer to [17]. To bring the references up to date, we mention some very recent work on longitudinal instability by Berti et al [8] and on near-extreme Stokes waves by Deconinck et al [20]. It is important to note that there are several other models of water waves for which the transverse instability has been studied rigorously, for instance for gravity-capillary waves.
As in [17] we now specify the parameters of our problem. Let $x$ and $y$ denote the horizontal variables and $z$ the vertical one. The problem depends in a significant way on the depth $h$. Consider the curve of Stokes waves traveling in the $x$-direction and with a given period, say $2\pi$ without loss of generality. This curve is parametrized by a small parameter $\varepsilon$ which represents the wave amplitude of the Stokes waves. Such a steady wave can be described in the moving $(x,z)$ plane (where $x-ct$ is replaced by $x$) by its free surface $S=\{(x,y,z)\ |\ z=\eta^{*}(x;\varepsilon)\}$ and by its velocity potential $\psi^{*}(x;\varepsilon)$ restricted to $S$.
We take the perturbation of $\eta^{*}$ to have the form $\overline{\eta}(x)e^{\lambda t+i\alpha y}$, where $\overline{\eta}$ has the same period $2\pi$ as the Stokes wave, $\lambda\in\mathbb{C}$ is the growth rate of the perturbation, and $\alpha\in\mathbb{R}$ is the transverse wave number of the perturbation. The problem is to find at least one value of $\alpha$ that leads to instability, that is, $\text{Re}\lambda>0$. After linearizing the nonlinear water wave system about a Stokes wave, introducing a “good-unknown,” and performing a conformal mapping change of variables, we find that the exponents $\lambda$ are eigenvalues of a linear Hamiltonian operator $\mathcal{L}_{h,\varepsilon,\beta}$, where $\beta=\alpha^{2}$. Of course, the operator depends in a significant way on the depth $h$.
For $\varepsilon=0$ and depth $h$, the eigenvalues of the linearized operator are
$\lambda^{0}_{\pm}(k,\beta)=i\left[\tanh^{1/2}(h)k\pm(k^{2}+\beta)^{1/4}\tanh^{1/2}\left(h\left(k^{2}+\beta\right)^{1/2}\right)\right],\quad k\in\mathbb{Z}.$ (1.1)
It is clear that the limit as $h\to\infty$ is simpler. Motivated by [30], we determine a resonant transverse wave number $\alpha_{*}$ so that the unperturbed operator $\mathcal{L}_{h,0,\beta_{*}}$ with $\beta_{*}=\alpha_{*}^{2}$ has an imaginary double eigenvalue $i\sigma$. This eigenvalue corresponds to the lowest possible resonance that generates a Type II transverse instability according to McLean [30], of which there are infinitely many higher-order resonances that potentially generate higher-order transverse instabilities. We expect however that higher-order transverse instabilities have smaller growth rates for small Stokes waves.
In order to capture the transverse instabilities we introduce a new small parameter $\delta$ for the perturbation of $\beta$ about $\beta_{*}$. Our main result is that the perturbed operator $\mathcal{L}_{h,\varepsilon,\beta_{*}+\delta}$ has eigenvalues $\lambda_{\pm}$ with non-zero real parts that bifurcate from $i\sigma$, stated more precisely in the following theorem.
Theorem 1.1.
Let the depth $h\in(0,\infty)$ be given. Except for a finite number $N$ of values of $h$, the following instability statement is true. There exist $\varepsilon_{\textrm{max}}>0$ and $\delta_{\text{max}}>0$ such that for all $\varepsilon\in(-\varepsilon_{\textrm{max}},\varepsilon_{\textrm{max}})$ and $\delta\in(-\delta_{\text{max}},\delta_{\text{max}})$, the operator $\mathcal{L}_{h,\varepsilon,\beta_{*}+\delta}$ has a pair of eigenvalues
$\lambda_{\pm}$ $=i\Big{(}\sigma+\frac{1}{2}T(\varepsilon,\delta)\Big{)}\pm\frac{1}{2}\sqrt{\Delta(\varepsilon,\delta)},$ (1.2)
where $T$ and $\Delta$ are real-valued, real-analytic functions such that $T(\varepsilon,\delta)=O\left(\delta\right)$ and $\Delta(\varepsilon,\delta)=O\left(\delta^{2}\right)$ as $(\varepsilon,\delta)\rightarrow(0,0)$. Furthermore, there exist $\kappa_{0}\in\mathbb{R}$ and $\kappa_{1}>0$ such that for
$\delta=\delta(\varepsilon,\theta)=\kappa_{0}\varepsilon^{2}+\theta\varepsilon^{3}\quad\textrm{with}\quad|\theta|<\kappa_{1},$ (1.3)
we have $\Delta(\varepsilon,\delta(\varepsilon,\theta))>0$ for sufficiently small $\varepsilon$. Thus the eigenvalue $\lambda_{+}$ has positive real part provided $\delta=\delta(\varepsilon,\theta)$ with $|\theta|<\kappa_{1}$ and $\varepsilon$ is sufficiently small. Moreover, $\text{Re}\lambda_{+}=O\left(\varepsilon^{3}\right)$ as $\varepsilon\rightarrow 0$ for each $\theta$. This means that there exist transverse perturbations of the given Stokes wave whose amplitudes grow temporally like $e^{t\,\text{Re}\lambda_{+}}$.
Remark 1.2.
In the corresponding longitudinal result in [8], there could be infinitely many exceptional values of the depth, whereas in the transverse case Theorem 1.1 asserts that only finitely many such values can exist. In fact, numerical computation shows that $N=1$, that is, there is only one such exceptional value of the depth. See Figure 3 in Section 5. It should also be noted that small Stokes waves at the exceptional depth(s) could still be susceptible to higher-order transverse instabilities.
Fixing $\varepsilon$, substituting $\delta=\delta(\varepsilon,\theta)$ into (1.2) and dropping terms of $O\left(\varepsilon^{4}\right)$ and smaller, we obtain an asymptotic expansion of the unstable eigenvalues. As $\theta$ varies, we find that the eigenvalues lie approximately on an ellipse (an ‘isola’) that is centered on the imaginary axis. Figure 1 illustrates this ellipse for three different depths. As $\varepsilon$ varies, the center of the ellipse drifts from a double eigenvalue on the imaginary axis like $O\left(\varepsilon^{2}\right)$, while its semi-major and semi-minor axes scale like $O\left(\varepsilon^{3}\right)$. We refer to Corollary 5.4 for the precise statement.
Remark 1.3.
Figure 1 compares our transverse instability isolae with the numerically computed eigenvalues for a Stokes wave with amplitude $\varepsilon=0.01$ at the three depths $h=1$, $3/2$, and $2$. The eigenvalues are computed using the Floquet-Fourier-Hill method applied to the Ablowitz-Fokas-Musslimani formulation of the finite-depth water wave problem. Full details of this method are presented in [19]. The distance between our transverse instability isolae and the numerically computed eigenvalues is $O\left(\varepsilon^{4}\right)$, as expected, since we do not calculate beyond $O\left(\varepsilon^{3}\right)$ to construct our isolae. The strong agreement between our isolae and the numerically computed eigenvalues to $O\left(\varepsilon^{3}\right)$ provides confidence both in our analytical results and in earlier numerical investigations of the transverse instability spectrum. Even better agreement could be found by retaining higher-order corrections of the unstable eigenvalues in a manner similar to [15].
Figure 1. A comparison of the transverse instability isola obtained in this work (orange curves) and numerical computations of the unstable eigenvalues of the transverse instability (blue dots) for a Stokes wave with amplitude $\varepsilon=0.01$ in water of depth $h=1$ (left), $h=3/2$ (middle), and $h=2$ (right). The center of the isola is subtracted from its imaginary component to show a sense of scale. In each plot the distance between the orange curves and blue dots is $O\left(\varepsilon^{4}\right)$.
We now turn to the main ideas in the proof of Theorem 1.1. It begins by finding an expression for $\mathcal{L}_{h,\varepsilon,\beta}$ by a method analogous to that in [32] and [17]. It continues by following the method of [4] that uses a Kato similarity transformation to reduce the relevant spectral data of $\mathcal{L}_{h,\varepsilon,\beta}$ to a $2\times 2$ matrix $\textrm{L}_{h,\varepsilon,\delta}$ with the property that $i\textrm{L}_{h,\varepsilon,\delta}$ is real and skew-adjoint. We prove that the entries of this matrix are real analytic functions of $\varepsilon$ and $\delta$ and we obtain convenient functional expressions for its eigenvalues, resulting in (1.2). Compared to the infinite depth case [17], these expressions are much more complicated. In order to conclude that $\Delta(\varepsilon,\beta)>0$ for $\delta=\delta(\varepsilon,\theta)$ and sufficiently small $\varepsilon$, we must expand the entries of the matrix in a power series up to third order in the pair $(\varepsilon,\delta)$. Second order is insufficient. The extremely arduous calculations require us to take advantage of Mathematica. They can be found in the companion Mathematica file CompanionToTransverseInstabilities.nb.
In Section 2 we introduce the Stokes waves and proceed with the linearization and the flattening by means of a conformal mapping from a strip $\Omega_{\varepsilon}=\mathbb{R}\times(-h_{\varepsilon},0)$ to the two-dimensional fluid domain of the Stokes wave. The main result that is required here is Theorem 2.6, which is devoted to the three-dimensional Dirichlet-Neumann operator $\mathcal{G}_{\varepsilon,\beta}$ that occurs in the linearized operator subject to the two-dimensional conformal mapping, and the proof of its analyticity in $\varepsilon$ and $\delta$. The lack of a trace theorem from $H^{1/2}$ of a domain to its boundary necessitates a careful alternative analysis. Compared to the infinite depth case [17], the proof of analyticity is considerably more difficult due to the dependence on $\varepsilon$ of the strip $\Omega_{\varepsilon}$.
Theorem 1.1 is then reduced to studying the eigenvalues of the linearized operator $\mathcal{L}_{h,\varepsilon,\beta}$, which has a Hamiltonian form and is reversible. Section 3 is devoted to a discussion of the first resonance on the imaginary axis when $\varepsilon=0$, the introduction of Kato’s perturbed basis, and the reduction to the study of the eigenvalues of a reduced $2\times 2$ matrix $\textrm{L}_{h,\varepsilon,\delta}$. In Section 4 we perform extremely tedious expansions of $\mathcal{L}_{h,\varepsilon,\beta_{*}+\delta}$ out to third order in both $\varepsilon$ and $\delta$, which are considerably more tedious than the corresponding expansions in [17, 32, 4]. We use the expansions of $\mathcal{L}_{h,\varepsilon,\beta_{*}+\delta}$ to compute the expansions of the Kato basis vectors and of the matrix $\textrm{L}_{h,\varepsilon,\delta}$. Finally in Section 5 we analyze the leading terms in the characteristic discriminant of $\textrm{L}_{h,\varepsilon,\delta}$. An important step is to prove that a key coefficient, which we call $b_{3,0}$, does not vanish. It is here that we must exclude the exceptional depths. By determining the precise asymptotics of $b_{3,0}$ as $h\to 0^{+}$ and $h\to\infty$, we are able to prove that there are at least one but at most only finitely many such exceptional depths. This concludes the proof of Theorem 1.1. The full equation that defines the isola of eigenvalues up to $O(\varepsilon^{4})$ is given in Corollary 5.4. Finally, we prove a trace result in Appendix A.1, which is used in the proof of the analyticity of the operator $\mathcal{G}_{\varepsilon,\beta}$.
2. Transverse perturbations of Stokes waves
2.1. Stokes waves
We consider the three-dimensional fluid domain
$D(t)=\{(x,y,z)\in\mathbb{R}^{3}\ :\ -h<z<\eta(x,y,t)\}$ (2.1)
with free surface $S(t)=\{(x,y,\eta(x,y,t)):(x,y)\in\mathbb{R}^{2}\}$ over a constant depth $h>0$. Assuming that the fluid is incompressible, inviscid and irrotational, the velocity field $\bf u$ admits a harmonic potential $\phi(x,y,z,t):D(t)\to\mathbb{R}$. That is, $\bf u=\nabla\phi$. Then $\phi$ and $\eta$ satisfy the water wave system
$\begin{cases}\Delta_{x,y,z}\phi=0\quad\text{ in }D(t),\\ \partial_{t}\phi+\tfrac{1}{2}|\nabla_{x,y,z}\phi|^{2}=-g\eta+P\quad\text{ on }S(t),\\ \partial_{t}\eta+\partial_{x}\phi\partial_{x}\eta=\partial_{y}\phi\quad\text{ on }S(t),\\ \partial_{z}\phi(x,y,-h)=0,\end{cases}$ (2.2)
where $P$ is the Bernoulli constant and $g>0$ is the constant acceleration due to gravity. The second equation is Bernoulli’s, which follows from the pressure being constant along the free surface; the third equation expresses the kinematic boundary condition that particles on the surface remain there; the last condition asserts that there is no flux of water out of the bottom boundary $\{z=-h\}$. For convenience we will take $P=0$.
In order to reduce the system to the free surface $S$, we introduce the Dirichlet-Neumann operator $G(\eta)$ associated to $D$, namely,
$(G(\eta)f)(x,y):=\partial_{z}\vartheta(x,y,\eta(x,y))-\nabla_{x,y}\vartheta(x,y,\eta(x,y))\cdot\nabla_{x,y}\eta(x,y),$ (2.3)
where $\vartheta(x,y,z)$ solves the elliptic problem
$\begin{cases}\Delta_{x,y,z}\vartheta=0\quad\text{in}~{}D,\\ \vartheta|_{S}=f(x,y),\\ \partial_{z}\vartheta(x,y,-h)=0.\end{cases}$ (2.4)
We define $\psi$, not as the stream function, but as the trace of the velocity potential on the free surface, $\psi(x,y,t)=\phi(x,y,\eta(x,y,t),t)$. Then, in the moving frame with speed $(c,0)\in\mathbb{R}^{2}$, the gravity water wave system written in the Zakharov-Craig-Sulem formulation [38, 14] is
$\begin{cases}\partial_{t}\eta=c\partial_{x}\eta+G(\eta)\psi=:F_{1}(\eta,\psi,c),\\ \partial_{t}\psi=c\partial_{x}\psi-\frac{1}{2}|\nabla_{x,y}\psi|^{2}+\frac{1}{2}\frac{\big{(}G(\eta)\psi+\nabla_{x,y}\psi\cdot\nabla_{x,y}\eta\big{)}^{2}}{1+|\nabla_{x,y}\eta|^{2}}-g\eta=:F_{2}(\eta,\psi,c),\end{cases}$ (2.5)
By the change of variables
$c\to\sqrt{g}{c},\quad\eta(x,y,t)\to\eta(x,y,\sqrt{g}t),\quad\psi(x,y,t)\to\sqrt{g}\psi(,x,y,\sqrt{g}t),$
we henceforth assume without loss of generality that $g=1$.
By a Stokes wave we mean a periodic steady solution of (2.5) that is independent of $y$. Without loss of generality, we henceforth consider $2\pi$-periodic Stokes waves. As has been known for over a century, there exists a curve of small Stokes waves parametrized analytically by the small amplitude $\varepsilon$.
Theorem 2.1.
For any physical depth $h>0$ and any index $s>5/2$, there exist $\varepsilon_{St}(s)>0$ depending on $h$ and a unique family of Stokes waves
$\big{(}\eta^{*}(x;\varepsilon),\psi^{*}(x;\varepsilon),c^{*}(\varepsilon)\big{)}\in H^{s}(\mathbb{T})\times H^{s}(\mathbb{T})\times\mathbb{R}$
parametrized by $|\varepsilon|<\varepsilon_{St}(s)$, such that
  • (i) the mapping $(-\varepsilon_{St}(s),\varepsilon_{St}(s))\ni\varepsilon\to\big{(}\eta^{*}(\cdot;\varepsilon),\psi^{*}(\cdot;\varepsilon),c^{*}(\varepsilon)\big{)}\in H^{s}(\mathbb{T})\times H^{s}(\mathbb{T})\times\mathbb{R}$ is analytic;
  • (ii) $\eta^{*}(\cdot;\varepsilon)$ is even and has average zero, and $\psi^{*}(\cdot;\varepsilon)$ is odd.
Proof.
See Theorem 1.3 in [5] for instance. ∎
We will need the following third-order expansions in $\varepsilon$ of the above Stokes waves.
Proposition 2.2.
The unique analytic family of Stokes waves given by Theorem 2.1 has the expansions
$c^{*}(\varepsilon)$ $=c_{0}+c_{2}\varepsilon^{2}+O(\varepsilon^{4}),$ (2.6)
$\eta^{*}(x;\varepsilon)$ $=\varepsilon\cos x+\varepsilon^{2}\Big{[}\eta_{2,0}+\eta_{2,2}\cos(2x)\Big{]}+\varepsilon^{3}\Big{[}\eta_{3,1}\cos x+\eta_{3,3}\cos(3x)\Big{]}+O(\varepsilon^{4}),$ (2.7)
$\psi^{*}(x;\varepsilon)$ $=\varepsilon c_{0}^{-1}\sin x+\varepsilon^{2}\psi_{2,2}\sin(2x)+\varepsilon^{3}\Big{[}\psi_{3,1}\sin x+\psi_{3,3}\sin(3x)\Big{]}+O(\varepsilon^{4}),$ (2.8)
where
$c_{0}=\sqrt{\tanh(h)},\qquad$ $c_{2}=\frac{-12c_{0}^{12}+13c_{0}^{8}-12c_{0}^{4}+9}{16c_{0}^{7}},$ (2.9)
$\eta_{2,0}=\frac{c_{0}^{4}-1}{4c_{0}^{2}},\quad\eta_{2,2}=\frac{-c_{0}^{4}+3}{4c_{0}^{6}},\quad$ $\eta_{3,1}=\frac{-2c_{0}^{12}+3c_{0}^{8}+3}{16c_{0}^{8}(1+c_{0}^{2})},\quad\eta_{3,3}=\frac{-3c_{0}^{12}+9c_{0}^{8}-9c_{0}^{4}+27}{64c_{0}^{12}},$ (2.10)
$\psi_{2,2}=\frac{c_{0}^{8}+3}{8c_{0}^{7}},\quad\psi_{3,1}=$ $\frac{2c_{0}^{12}-8c_{0}^{8}-3}{16c_{0}^{7}(1+c_{0}^{2})},\quad\psi_{3,3}=\frac{-9c_{0}^{12}+19c_{0}^{8}+5c_{0}^{4}+9}{64c_{0}^{13}}.$ (2.11)
Proof.
See (1.9) in [7]. ∎
2.2. Transverse perturbations, linearization, and transformations
Let us fix a Stokes wave $(\eta^{*},\psi^{*},c^{*})$ with amplitude $\varepsilon$ and perturb it by a two-dimensional perturbation:
$\eta(x,y)=\eta^{*}(x)+\nu\overline{\eta}(x,y),\quad\psi(x,y)=\psi^{*}(x)+\nu\overline{\psi}(x,y),\quad|\nu|\ll 1.$ (2.12)
To prepare for the linearization of the dynamics of the full water wave equations, we recall the shape-derivative formula for the derivative of $G(\eta)\psi$ with respect to $\eta$.
Theorem 2.3.
Let $\eta:\mathbb{T}^{2}\to\mathbb{R}$ and $\psi:\mathbb{T}^{2}\to\mathbb{R}$. The shape-derivative of $G(\eta)\psi$ with respect to $\eta$ is denoted by
$G^{\prime}(\eta)\overline{\eta}\psi=\lim_{h\to 0}\frac{1}{h}\left(G(\eta+h\overline{\eta})\psi-G(\eta)\psi\right).$ (2.13)
We have
$G^{\prime}(\eta)\overline{\eta}\psi=-G(\eta)(\overline{\eta}B(\eta)\psi)-\operatorname{div}(\overline{\eta}V(\eta)\psi),$ (2.14)
where
$B(\eta)\psi:=\frac{G(\eta)\psi+\nabla\eta\cdot\nabla\psi}{1+|\nabla\eta|^{2}},\quad V(\eta)\psi:=\nabla\psi-B\nabla\psi.$ (2.15)
Proof.
See Theorem 3.20 in [25] and Proposition 2.11 in [1]. ∎
Using the formulas (2.14), (2.15), and the fact that the Stokes wave is independent of the transverse variable $y$, we linearize (2.5) around it to obtain
$\partial_{t}\overline{\eta}=\frac{\delta F_{1}(\eta^{*},\psi^{*},c^{*})}{\delta(\eta,\psi)}(\overline{\eta},\overline{\psi})=\partial_{x}\big{(}(c^{*}-V^{*})\overline{\eta}\big{)}+G(\eta^{*})(\overline{\psi}-{B^{*}}\overline{\eta}),$ (2.16)
$\partial_{t}\overline{\psi}=\frac{\delta F_{2}(\eta^{*},\psi^{*},c^{*},P^{*})}{\delta(\eta,\psi)}(\overline{\eta},\overline{\psi})=(c^{*}-{V^{*}})\partial_{x}\overline{\psi}+{B^{*}}G(\eta^{*})(\overline{\psi}-{B^{*}}\overline{\eta})-{B^{*}}\partial_{x}{V^{*}}\overline{\eta}-g\overline{\eta},$ (2.17)
where
$B^{*}:=B(\eta^{*})\psi^{*}=\frac{G(\eta^{*})\psi^{*}+\partial_{x}\psi^{*}\partial_{x}\eta^{*}}{1+|\partial_{x}\eta^{*}|^{2}},\qquad V^{*}:=V(\eta^{*})\psi^{*}=\partial_{x}\psi^{*}-B^{*}\partial_{x}\eta^{*}.$ (2.18)
A similar derivation was done in Lemma 3.2, [32] for perturbations $\overline{\eta}$ and $\overline{\psi}$ depending only on $x$. It is worthwhile noting that in fact $B^{*}$ and $V^{*}$ are the vertical and horizontal components, respectively, of the fluid velocity of the Stokes wave at the free surface $\{z=\eta^{*}(x)\}$. Where $G(\eta^{*})$ acts on functions of $(x,y)$ we consider $\eta^{*}(x,y)\equiv\eta^{*}(x)$. We change variables to the so-called Alinhac’s good unknowns
$v_{1}(x,y)=\overline{\eta}\quad\text{and}~{}v_{2}(x,y)=\overline{\psi}-{B^{*}}\overline{\eta}$ (2.19)
satisfying
$\partial_{t}v_{1}=\partial_{x}\big{(}(c^{*}-{V^{*}})v_{1}\big{)}+G(\eta^{*})v_{2},$ (2.20)
$\partial_{t}v_{2}=-\big{(}g+({V^{*}}-c^{*})\partial_{x}{B^{*}}\big{)}v_{1}+(c^{*}-{V^{*}})\partial_{x}v_{2}.$ (2.21)
Choosing a simple form for the transverse perturbations, we consider perturbations $\overline{\eta}$ and $\overline{\psi}$ that have wave number $\alpha\in\mathbb{R}\setminus\{0\}$ in the transverse variable $y$:
$\overline{\eta}(x,y)=\widetilde{\eta}(x)e^{i\alpha y},\quad\overline{\psi}(x,y)=\widetilde{\psi}(x)e^{i\alpha y}.$ (2.22)
Here $\widetilde{\eta}$ and $\widetilde{\psi}$ are $2\pi$-periodic functions, so the perturbations have the same longitudinal wave length as that of the Stokes wave. Consequently, the good unknowns have the form
$v_{1}(x,y)=\widetilde{\eta}(x)e^{i\alpha y}:=\widetilde{v_{1}}(x)e^{i\alpha y},\quad v_{2}(x,y)=(\widetilde{\psi}-{B^{*}}\widetilde{\eta})e^{i\alpha y}:=\widetilde{v_{2}}(x)e^{i\alpha y}.$ (2.23)
Then, in the linearized system (2.20)-(2.21), the most difficult term to analyze is $G(\eta^{*})(\widetilde{v_{2}}(x)e^{i\alpha y})$. To handle this term, we shall flatten the surface $\{z=\eta^{*}(x)\}$ using a Riemann mapping. For the sake of clarity, from now on we denote the independent variables in the physical space by $(X,y,Z)$. The Riemann mapping is as follows.
Theorem 2.4.
There exists a holomorphic bijection $X(x,z)+iZ(x,z)$ from a strip $\Omega_{\varepsilon}=\{(x,z)\in\mathbb{R}^{2}:-h_{\varepsilon}<z<0\}$ onto $\Omega=\{(X,Z)\in\mathbb{R}^{2}:-h<Z<\eta^{*}(X)\}$ with the following properties.
  • (i)
    $X(x+2\pi,z)=2\pi+X(x,z)$ and $Z(x+2\pi,z)=Z(x,z)$ for all $(x,z)\in\Omega_{\varepsilon}$;
    $X$ is odd in $x$ and $Z$ is even in $x$;
  • (ii)
    $X+iZ$ maps $\{(x,0):x\in\mathbb{R}\}$ onto the surface $S=\{(x,\eta^{*}(x)):x\in\mathbb{R}\}$ and maps $\{(x,-h_{\varepsilon}):x\in\mathbb{R}\}$ onto the bottom $\{(x,-h):x\in\mathbb{R}\}$.
  • (iii)
    Defining the “Riemann stretch” as
    $\zeta(x)=X(x,0),$ (2.24)
    we have the Fourier expansions
    $X(x,z)=x-\frac{i}{2\pi}\sum_{k\in\mathbb{Z}\setminus\{0\}}e^{ikx}\mathrm{sign}(k)\frac{\cosh(|k|(z+h_{\varepsilon})}{\sinh(|k|h_{\varepsilon})}\widehat{\eta^{*}\circ\zeta}(k)$ (2.25)
    $Z(x,z)=z+h_{\varepsilon}-h+\frac{1}{2\pi}\sum_{k\neq 0}e^{ikx}\frac{\sinh(|k|(z+h_{\varepsilon})}{\sinh(|k|h_{\varepsilon})}\widehat{\eta^{*}\circ\zeta}(k).$ (2.26)
  • (iv)
    The “conformal depth” $h_{\varepsilon}$ is given by
[…]
Here we have defined the Fourier transform as
[…]
Proof.
1. Assume first that there exist $h_{\varepsilon}>0$ and a conformal mapping $(X,Z):\Omega\to\Omega_{\varepsilon}$ such that (i) and (ii) hold. We will prove (iii) and (iv). For brevity, denote $f(x)=X(x,0)-x$. We have $\Delta X=0$ and $\partial_{z}X(x,-h_{\varepsilon})=0$ as well as $\Delta Z=0$ and $Z(x,-h_{\varepsilon})=-h$. By solving the Laplace problem for $X$ and using the Cauchy-Riemann equations, we find
$X(x,z)-x=\frac{1}{2\pi}\sum_{k}\hat{f}(k)\frac{\cosh(|k|(z+h_{\varepsilon})}{\cosh(|k|h_{\varepsilon})}e^{ikx}$ (2.28)
and
$Z(x,z)=z+h_{\varepsilon}-h+\frac{i}{2\pi}\sum_{k\neq 0}\hat{f}(k)\ \mathrm{sign}(k)\frac{\sinh(|k|(z+h_{\varepsilon})}{\cosh(|k|h_{\varepsilon})}e^{ikx}.$ (2.29)
Furthermore, on top we have
$\eta^{*}(\zeta(x))=Z(x,0)=h_{\varepsilon}-h+\frac{i}{2\pi}\sum_{k\neq 0}\hat{f}(k)\ \mathrm{sign}(k)\tanh(|k|h_{\varepsilon})e^{ikx},$ (2.30)
where $\zeta(x)=X(x,0)$. Thus $h_{\varepsilon}-h$ equals the average of $\eta^{*}\circ\zeta$, which is the same as (2.27). For $k\neq 0$, (2.30) gives us
$\hat{f}(k)=-i\frac{\mathrm{sign}(k)}{\tanh(|k|h_{\varepsilon})}\ \widehat{\eta^{*}\circ\zeta}(k).$ (2.31)
Substituting this formula for $\hat{f}$ into (2.28) and (2.29), and noting that $X(x,z)-x$ is odd in $x$, we obtain (2.25) and (2.26).
Thus we have derived the formulas (2.25), (2.26), and (2.27). Moreover, by evaluating (2.25) at $z=0$ and invoking (2.27), we obtain
$g(x):=\zeta(x)-x=\frac{i}{2\pi}\sum_{k\neq 0}e^{ikx}\frac{\mathrm{sign}(k)}{\tanh(|k|h_{\varepsilon})}\widehat{\eta^{*}\circ(I+g}(k),$ (2.32)
where $g$ is periodic and $Ix=x$.
2. Now we prove the existence of $h_{\varepsilon}$ and $(X,Z)$. Starting with arbitrary $g\in H^{s}(\mathbb{T})$ and $s\geq 1$, we define $w=\eta^{*}\circ(I+g):=\mathscr{G}_{0}(g)$, which occurs at the right end of (2.32). We notice in view of (2.32) that $g=\zeta-\operatorname{I}$ can be defined as the fixed point of the nonlinear mapping $g\to\mathscr{F}_{0}\mathscr{G}_{0}$, where
$\mathscr{F}_{0}(w):=-\frac{i}{2\pi}\sum_{k\neq 0}e^{ikx}\frac{\mathrm{sign}(k)}{\tanh(|k|(h+(2\pi)^{-1}\hat{w}(0)))}\hat{w}(k).$
There exists $r=r(h,s)$ sufficiently small so that if $\|w\|_{H^{s}(\mathbb{T})}\leq r_{0}$ then $(2\pi)^{-1}|\hat{w}(0)|\leq\frac{h}{2}$ and we have
$\|\mathscr{F}_{0}(w)\|_{H^{s}(\mathbb{T})}\leq C_{1}\|w\|_{H^{s}(\mathbb{T})},\quad C_{1}=C_{1}(s,h).$
On the other hand, there exists $C_{2}>0$ independent of $\varepsilon\in(0,1)$ such that for any $g\in H^{s}(\mathbb{T})$, we have
$\|\eta^{*}(\operatorname{I}+g)\|_{H^{s}(\mathbb{T})}\leq C_{2}\varepsilon(1+\|g\|_{H^{s}(\mathbb{T})}).$
Thus $\mathscr{F}_{0}\mathscr{G}_{0}$ maps the ball $B(0,r_{0})\subset H^{s}(\mathbb{T})$ to itself provided $\varepsilon<\frac{r}{C_{1}C_{2}(1+r)}$.
For the contraction of $\mathscr{F}_{0}\mathscr{G}_{0}$ we consider $w_{1}$, $w_{2}\in B(0,r)$. Using the mean value theorem for $\tanh(|k|(h+\cdot)$, we find
$\left|\frac{\hat{w_{1}}(k)}{\tanh(|k|(h+(2\pi)^{-1}\hat{w_{1}}(0)))}-\frac{\hat{w_{2}}(k)}{\tanh(|k|(h+(2\pi)^{-1}\hat{w_{2}}(0)))}\right|\leq C|\hat{w_{1}}(k)-\widehat{w_{2}}(k)|$
$\qquad+C|\hat{w_{1}}(k)\frac{|k|}{\cosh^{2}(|k|(h+\theta))}|\hat{w_{1}}(0)-\hat{w_{2}}(0)|$
for some $\theta$ between $(2\pi)^{-1}\hat{w_{1}}(0)$ and $(2\pi)^{-1}\hat{w_{2}}(0)$. In particular, $|\theta|\leq\frac{h}{2}$, and hence
$\frac{|k|}{\cosh^{2}(|k|(h+\theta))}\leq C$
since $\cosh(x)$ decays exponentially as $x\to\infty$. We deduce that
$\|\mathscr{F}_{0}(w_{1})-\mathscr{F}_{0}(w_{2})\|_{H^{s}(\mathbb{T})}\leq C_{3}\|w_{1}-w_{2}\|_{H^{s}(\mathbb{T})}(1+\|w_{1}\|_{H^{s}(T)}),\quad C_{3}=C_{3}(s,h).$
This implies
$\|\mathscr{F}_{0}\mathscr{G}_{0}(g_{1}))-\mathscr{F}_{0}\mathscr{G}_{0}(g_{2}))\|_{H^{s}(\mathbb{T})}$ $\leq C_{3}\|\eta^{*}(\operatorname{I}+g_{1})-\eta^{*}(\operatorname{I}+g_{2})\|_{H^{s}(\mathbb{T})}(1+\|\eta^{*}(\operatorname{I}+g_{1})\|_{H^{s}(\mathbb{T})})$
$\leq C_{4}\varepsilon\|g_{1}-g_{2}\|_{H^{s}(\mathbb{T})}(1+r).$
Therefore, $\mathscr{F}_{0}\mathscr{G}_{0}$ is a contraction on $B(0,r)\subset H^{s}(\mathbb{T})$ provided
$\varepsilon<\min\left\{\frac{r}{C_{1}C_{2}(1+r)},\frac{1}{2C_{4}(1+r)}\right\}.$
We conclude that $g=\zeta-\operatorname{I}$ is the unique fixed point of $\mathscr{F}_{0}\mathscr{G}_{0}$ in $B(0,r)$. Using $\zeta$ we define $h_{\varepsilon}$, $X$, and $Z$ by the formulas (2.27), (2.25), and (2.26). Then it can be directly checked that $X$ and $Z$ satisfy the desired properties. ∎
Remark 2.5.
$X$, $Z$, and $\zeta$ are analytic in the parameter $\varepsilon$ with values in Sobolev spaces. Indeed, we see that the right side of (2.32) is analytic in $\varepsilon$ with values in Sobolev spaces since $\eta^{*}$ is so. The analyticity of $\zeta$ in $\varepsilon$ then follows from the Analytic Implicit Function Theorem. Next, we return to the formula (2.25) in which $\eta^{*}\circ\zeta$ is now analytic in $\varepsilon$ with values in any Sobolev space provided $\varepsilon$ is small enough. Consequently $\widehat{\eta^{*}\circ\zeta}(k)$ is analytic in $\varepsilon$ for all $k$, and the series in (2.25) converges absolutely and is analytic with values in Sobolev spaces.
2.3. The transformed Dirichlet-Neumann operator
This section is devoted to the proof of the following theorem.
Theorem 2.6.
For fixed $\alpha\neq 0$, denote $\beta=\alpha^{2}$. Fix any depth $h\in(0,\infty)$. There exists $\iota>0$ such that for each $\varepsilon\in(-\iota,\iota)$ there exists a bounded linear operator $\mathcal{G}_{\varepsilon,\beta}\in\mathcal{L}(H^{\frac{1}{2}}(\mathbb{T}),H^{-\frac{1}{2}}(\mathbb{T}))$ with the following properties. Let $s\geq\frac{1}{2}$.
  • (i)
    There exists $0<\varepsilon_{0}(s)\leq\iota$ such that for all $\varepsilon\in(-\varepsilon_{0}(s),\varepsilon_{0}(s))$ the operator $\mathcal{G}_{\varepsilon,\beta}$ is bounded from $\mathcal{L}(H^{s}(\mathbb{T})$ to $H^{s-1}(\mathbb{T}))$. In addition, we have
    $(\mathcal{G}_{\varepsilon,\beta}(\bar{f}(-\cdot)))(x)=\overline{(\mathcal{G}_{\varepsilon,\beta}f)(-x)}$ (2.33)
    and
    $\mathcal{G}_{0,\beta}=\sqrt{|D|^{2}+\beta}\ \tanh(h\sqrt{|D|^{2}+\beta}).$ (2.34)
    If $s=1$, the operator $\mathcal{G}_{\varepsilon,\beta}$ is self-adjoint on $L^{2}(\mathbb{T})$.
  • (ii)
    The mapping
    $(-\varepsilon_{0}(s),\varepsilon_{0}(s))\times(0,\infty)\ni(\varepsilon,\beta)\mapsto\mathcal{G}_{\varepsilon,\beta}\in\mathcal{L}(H^{s}(\mathbb{T}),H^{s-1}(\mathbb{T}))$ (2.35)
    is analytic.
  • (iii)
    If $s\geq 1$ and $f_{0}\in H^{s}(\mathbb{T})$, we have the identity
    $\mathcal{G}_{\varepsilon,\beta}(f_{0}\circ\zeta)(x)=e^{-i\alpha y}\zeta^{\prime}(x)\big{[}G(\eta^{*})(f_{0}e^{i\alpha y})\big{]}|_{(\zeta(x),y)}$ (2.36)
    for a.e. $x\in\mathbb{T}$ and for all $y\in\mathbb{R}$.
We emphasize that the operator $\mathcal{G}_{\varepsilon,\beta}$ and the constants $\iota$ and $\varepsilon_{0}(s)$ all depend on the depth $h$.
Remark 2.7.
In the infinite depth case, Theorem 2.6 was proven in Theorem 2.7, [17]. The assertions (i) and (iii) can be obtained similarly to the infinite depth case. On the other hand, the analyticity of $\mathcal{G}_{\varepsilon,\beta}$ in $(\varepsilon,\beta)$ is more difficult to establish due to the dependence on $\varepsilon$ of the flat domain $\Omega_{\varepsilon}$. We will be led to proving the analyticity of solutions to the second-order elliptic problem (2.44) where the highest order term has an $\varepsilon$-dependent coefficient. Moreover, the case $s\in[-\frac{1}{2},1]$ will be deduced by interpolation from the cases $s=-\frac{1}{2}$ and $s>1$, where $s=-\frac{1}{2}$ requires the nonstandard trace result in Lemma A.2.
Proof.
We assume without loss of generality that $\alpha>0$ and denote $\mathbb{T}^{2}_{\alpha}=\mathbb{T}\times(\mathbb{R}/2\pi\alpha\mathbb{Z})$. Let $(X,Z):\Omega_{\varepsilon}\to\Omega$ be the Riemann mapping in Theorem 2.4. It has the Jacobian $\mathcal{J}=\mathcal{J}(x,z,\varepsilon)=|\partial_{x}X|^{2}+|\partial_{z}X|^{2}$. For an arbitrary function $f\in H^{\frac{1}{2}}(\mathbb{T})$, we will define the function $\Theta_{f}$ by solving
$\begin{cases}\Delta_{x,z}\Theta_{f}-\beta\mathcal{J}\Theta_{f}=0\quad\text{ in}\Omega_{\varepsilon}=\{(x,z)\in\mathbb{R}^{2}:x\in\mathbb{T},~{}-h_{\varepsilon}<z<0\},\\ \Theta_{f}(x,0)=f(x),\\ \partial_{z}\Theta_{f}(x,-h_{\varepsilon})=0.\end{cases}$ (2.37)
Then we define the operator $\mathcal{G}_{\varepsilon,\beta}$ by
$\mathcal{G}_{\varepsilon,\beta}f=\partial_{z}\Theta_{f}(\cdot,0),$ (2.38)
where $\partial_{z}\Theta_{f}(\cdot,0)$ is the trace $\gamma(\nabla\Theta_{f})$ given in Proposition A.1. As shown in the proof of Theorem 2.7 in [17], for any $s\geq\frac{1}{2}$ there exists $\varepsilon_{0}(s)>0$ such that
$\mathcal{G}_{\varepsilon,\beta}\in\mathcal{L}(H^{s}(\mathbb{T}),H^{s-1}(\mathbb{T}))\quad\forall\varepsilon\in(-\varepsilon_{0}(s),\varepsilon_{0}(s));$ (2.39)
moreover, $\mathcal{G}_{\varepsilon,\beta}$ is self-adjoint on $L^{2}$ and it satisfies (2.33) thanks to the fact that $\eta^{*}$ is even.
In order to prove (2.36), we consider the problem
$\begin{cases}\Delta_{X,Y,Z}\vartheta=0\quad\text{in }\{(X,Y,Z)\in\mathbb{R}^{3}:(X,Y)\in\mathbb{T}^{2}_{\alpha},~{}-h<Z<\eta^{*}(X)\},\\ \vartheta(X,Y,\eta^{*}(X))=f_{0}(X)e^{i\alpha Y},\\ \partial_{Z}\vartheta(X,Y,-h)=0.\end{cases}$ (2.40)
For clarity, we are now denoting the physical variables by $(X,Y,Z)$. By (2.3), we have
$G(\eta^{*})(f_{0}e^{i\alpha Y})(X,Y)=\partial_{z}\vartheta(X,Y,\eta^{*}(X))-\partial_{x}\vartheta(X,Y,\eta^{*}(X))\partial_{x}\eta^{*}(X).$ (2.41)
The solution of (2.40) is given by $\vartheta(X,Y,Z)=\theta(X,Z)e^{i\alpha Y}$, where $\theta(X,Z)$ solves
$\begin{cases}\Delta_{X,Z}\theta-\alpha^{2}\theta=0\quad\text{in }\{(X,Z)\in\mathbb{R}^{2}:X\in\mathbb{T},~{}-h<Z<\eta^{*}(X)\},\\ \theta(X,\eta^{*}(X))=f_{0}(X),\\ \partial_{Z}\theta(X,-h)=0.\end{cases}$ (2.42)
Using the Riemann mapping $(X,Z):\Omega_{\varepsilon}\to\Omega$ in Proposition 2.4 to flatten the domain $\Omega$, we find that $\Theta(x,z):=\theta(X(x,z),Z(x,z))$ solves (2.37) with $f=f_{0}\circ\zeta$, i.e. $\Theta=\Theta_{f_{0}\circ\zeta}$. Then, for $s>1$, the identity (2.36) follows by using (2.41), the chain rule, and the Cauchy-Riemann equations. The case $s=1$ is treated by an approximation argument. We refer to the proof of Theorem 2.7 in [17] for further details.
Assume now that $\varepsilon=0$, i.e. $\eta^{*}=0$. Then we have $\mathcal{J}=1$ and $h_{\varepsilon}=h$ in view of (2.27). Consequently, by solving (2.37) using Fourier transform in $x$, we find
$\partial_{z}\Theta_{f}(\cdot,0)=\sqrt{|D|^{2}+\beta}\ \tanh(h\sqrt{|D|^{2}+\beta})f.$ (2.43)
This proves (2.34).
The remainder of this proof is devoted to the analyticity (2.35). We fix $s\geq\frac{1}{2}$ unless stated otherwise. We first note that since $\eta^{*}$ is analytic in $\varepsilon\in(-\varepsilon_{0}(s),\varepsilon_{0}(s))$ with values in Sobolev spaces, so are $\zeta$, $(X,Z)-\operatorname{I}$, $\mathcal{J}$, and $h_{\varepsilon}$. For fixed $f\in H^{s}(\mathbb{T})$, we let $\Theta(\varepsilon,\beta)\equiv\Theta(x,z;\varepsilon,\beta)$ be the solution of (2.37), which is posed on $\Omega_{\varepsilon}=\{(x,z):x\in\mathbb{T},~{}z\in(-h_{\varepsilon},0)\}$. Compared to the infinite depth case considered in [17], the extra complication comes from the $\varepsilon$-dependence of the domain $\Omega_{\varepsilon}$. To cope with this we make the simple rescaling
$\Theta(x,z;\varepsilon,\beta)=\Psi\left(x,\frac{h}{h_{\varepsilon}}z;\varepsilon,\beta\right),\quad\mathcal{J}(x,z;\varepsilon)=\tilde{J}\left(x,\frac{h}{h_{\varepsilon}}z;\varepsilon\right),$
so that $\Psi\equiv\Psi(\varepsilon,\beta):\Omega_{0}\to\mathbb{R}$ satisfies
$\begin{cases}\left(\partial_{x}^{2}+\big{(}\frac{h}{h_{\varepsilon}}\big{)}^{2}\partial_{z}^{2}\right)\Psi-\beta\tilde{J}\Psi=0\quad\text{ in }\Omega_{0}\ ,\\ \Psi(x,0)=f(x),\quad\partial_{z}\Psi(x,-h)=0,\end{cases}$ (2.44)
where $\Omega_{0}:=\{(x,z):x\in\mathbb{T},~{}z\in(-h,0)\}$. Thus the $\varepsilon$-dependence of $\Omega_{\varepsilon}$ is now encoded in the coefficient $\frac{h}{h_{\varepsilon}}$ of the highest order term $\partial_{z}^{2}\Psi$. Recalling (2.38) and using the chain rule, we obtain
$\mathcal{G}_{\varepsilon,\beta}f=\frac{h}{h_{\varepsilon}}\partial_{z}\Psi(\varepsilon,\beta)|_{z=0}=\frac{h}{h_{\varepsilon}}\tilde{\gamma}(\nabla\Psi(\varepsilon,\beta)),$ (2.45)
where $\tilde{\gamma}$ is defined in Lemma A.2. We divide the rest of the proof into several steps.
1. We claim that the mapping
$(-\varepsilon_{0}(s),\varepsilon_{0}(s))\times(0,\infty)\ni(\varepsilon,\beta)\mapsto\Psi(\varepsilon,\beta)\in H^{s+\frac{1}{2}}(\Omega_{0})$ (2.46)
is $C^{\infty}$ for $s\geq\frac{1}{2}$.
For sufficiently small $\varepsilon_{0}(s)$, we have $h_{\varepsilon}\geq\frac{1}{2}h$ and $\tilde{J}\geq\frac{1}{2}$. Then, elliptic estimates yield the uniform (in $\varepsilon$) bound
$\|\Psi(\varepsilon,\beta)\|_{H^{s+\frac{1}{2}}(\Omega_{0})}\leq C_{1}(\beta,s)\|f\|_{H^{s}(\mathbb{T})}\quad\forall\varepsilon\in(-\varepsilon_{0}(s),\varepsilon_{0}(s)),$ (2.47)
where $C_{1}(\beta,s)$ is bounded for $\beta$ in compact subsets of $(0,\infty)$. For fixed $\varepsilon\in(-\varepsilon_{0}(s),\varepsilon_{0}(s))$, the differentiability of $\Psi(\varepsilon,\beta)$ in $\beta$ can be proven analogously to the proof of Theorem 2.7, [17]. On the other hand, for fixed $\beta>0$, the differentiability of $\Psi(\varepsilon,\beta)$ in $\varepsilon$ requires more attention since $\varepsilon$ appears in the highest order term $\frac{h}{h_{\varepsilon}}\partial_{z}^{2}\Psi$. We let $u_{\varepsilon,\beta}$ be the solution of the inhomogeneous problem
$\begin{cases}(\partial_{x}^{2}+\big{(}\frac{h}{h_{\varepsilon}}\big{)}^{2}\partial_{z}^{2})u_{\varepsilon,\beta}-\beta\tilde{J}(\varepsilon)u_{\varepsilon,\beta}=\beta[\frac{d}{d\varepsilon}\tilde{J}(\varepsilon)]\Psi(\varepsilon,\beta)-\partial_{z}^{2}\Psi(\varepsilon,\beta)\frac{d}{d\varepsilon}\big{(}\frac{h}{h_{\varepsilon}}\big{)}^{2}:=H_{\varepsilon,\beta}\quad\text{in }\Omega_{0},\\ u_{\varepsilon,\beta}(\cdot,0)=0,\quad\partial_{z}u_{\varepsilon,\beta}(\cdot,-h)=0,\end{cases}$ (2.48)
which is obtained by formally differentiating (2.44) in $\varepsilon$.
We will prove that
$\|u_{\varepsilon,\beta}\|_{H^{s+\frac{1}{2}}}\leq C_{2}(\beta,s)\|f\|_{H^{s}(\mathbb{T})}\quad\forall\varepsilon\in(-\varepsilon_{0}(s),\varepsilon_{0}(s)),$ (2.49)
where $C_{2}(\beta,s)$ is independent of $\varepsilon$ and is bounded for $\beta$ in compact subsets of $(0,\infty)$. Consider first the smoother case $s\geq\frac{3}{2}$. Then, $s-\frac{3}{2}\geq 0$ and (2.47) implies that $\|H_{\varepsilon,\beta}\|_{H^{s-\frac{3}{2}}(\Omega_{0})}\leq C(\beta,s)\|f\|_{H^{s}(\mathbb{T})}$ uniformly in $\varepsilon$. Thus (2.49) follows from the standard elliptic estimates.
It remains to prove (2.49) for $s=\frac{1}{2}$, since the mapping $f\mapsto u_{\varepsilon,\beta}$ is linear so that we can use interpolation. We do know that $\|\Psi(\varepsilon,\beta)\|_{H^{1}(\Omega_{0})}\leq C_{1}\|f\|_{H^{\frac{1}{2}}(\mathbb{T})}$ as in (2.47). Suppose for the moment that $u=u_{\varepsilon,\beta}$ is a smooth solution of (2.48). Let us denote
$H^{1,0}(\Omega_{0})=\{\varphi\in H^{1}(\Omega_{0}):~{}u(\cdot,0)=0\},$
which is a Hilbert space with the inner product $(\varphi,\tilde{\varphi})_{H^{1,0}}:=\int_{\Omega_{0}}\nabla\varphi\cdot\nabla\tilde{\varphi}$. Multiplying (2.48) by $\varphi\in H^{1,0}(\Omega_{0})$ and integrating by parts using the homogeneous Neumann boundary condition in (2.48), we find that
[…]
$\quad+\left\langle\partial_{z}^{2}\Psi(\varepsilon,\beta)\frac{d}{d\varepsilon}\big{(}\frac{h}{h_{\varepsilon}}\big{)}^{2},\varphi\right\rangle_{(H^{1,0})^{\prime},H^{1,0}},$
where, since $\partial_{z}\Psi(\varepsilon,\beta)|_{z=-h}=0$, we denote
$\left\langle\partial_{z}^{2}\Psi(\varepsilon,\beta)\frac{d}{d\varepsilon}\big{(}\frac{h}{h_{\varepsilon}}\big{)}^{2},\varphi\right\rangle_{(H^{1,0})^{\prime},H^{1,0}}:=-\int_{\Omega_{0}}\partial_{z}\Psi(\varepsilon,\beta)\frac{d}{d\varepsilon}\big{(}\frac{h}{h_{\varepsilon}}\big{)}^{2}\partial_{z}\varphi dxdz.$
From the above considerations, we define $u\in H^{1,0}(\Omega_{0})$ to be a weak solution of (2.48) if
[…]
$\quad-\int_{\Omega_{0}}\partial_{z}\Psi(\varepsilon,\beta)\frac{d}{d\varepsilon}\big{(}\frac{h}{h_{\varepsilon}}\big{)}^{2}\partial_{z}\varphi dxdz\quad\forall\varphi\in H^{1,0}(\Omega_{0}).$
Since $\beta>0$, $h_{\varepsilon}\geq\frac{1}{2}h$, and $\tilde{J}\geq\frac{1}{2}$, the left-hand side of (2.50) is a coercive continuous bilinear form on $H^{1,0}(\Omega_{0})$. On the other hand, the right-hand side of (2.50) defines a continuous linear functional $T$ on $H^{1,0}(\Omega_{0})$ with norm bounded by
$\|T\|\leq C\|\Psi(\varepsilon,\beta)\|_{H^{1}(\Omega_{0})}\leq C^{\prime}\|f\|_{H^{\frac{1}{2}}}\ ,$
where $C$ and $C^{\prime}$ are independent of $\varepsilon$. The Lax-Milgram theorem then yields the existence and uniqueness of a weak solution $u_{\varepsilon,\beta}$ to (2.48). Moreover, it satisfies $\|u_{\varepsilon,\beta}\|_{H^{1}(\Omega)}\leq C\|T\|$, and thus (2.49) holds for $s=\frac{1}{2}$.
Now we fix $(\varepsilon,\beta)\in(-\varepsilon_{0}(s),\varepsilon_{0}(s))\times(0,\infty)$. For small $\mu$, the expression $v_{\mu}:=\frac{\Psi(\varepsilon+\mu,\beta)-\Psi(\varepsilon,\beta)}{\mu}-u_{\varepsilon,\beta}\in H^{1,0}(\Omega_{0})$ satisfies, in the weak sense analogously to (2.50),
$\begin{cases}(\partial_{x}^{2}+\big{(}\frac{h}{h_{\varepsilon}}\big{)}^{2}\partial_{z}^{2})v_{\mu}-\beta\tilde{J}(\varepsilon+\mu)v_{\mu}=F_{\mu}\quad\text{in }\Omega_{0},\\ v_{\mu}(\cdot,0)=0,\quad\partial_{z}v_{\mu}(\cdot,-h)=0,\end{cases}$ (2.51)
where
$F_{\mu}$ $=\beta\left[\frac{\tilde{J}(\varepsilon+\mu)-\tilde{J}(\varepsilon)}{\mu}-\frac{d}{d\varepsilon}\tilde{J}(\varepsilon)\right]\Psi(\varepsilon,\beta)+\beta\left[\tilde{J}(\varepsilon+\mu)-\tilde{J}(\varepsilon)\right]u_{\varepsilon,\beta}$ (2.52)
$\quad-\left[\frac{1}{\mu}\left(\Big{(}\frac{h}{h_{\varepsilon+\mu}}\Big{)}^{2}-\Big{(}\frac{h}{h_{\varepsilon}}\Big{)}^{2}\right)-\frac{d}{d\varepsilon}\big{(}\frac{h}{h_{\varepsilon}}\big{)}^{2}\right]\partial_{z}^{2}\Psi(\varepsilon,\beta)-\left(\Big{(}\frac{h}{h_{\varepsilon+\mu}}\Big{)}^{2}-\Big{(}\frac{h}{h_{\varepsilon}}\Big{)}^{2}\right)\partial_{z}^{2}u_{\varepsilon,\beta}.$
As above, we first consider the case $s\geq\frac{3}{2}$. Then the uniform estimates (2.47) and (2.49) imply
$\|F_{\mu}\|_{H^{s-\frac{3}{2}}(\Omega_{0})}\leq\mu C_{3}(\beta,s)\|f\|_{H^{s}(\mathbb{T})},\quad s\geq\frac{3}{2}.$ (2.53)
Consequently, elliptic regularity theory yields
$\|v_{\mu}\|_{H^{s+\frac{1}{2}}(\Omega_{0})}\leq\mu C_{4}(\beta,s)\|f\|_{H^{s}(\mathbb{T})},\quad s\geq\frac{3}{2}.$ (2.54)
Now consider the case $s=\frac{1}{2}$. Since $\partial_{z}\Psi(\varepsilon,\beta)|_{z=0}=\partial_{z}u_{\varepsilon,\beta}|_{z=0}=0$, we can argue as in the proof of (2.49) for $s=\frac{1}{2}$ to obtain (2.54) for $s=\frac{1}{2}$. Of course, because the mappings
$f\mapsto\Psi(\varepsilon,\beta)\mapsto u_{\varepsilon,\delta}\quad\text{and}\quad(\Psi(\varepsilon,\beta),u_{\varepsilon,\beta})\mapsto F_{\mu}\mapsto v_{\mu}$
are linear, the mapping $f\mapsto v_{\mu}$ is also linear. Thus by interpolation, (2.54) also holds for all $s\geq\frac{1}{2}$. Letting $\mu\to 0$ in (2.54), we conclude that the mapping (2.46) is differentiable in $\varepsilon\in(-\varepsilon_{0}(s),\varepsilon_{0}(s))$ for fixed $\beta\in(0,\infty)$. An induction argument shows that this mapping is $C^{\infty}$ in $(\varepsilon,\beta)$.
2. This step is devoted to the analyticity of $\Theta(\varepsilon,\beta)$. Fix an arbitrary number $\beta^{0}\in(0,\infty)$. We will prove the analyticity in a neighborhood of $\varepsilon=0$ and $\beta=\beta^{0}$. If we choose $\varepsilon_{0}(s)$ small enough, then $\tilde{J}(\varepsilon)$ is analytic in $\varepsilon$ with values in $H^{s+10}(\mathbb{T})$, say. Consequently, for any subinterval $I=[-a,a]\subset(-\varepsilon_{0}(s),\varepsilon_{0}(s))$, there exists a constant $M=M(I)>1$ such that for all $\varepsilon\in I$ we have the $\varepsilon$-expansions
$\tilde{J}(x,z;\varepsilon)=\sum_{j=0}^{\infty}\varepsilon^{j}\chi_{j}(x,z),\quad\big{(}\frac{h}{h_{\varepsilon}}\big{)}^{2}=\sum_{j=0}^{\infty}\varepsilon^{j}\nu_{j}$
with
$\|\chi_{j}\|_{H^{s+10}(\Omega_{\varepsilon})}\leq M^{j+1},\quad|\nu_{j}|\leq M^{j+1}.$ (2.55)
Moreover, $\chi_{0}=\nu_{0}=1$ and $\frac{h}{h_{\varepsilon}}\geq\frac{1}{2}$ if $\varepsilon_{0}(s)$ is small enough. Our goal in this part is to prove that $\Psi(\varepsilon,\beta)$ can be expanded in powers of $\varepsilon$ and $\beta-\beta^{0}$ as a convergent series
$\Psi(\varepsilon,\beta)=\sum_{j=0}^{\infty}\sum_{\ell=0}^{\infty}\varepsilon^{j}(\beta-\beta^{0})^{\ell}\Psi_{j,\ell}\quad\text{in }H^{s+\frac{1}{2}}(\Omega_{0})$ (2.56)
near $(\varepsilon,\beta)=(0,\beta^{0})$.
We begin formally. From (2.44) the coefficients $\Psi_{j,\ell}$ must formally satisfy the following problems. First, $\Psi_{0,0}$ satisfies
$\begin{cases}\Delta_{x,z}\Psi_{0,0}-\beta^{0}\Psi_{0,0}=0\quad\text{in~{}}\Omega_{0},\\ \Psi_{0,0}(x,0)=f(x),\quad\partial_{z}\Psi_{0,0}(x,-h)=0.\end{cases}$ (2.57)
Next, for $\ell=0$ and $j\geq 1$, $\Psi_{j,0}$ satisfies
$\begin{cases}\Delta_{x,z}\Psi_{j,0}-\beta^{0}\Psi_{j,0}=\beta^{0}\sum_{k=0}^{j-1}\chi_{j-k}\Psi_{k,0}-\sum_{k=0}^{j-1}\nu_{j-k}\partial_{z}^{2}\Psi_{k,0}\quad\text{in~{}}\Omega_{0},\\ \Psi_{j,0}(\cdot,0)=0,\quad\partial_{z}\Psi_{j,0}(\cdot,-h)=0.\end{cases}$ (2.58)
Finally, for $\ell\geq 1$ and $j\geq 1$, $\Psi_{j,\ell}$ satisfies
$\begin{cases}\Delta_{x,z}\Psi_{j,\ell}-\beta^{0}\Psi_{j,\ell}=\sum_{k=0}^{j}\chi_{j-k}\Psi_{k,\ell-1}+\beta^{0}\sum_{k=0}^{j-1}\chi_{j-k}\Psi_{k,\ell}-\sum_{k=0}^{j-1}\nu_{j-k}\partial_{z}^{2}\Psi_{k,\ell}\quad\text{in~{}}\Omega_{0},\\ \Psi_{j,\ell}(\cdot,0)=0,\quad\partial_{z}\Psi_{j,\ell}(\cdot,-h)=0\end{cases}$ (2.59)
with the convention that $\sum_{k=0}^{j-1}=0$ when $j=0$.
Now let us estimate $\Psi_{j,\ell}$ by elliptic theory. First, from (2.57) we have
$\|\Psi_{0,0}\|_{H^{s+\frac{1}{2}}(\Omega_{0})}\leq M_{0}\|f\|_{H^{s}(\mathbb{T})},\quad M_{0}=M_{0}(\beta^{0},s,h).$ (2.60)
Next, for any $s\geq\frac{1}{2}$ the problems (2.57), (2.58) and (2.59) should be understood in the weak sense analogous to (2.50). Precisely, $\Psi_{0,0}$ is the unique function in $H^{1}(\Omega)$ satisfying $\Psi_{0,0}|_{z=0}=f$ and
$-\int_{\Omega_{0}}\partial_{x}\Psi_{0,0}\cdot\partial_{x}\varphi+\partial_{z}\Psi_{0,0}\partial_{z}\varphi+\beta^{0}\Psi_{0,0}\varphi=0\quad\forall\varphi\in H^{1,0}(\Omega_{0}),$ (2.61)
$\Psi_{j,0}$ ($j\geq 1$) is the unique function in $H^{1,0}(\Omega_{0})$ satisfying
$-\int_{\Omega_{0}}\partial_{x}\Psi_{j,0}\cdot\partial_{x}\varphi+\partial_{z}\Psi_{j,0}\partial_{z}\varphi+\beta^{0}\Psi_{j,0}\varphi$ (2.62)
$\qquad=\int_{\Omega_{0}}\beta^{0}\sum_{k=0}^{j-1}\chi_{j-k}\Psi_{k,0}\varphi+\int_{\Omega_{0}}\sum_{k=0}^{j-1}\nu_{j-k}\partial_{z}\Psi_{k,0}\partial_{z}\varphi\quad\forall\varphi\in H^{1,0}(\Omega_{0}),$
and $\Psi_{j,\ell}$ ($\ell\geq 1$) is the unique function in $H^{1,0}(\Omega_{0})$ satisfying
$-\int_{\Omega_{0}}\partial_{x}\Psi_{j,\ell}\cdot\partial_{x}\varphi+\partial_{z}\Psi_{j,\ell}\partial_{z}\varphi+\beta^{0}\Psi_{j,\ell}\varphi$ (2.63)
$\qquad=\int_{\Omega_{0}}\left(\sum_{k=0}^{j}\chi_{j-k}\Psi_{k,\ell-1}+\beta^{0}\sum_{k=0}^{j-1}\chi_{j-k}\Psi_{k,\ell}\right)\varphi+\int_{\Omega_{0}}\sum_{k=0}^{j-1}\nu_{j-k}\partial_{z}\Psi_{k,\ell}\partial_{z}\varphi\quad\forall\varphi\in H^{1,0}(\Omega_{0}).$
By product estimates in Sobolev spaces, there exists a constant $M_{1}=M_{1}(\beta^{0},s,h)>1$ such that
$\|\Psi_{j,0}\|_{H^{s+\frac{1}{2}}(\Omega_{0})}\leq M_{1}\sum_{k=0}^{j-1}\big{(}\beta^{0}\|\chi_{j-k}\|_{H^{s+10}(\Omega_{0})}+|\nu_{j-k}|\big{)}\|\Psi_{k,0}\|_{H^{s+\frac{1}{2}}(\Omega_{0})}\quad\text{for~{}}j\geq 1$ (2.64)
and
$\|\Psi_{j,\ell}\|_{H^{s+\frac{1}{2}}(\Omega_{0})}$ $\leq M_{1}\sum_{k=0}^{j}\|\chi_{j-k}\|_{H^{s+10}(\Omega_{0})}\|\Psi_{k,\ell-1}\|_{H^{s+\frac{1}{2}}(\Omega_{0})}$ (2.65)
$\quad+M_{1}\sum_{k=0}^{j-1}\big{(}\beta^{0}\|\chi_{j-k}\|_{H^{s+10}(\Omega_{0})}+|\nu_{j-k}|\big{)}\|\Psi_{k,\ell}\|_{{}_{H^{s+\frac{1}{2}}(\Omega_{0})}}\quad\text{for~{}}\ell\geq 1.$
Lemma 2.8.
Choosing any number $A$ that satisfies
$A>2+M+M_{1}+(\beta^{0}+1)M_{1}M\ ,$ (2.66)
we have
$\|\Psi_{j,\ell}\|_{H^{s+\frac{1}{2}}(\Omega_{0})}\leq\|\Psi_{0,0}\|_{H^{s+\frac{1}{2}}(\Omega_{0})}(AM)^{j+\ell}\quad\text{if}\quad j+\ell\geq 1.$ (2.67)
Proof.
We will prove the claim (2.67) by induction on $\ell$ and $j$. We begin with $\ell=0$ and $j=1$. Combining (2.64), (2.55), and (2.66) yields
$\|\Psi_{1,0}\|_{H^{s+\frac{1}{2}}(\Omega_{0})}$ $\leq M_{1}(\beta^{0}\|\chi_{0}\|_{H^{s+10}(\Omega_{0})}+|\nu_{0}|)\|\Psi_{0,0}\|_{H^{s+\frac{1}{2}}(\Omega_{0})}$
$\leq\|\Psi_{0,0}\|_{H^{s+\frac{1}{2}}(\Omega_{0})}(\beta^{0}+1)M_{1}M\leq\|\Psi_{0,0}\|_{H^{s+\frac{1}{2}}(\Omega_{0})}(AM)\quad(M>1).$
This proves (2.67) for $(j,\ell)=(1,0)$. Letting $\ell=0$ and $j\geq 1$, we assume by induction on $j$ that (2.67) is valid up to $j-1$. Then we deduce from (2.64), (2.55), and (2.66) that
$\|\Psi_{j,0}\|_{H^{s+\frac{1}{2}}(\Omega_{0})}$ $\leq M_{1}\sum_{k=0}^{j-1}\big{(}\beta^{0}M^{j-k+1}+M^{j-k+1}\big{)}\|\Psi_{0,0}\|_{H^{s+\frac{1}{2}}(\Omega_{0})}(AM)^{k}$
$=\|\Psi_{0,0}\|_{H^{s+\frac{1}{2}}(\Omega_{0})}(AM)^{j}(\beta^{0}+1)M_{1}M\frac{1-A^{-j}}{A-1}$
$\leq\|\Psi_{0,0}\|_{H^{s+\frac{1}{2}}(\Omega_{0})}(AM)^{j}.$
Thus (2.67) holds for $\ell=0$ and $j\geq 1$. Next, we consider $\ell=1$. The case $j=0$ follows directly from (2.65) and the fact that $A>M_{1}$. Assume by induction on $j$ that (2.67) holds up to $j-1$, for $j\geq 2$. In view of (2.65), (2.55), (2.66), and the case $\ell=0$, we have
$\|\Psi_{j,1}\|_{H^{s+\frac{1}{2}}(\Omega_{0})}$ $\leq M_{1}\sum_{k=0}^{j}\|\chi_{j-k}\|_{H^{s+10}(\Omega_{0})}\|\Psi_{k,0}\|_{H^{s+\frac{1}{2}}(\Omega_{0})}$
$\quad+M_{1}\sum_{k=0}^{j-1}\big{(}\beta^{0}\|\chi_{j-k}\|_{H^{s+10}(\Omega_{0})}+|\nu_{j-k}|\big{)}\|\Psi_{k,1}\|_{{}_{H^{s+\frac{1}{2}}(\Omega_{0})}}$
$\leq M_{1}\sum_{k=0}^{j}M^{j-k+1}(AM)^{k}\|\Psi_{0,0}\|_{H^{s+\frac{1}{2}}(\Omega_{0})}$
$\quad+M_{1}\sum_{k=0}^{j-1}(\beta^{0}+1)M^{j-k+1}(AM)^{k+1}\|\Psi_{0,0}\|_{H^{s+\frac{1}{2}}(\Omega_{0})}\quad(\text{by induction})$
$=M_{1}M^{j+1}\|\Psi_{0,0}\|_{H^{s+\frac{1}{2}}(\Omega_{0})}\frac{A^{j+1}-1}{A-1}+(\beta^{0}+1)M_{1}M^{j+2}\frac{A^{j+1}-A}{A-1}$
$\leq(AM)^{j+1}\|\Psi_{0,0}\|_{H^{s+\frac{1}{2}}(\Omega_{0})}\big{(}M_{1}+(\beta^{0}+1)M_{1}M\big{)}\frac{1}{A-1}$
$\leq(AM)^{j+1}\|\Psi_{0,0}\|_{H^{s+\frac{1}{2}}(\Omega_{0})}.$
This proves (2.67) for $\ell=1$ and $j\geq 0$. The same reasoning allows us to prove by induction on $\ell$ that (2.67) holds for all $\ell\geq 1$. ∎
Due to (2.67) and (2.60), it follows that the series (2.56) converges in $H^{s+\frac{1}{2}}(\Omega_{0})$ provided $|\varepsilon|,~{}|\beta-\beta^{0}|<(AM)^{-1}$. Denote the sum of this series by $\tilde{\Psi}\in H^{s+\frac{1}{2}}(\Omega_{0})$. To conclude the proof of (2.56), it suffices show that $\tilde{\Psi}=\Psi$. For this purpose it suffices to consider the case $s=\frac{1}{2}$. Define the partial sums
$\Psi^{(m,n)}=\sum_{j=0}^{m}\sum_{\ell=0}^{n}\varepsilon^{j}(\beta-\beta^{0})^{\ell}\Psi_{j,\ell},\quad\chi^{(m)}=\sum_{j=0}^{m}\varepsilon^{j}\chi_{j},\quad H^{(m)}=\sum_{j=0}^{m}\varepsilon^{j}\nu_{j}.$ (2.68)
Using the fact that $\chi_{0}=\nu_{0}=1$, we deduce from the weak equations (2.61), (2.62) and (2.63) that
$-\int_{\Omega_{0}}\partial_{x}\Psi^{(m,n)}\partial_{x}\varphi=\int_{\Omega_{0}}(\beta-\beta^{0})\chi^{(m)}\Psi^{(m,n-1)}\varphi+\beta^{0}\chi^{(m)}\Psi^{(m,n)}\varphi+H^{(m)}\Psi^{(m,n)}\partial_{z}\varphi\quad\ \forall\varphi\in H^{1,0}(\Omega_{0}).$ (2.69)
Indeed, using the convention that $\sum_{k=0}^{j-1}=0$ for $j=0$, we have for all $\varphi\in H^{1,0}(\Omega_{0})$ that
$-\int_{\Omega}\partial_{x}\Psi^{(m,n)}\partial_{x}\varphi$ $=\int_{\Omega}\partial_{z}\Psi^{(m,n)}\partial_{z}\varphi+\int_{\Omega}\beta^{0}\Psi^{(m,n)}\varphi$
$\quad+\beta^{0}\int_{\Omega}\sum_{j=0}^{m}\varepsilon^{j}\sum_{k=0}^{j-1}\chi_{j-k}\Psi_{k,0}\varphi+\int_{\Omega}\sum_{j=0}^{m}\varepsilon^{j}\sum_{k=0}^{j-1}\nu_{j-k}\partial_{z}\Psi_{k,0}\partial_{z}\varphi$
$\quad+\int_{\Omega}\sum_{j=0}^{m}\sum_{\ell=1}^{n}\varepsilon^{j}(\beta-\beta^{0})^{\ell}\sum_{k=0}^{j}\chi_{j-k}\Psi_{k,\ell-1}\varphi$
$\quad+\beta^{0}\int_{\Omega}\sum_{j=0}^{m}\sum_{\ell=1}^{n}\varepsilon^{j}(\beta-\beta^{0})^{\ell}\sum_{k=0}^{j-1}\chi_{j-k}\Psi_{k,\ell}\varphi+\int_{\Omega}\sum_{j=0}^{m}\sum_{\ell=1}^{n}\varepsilon^{j}(\beta-\beta^{0})^{\ell}\sum_{k=0}^{j-1}\nu_{j-k}\partial_{z}\Psi_{k,\ell}\partial_{z}\varphi$
$=\int_{\Omega}\partial_{z}\Psi^{(m,n)}\partial_{z}\varphi+\int_{\Omega}\beta^{0}\Psi^{(m,n)}\varphi$
$\quad+\int_{\Omega}\sum_{j=0}^{m}\sum_{\ell=1}^{n}\varepsilon^{j}(\beta-\beta^{0})^{\ell}\sum_{k=0}^{j}\chi_{j-k}\Psi_{k,\ell-1}\varphi$
$\quad+\beta^{0}\int_{\Omega}\sum_{j=0}^{m}\sum_{\ell=0}^{n}\varepsilon^{j}(\beta-\beta^{0})^{\ell}\sum_{k=0}^{j-1}\chi_{j-k}\Psi_{k,\ell}\varphi+\int_{\Omega}\sum_{j=0}^{m}\sum_{\ell=0}^{n}\varepsilon^{j}(\beta-\beta^{0})^{\ell}\sum_{k=0}^{j-1}\nu_{j-k}\partial_{z}\Psi_{k,\ell}\partial_{z}\varphi$
$=(\beta-\beta^{0})\quad\int_{\Omega}\sum_{j=0}^{m}\sum_{\ell^{\prime}=1}^{n-1}\varepsilon^{j}(\beta-\beta^{0})^{\ell^{\prime}}\sum_{k=0}^{j}\chi_{j-k}\Psi_{k,\ell^{\prime}}\varphi$
$\quad+\beta^{0}\int_{\Omega}\sum_{j=0}^{m}\sum_{\ell=0}^{n}\varepsilon^{j}(\beta-\beta^{0})^{\ell}\sum_{k=0}^{j}\chi_{j-k}\Psi_{k,\ell}\varphi+\int_{\Omega}\sum_{j=0}^{m}\sum_{\ell=0}^{n}\varepsilon^{j}(\beta-\beta^{0})^{\ell}\sum_{k=0}^{j}\nu_{j-k}\partial_{z}\Psi_{k,\ell}\partial_{z}\varphi$
$=\int_{\Omega_{0}}(\beta-\beta^{0})\chi^{(m)}\Psi^{(m,n-1)}\varphi+\beta^{0}\chi^{(m)}\Psi^{(m,n)}\varphi+H^{(m)}\Psi^{(m,n)}\partial_{z}\varphi,$
which is (2.69). Letting $m,~{}n\to\infty$ in (2.69) yields
$-\int_{\Omega_{0}}\partial_{x}\tilde{\Psi}\partial_{x}\varphi=\int_{\Omega_{0}}(\beta-\beta^{0})\tilde{J}\tilde{\Psi}\varphi+\beta^{0}\tilde{J}\tilde{\Psi}\varphi+\big{(}\frac{h}{h_{\varepsilon}}\big{)}^{2}\partial_{z}\Psi\partial_{z}\varphi\quad\forall\varphi\in H^{1,0}(\Omega_{0}),$
and hence
$\int_{\Omega_{0}}\partial_{x}\tilde{\Psi}\partial_{x}\varphi+\big{(}\frac{h}{h_{\varepsilon}}\big{)}^{2}\partial_{z}\Psi\partial_{z}\varphi=\int_{\Omega_{0}}\beta\tilde{J}\tilde{\Psi}\varphi\quad\forall\varphi\in H^{1,0}(\Omega_{0}).$ (2.70)
By the continuity from $H^{1}(\Omega_{0})$ to $H^{\frac{1}{2}}(\mathbb{T})$ of the trace $\Psi_{j,\ell}\mapsto\Psi_{j,\ell}|_{z=0}$, we also have $\Psi|_{z=0}=\Psi_{0,0}|_{z=0}=f$. This and (2.70) imply that $\tilde{\Psi}$ is a weak solution of (2.44). Thus $\tilde{\Psi}=\Psi$ by the uniqueness of (2.44). This completes the proof of (2.56).
3. Now we make the connection to $\mathcal{G}_{\varepsilon,\beta}$, which is given by (2.45). By abuse of notation we may regard $\Psi(\varepsilon,\beta)$ and $\Psi_{j,\ell}$ as linear operators acting on $f$. It follows from (2.67) and (2.60) that
$\forall s\geq\frac{1}{2},\quad\|\Psi_{j,\ell}\|_{H^{s+\frac{1}{2}}(\Omega_{0})}\leq M_{0}(AM)^{j+\ell}\|f\|_{H^{s}(\mathbb{T})},$ (2.71)
where $M_{0}$, $M$ and $A$ depend only on $(\beta^{0},s,h)$.
In the smoother case $s>1$, we combine the analyticity (2.56) of $\Psi(\varepsilon,\beta)$ in $H^{s+\frac{1}{2}}(\Omega_{0})$ with the standard continuity from $H^{s-\frac{1}{2}}(\Omega_{0})$ to $H^{s-1}(\mathbb{T})$ of the trace of $\partial_{z}\Psi(\varepsilon,\beta)$ and $\partial_{z}\Psi_{j,\ell}$ onto $\{z=0\}$, to obtain
$\tfrac{h_{\varepsilon}}{h}\mathcal{G}_{\varepsilon,\beta}=\partial_{z}\Psi(\varepsilon,\beta)|_{z=0}=\sum_{j=0}^{\infty}\sum_{\ell=0}^{\infty}\varepsilon^{j}(\beta-\beta^{0})^{\ell}\partial_{z}\Psi_{j,\ell}|_{z=0}\quad\text{in }H^{s-1}(\mathbb{T}).$
Thus the mapping (2.35) is analytic. Because the trace theorem fails for $s\leq 1$, we shall provide a different argument for $s=\frac{1}{2}$ and then conclude the proof by interpolation.
Now we fix $s=\frac{1}{2}$. Then the coefficient $\Psi_{j,\ell}$ belongs to $H^{1}(\Omega_{0})$ and satisfies the bound (2.67). First, since $\Psi_{0,0}=\Theta|_{(\varepsilon,\beta)=(0,\beta^{0})}$, in view of the explicit formula (2.43), we have $\partial_{z}\Psi_{0,0}|_{z=0}\in H^{-\frac{1}{2}}(\mathbb{T})$ and $\|\partial_{z}\Psi_{0,0}|_{z=0}\|_{H^{-\frac{1}{2}}}\leq C\|f\|_{H^{\frac{1}{2}}}$. Next, for general $j$ and $k$ we will appeal to Lemma A.2 to make sense of $\partial_{z}\Psi_{j,k}|_{z=0}$. For brevity, we define
$w_{j,\ell}:=\sum_{k=0}^{j}\nu_{j-k}\Psi_{k,\ell}.$
In case $\ell=0$ and $j\geq 1$, we add $\Delta_{x,z}\sum_{k=0}^{j-1}\nu_{j-k}\Psi_{k,0}$ to both sides of (2.58) and use $\nu_{0}=\chi_{0}=1$ to obtain
$\Delta_{x,z}w_{j,0}=\beta^{0}\sum_{k=0}^{j}\chi_{j-k}\Psi_{k,0}+\partial_{x}^{2}\sum_{k=0}^{j-1}\nu_{j-k}\Psi_{k,0}\quad\text{in~{}}\Omega_{0}.$ (2.72)
More generally, for $\ell\geq 1$ we add $\Delta_{x,z}\sum_{k=0}^{j-1}\nu_{j-k}\Psi_{k,\ell}$ to both sides of (2.59) to obtain
$\Delta_{x,z}w_{j,\ell}=\sum_{k=0}^{j}\chi_{j-k}\Psi_{k,\ell-1}+\beta^{0}\sum_{k=0}^{j}\chi_{j-k}\Psi_{k,\ell}+\partial_{x}^{2}\sum_{k=0}^{j-1}\nu_{j-k}\Psi_{k,\ell}\quad\text{in~{}}\Omega_{0}.$ (2.73)
Hence, applying the estimate (A.7) in Lemma A.2 to $u=\nabla w_{j,\ell}$, we have $\tilde{\gamma}(\nabla w_{j,\ell})\in(H^{-\frac{1}{2}}(\mathbb{T}))^{2}$ and the estimate
$\|\tilde{\gamma}(\nabla w_{j,\ell})\|_{(H^{-\frac{1}{2}}(\mathbb{T}))^{2}}$ $\leq C\|\nabla w_{j,\ell}\|_{L^{2}(\Omega_{0})}+C\left\|\sum_{k=0}^{j}\chi_{j-k}\Psi_{k,\ell-1}+\beta^{0}\sum_{k=0}^{j}\chi_{j-k}\Psi_{k,\ell}\right\|_{L^{2}(\Omega_{0})}$
$\qquad+C\left\|\partial_{x}\sum_{k=0}^{j-1}\nu_{j-k}\Psi_{k,\ell}\right\|_{L^{2}(\Omega_{0})}$
with the convention that $\Psi_{j,-1}=0$, where $C$ depends only on $h$. Then, using the definition of $w_{j,\ell}$ and invoking the estimates (2.67) and (2.60) for $\|\Psi_{j,\ell}\|_{H^{1}(\Omega_{0})}$, as well as the bounds (2.55) for $\chi_{j-k}$ and $\nu_{j-k}$, we find
$\|\tilde{\gamma}(\nabla w_{j,\ell})\|_{(H^{-\frac{1}{2}}(\mathbb{T}))^{2}}\leq\frac{C_{1}}{A-1}\|f\|_{H^{\frac{1}{2}}(\mathbb{T})}(AM)^{j+\ell},\quad j,~{}\ell\geq 0,$ (2.74)
recalling that $A>1$ and $M>1$, where $C_{1}$ depends only $(\beta^{0},h)$.
From (2.60) we have $\|\Psi_{0,0}\|_{H^{1}(\Omega_{0})}\leq M_{0}\|f\|_{H^{\frac{1}{2}}(\mathbb{T})}$, $M_{0}=M_{0}(\beta^{0},h)$. Then, the first equation in (2.57) implies $\Delta\Psi_{0,0}\in H^{1}(\Omega_{0})\subset L^{2}(\Omega_{0})$, so that $\nabla\Psi_{0,0}\in H_{\operatorname{div}}(\Omega_{0})$. Therefore, we can apply Proposition A.1 and Lemma A.2 to have $\tilde{\gamma}(\nabla\Psi_{0,0})={\gamma}(\nabla\Psi_{0,0})$ and $\|\tilde{\gamma}(\nabla\Psi_{0,0})\|_{H^{-\frac{1}{2}}(\mathbb{T})}\leq C_{2}\|f\|_{H^{\frac{1}{2}}(\mathbb{T})}$ for some constant $C_{2}=C_{2}(\beta^{0},h)\geq C_{1}$.
Now we prove by induction on $j+\ell$ that
$\|\tilde{\gamma}(\nabla\Psi_{j,\ell})\|_{(H^{-\frac{1}{2}}(\mathbb{T}))^{2}}\leq C_{2}\|f\|_{H^{\frac{1}{2}}}(AM)^{j+\ell},\quad j,~{}\ell\geq 0.$ (2.75)
Indeed, the case $j+\ell=0$ is true. Suppose by induction that (2.75) holds for all pairs $(j,\ell)$ with $j+\ell\leq L-1$, and consider $j+\ell=L$. By the definition of $w_{j,\ell}$ we have $\Psi_{j,\ell}=w_{j,\ell}-\sum_{k=0}^{j-1}\nu_{j-k}\Psi_{k,\ell}$, so that (2.74) and the induction hypothesis yield
$\|\tilde{\gamma}(\nabla\Psi_{j,\ell})\|_{(H^{-\frac{1}{2}}(\mathbb{T}))^{2}}$ $\leq\|\tilde{\gamma}(\nabla w_{j,\ell})\|_{(H^{-\frac{1}{2}}(\mathbb{T}))^{2}}+\sum_{k=0}^{j-1}|\nu_{j-k}\|\tilde{\gamma}(\nabla\Psi_{k,\ell})\|_{(H^{-\frac{1}{2}}(\mathbb{T}))^{2}}$
$\leq\frac{C_{2}}{A-1}\|f\|_{H^{\frac{1}{2}}(\mathbb{T})}(AM)^{j+\ell}+\sum_{k=0}^{j-1}M^{j-k+1}C_{2}\|f\|_{H^{\frac{1}{2}}}(AM)^{k+\ell}$
$\leq\frac{C_{2}}{A-1}\|f\|_{H^{\frac{1}{2}}(\mathbb{T})}(AM)^{j+\ell}+\frac{C_{2}M}{A-1}\|f\|_{H^{\frac{1}{2}}}(AM)^{j+\ell}$
$\leq C_{2}\|f\|_{H^{\frac{1}{2}}}(AM)^{j+\ell},$
where in the last inequality we have used the fact from (2.66) that $A-1>1+M$. The proof of (2.75) is complete.
Combining (2.56) with (2.75), we deduce that
the series $\text{the series ~{}}\sum_{j=0}^{\infty}\sum_{\ell=0}^{\infty}\varepsilon^{j}(\beta-\beta^{0})^{\ell}\tilde{\gamma}(\nabla\Psi_{j,\ell})\text{~{}converges to some~{}}g~{}\text{in~{}}(H^{-\frac{1}{2}}(\mathbb{T}))^{2}$ (2.76)
provided $|\varepsilon|<(AM)^{-1}$ and $|\beta-\beta^{0}|<(AM)^{-1}$. We will show that $g=\tilde{\gamma}(\nabla\Psi(\varepsilon,\beta))$. To that end we define
$u_{m,n}:=\begin{bmatrix}\partial_{x}\Psi^{(m,n)}\\ H^{(m)}\partial_{z}\Psi^{(m,n)}\end{bmatrix}=H^{(m)}\nabla\Psi^{(m,n)}+\begin{bmatrix}(1-H^{(m)})\partial_{x}\Psi^{(m,n)}\\ 0\end{bmatrix},$ (2.77)
where $H^{(m)}$ is given in (2.68). Combining Lemma A.2 (i)(c) with (2.76), we obtain
[…]
On the other hand, by the first equations in (2.57), (2.58) and (2.59), we find
$\operatorname{div}u_{m,n}=\partial_{x}^{2}\Psi^{(m,n)}+H^{(m)}\partial_{z}^{2}\Psi^{(m,n)}=\beta^{0}\chi^{(m)}\Psi^{(m,n)}+(\beta-\beta^{0})\chi^{(m)}\Psi^{(m,n-1)}\in H^{1}(\Omega_{0}),$
which can be proven similarly to (2.69). In addition,
$u:=\begin{bmatrix}\partial_{x}\Psi(\varepsilon,\beta)\\ \big{(}\frac{h}{h_{\varepsilon}}\big{)}^{2}\partial_{z}\Psi(\varepsilon,\beta)\end{bmatrix}$
satisfies $\operatorname{div}u=\beta\tilde{J}\Psi(\varepsilon,\beta)\in H^{1}(\Omega_{0})$ in light of (2.44). It follows that $u_{m,n},\ u\in H_{\operatorname{div}}(\Omega_{0})$, and hence (A.2) implies
$\|\gamma(u_{m,n}-u)\|_{(H^{-\frac{1}{2}}(\mathbb{T}))^{2}}\leq C\|u_{m,n}-u\|_{L^{2}(\Omega_{0})}+C\|\operatorname{div}(u_{m,n}-u)\|_{L^{2}(\Omega_{0})}\to 0\quad\text{as~{}}m,n\to\infty,$
where we have used the fact from (2.56) that $\Psi^{(m,n)}\to\Psi(\varepsilon,\beta)$ in $H^{1}(\Omega_{0})$. Recalling our definition (A.6) of $\tilde{\gamma}$ on $H_{\operatorname{div}}+\tilde{H}_{\operatorname{div}}$, we obtain
$\lim_{m,n\to\infty}\tilde{\gamma}(u_{m,n})=\tilde{\gamma}(u)=\big{(}\frac{h}{h_{\varepsilon}}\big{)}^{2}\tilde{\gamma}(\nabla\Psi(\varepsilon,\beta))\quad\text{in~{}}(H^{-\frac{1}{2}}(\mathbb{T}))^{2},$
where the second equality follows from a decomposition analogous to (2.77).
In conjunction with (2.78) this yields $g=\tilde{\gamma}(\nabla\Psi(\varepsilon,\beta))$ as claimed. It means that
$\tilde{\gamma}(\nabla\Psi(\varepsilon,\beta))=\sum_{j=0}^{\infty}\sum_{\ell=0}^{\infty}\varepsilon^{j}(\beta-\beta^{0})^{\ell}\tilde{\gamma}(\nabla\Psi_{j,\ell})\quad\text{in~{}}(H^{-\frac{1}{2}}(\mathbb{T}))^{2}.$ (2.79)
Due to (2.45), this yields the analyticity of the mapping (2.35) when $s=\frac{1}{2}$.
Finally, in order to generalize to $s\geq\frac{1}{2}$, we shall interpolate. We recall (2.71) and view $M_{0}$, $A$ and $M$ as functions of $s\geq\frac{1}{2}$. We fix any $s_{0}>1$ and use the standard trace theorem to obtain
$\|\gamma(\nabla\Psi_{j,\ell})\|_{(H^{s_{0}-1}(\mathbb{T}))^{2}}\leq N(s_{0})\|\Psi_{j,\ell}\|_{H^{s_{0}+\frac{1}{2}}(\Omega_{0})}\leq N(s_{0})M_{0}(s_{0})\big{(}A(s_{0})M(s_{0})\big{)}^{j+\ell}\|f\|_{H^{s_{0}}(\mathbb{T})}.$
Thus we can view $\gamma(\nabla\Psi_{j,\ell})=\tilde{\gamma}(\nabla\Psi_{j,\ell})$ as an operator in $\mathcal{L}(H^{s_{0}}(\mathbb{T}),H^{s_{0}-1}(\mathbb{T}))^{2})$. In light of this inequality and (2.75), we obtain by interpolation that
$\|\tilde{\gamma}(\nabla\Psi_{j,\ell})\|_{\mathcal{L}(H^{s}(\mathbb{T}),(H^{s-1}(\mathbb{T}))^{2})}\leq C_{2}^{\kappa}\big{(}A(\tfrac{1}{2})M(\tfrac{1}{2})\big{)}^{\kappa(j+\ell)}\big{(}N_{0}(s_{0})M_{0}(s_{0})\big{)}^{1-\kappa}\big{(}A(s_{0})M(s_{0})\big{)}^{(1-\kappa)(j+\ell)}$
for any $s=\kappa\frac{1}{2}+(1-\kappa)s_{0}\in(\frac{1}{2},s_{0})$, $\kappa\in(0,1)$. Consequently the series $\sum_{j=0}^{\infty}\sum_{\ell=0}^{\infty}\varepsilon^{j}(\beta-\beta^{0})^{\ell}\tilde{\gamma}(\nabla\Psi_{j,\ell})$ converges in $(H^{s-1}(\mathbb{T}))^{2})$ provided $|\varepsilon|$ and $|\beta-\beta^{0}|$ are small enough. The sum of this series must equal $\tilde{\gamma}(\nabla\Psi(\varepsilon,\beta))$ thanks to (2.79). Therefore, the mapping (2.35) is analytic in a neighborhood of $\varepsilon=0$ and $\beta=\beta^{0}$ for any $s\geq\frac{1}{2}$. The proof of Theorem 2.6 is now complete.
2.4. The reformulated instability problem
Theorem 2.6 allows us to precisely formulate the instability problem as follows. Starting from the Riemann stretch (2.24), it is convenient to first define the two auxiliary operators
$\zeta_{\sharp}f(x)=(f\circ\zeta)(x),\quad\zeta_{*}f(x)=\zeta^{\prime}(x)\zeta_{\sharp}f(x).$ (2.80)
Then we define the new unknowns
$w_{1}(x)=\zeta_{*}\widetilde{v_{1}},\quad w_{2}(x)=\zeta_{\sharp}\widetilde{v_{2}}.$ (2.81)
We apply $\zeta_{*}$ to (2.20) and $\zeta_{\sharp}$ to (2.21), and use Theorem 2.6 (iii) to have
$\zeta^{\prime}(x)\big{[}G(\eta^{*})(\widetilde{v_{2}}e^{i\alpha y})\big{]}|_{(\zeta(x),y)}=e^{i\alpha y}\mathcal{G}_{\varepsilon,\beta}(\widetilde{v_{2}}\circ\zeta)(x),$
that is, $\zeta_{*}G(\eta^{*})(v_{2})(x,y)=e^{i\alpha y}\mathcal{G}_{\varepsilon,\beta}(w_{2})(x)$. We thus obtain the equivalent linearized system
$\partial_{t}w_{1}=\partial_{x}\big{(}p(x)w_{1}\big{)}+\mathcal{G}_{\varepsilon,\beta}w_{2},$ (2.82)
$\partial_{t}w_{2}=-\frac{1+q(x)}{\zeta^{\prime}(x)}w_{1}+p(x)\partial_{x}w_{2},$ (2.83)
where the variable coefficients
$p:=\frac{c^{*}-\zeta_{\sharp}V^{*}}{\zeta^{\prime}},\quad q:=-p\partial_{x}(\zeta_{\sharp}B^{*})$ (2.84)
have auxiliary dependence on $\varepsilon$ and $h$.
For the spectral analysis, we seek solutions of the linearized system (2.82)–(2.83) of the form $w_{j}(x,t)=e^{\lambda t}u_{j}(x)$, thereby arriving at the eigenvalue problem
$\lambda U=\mathcal{L}_{\varepsilon,\beta}U:=\begin{bmatrix}\partial_{x}(p(x)\cdot)&\mathcal{G}_{\varepsilon,\beta}\\ -\frac{1+q(x)}{\zeta^{\prime}(x)}&p(x)\partial_{x}\end{bmatrix}U,\quad U=\begin{bmatrix}u_{1}\\ u_{2}\end{bmatrix}.$ (2.85)
Of course $\mathcal{L}_{\varepsilon,\beta}$ depends also on the depth $h$ but we will drop this dependence to ease notation. By virtue of Theorem 2.6 (i), we have $\mathcal{G}_{\varepsilon,\beta}\in\mathcal{L}(H^{1}(\mathbb{T}),L^{2}(\mathbb{T}))$ is self-adjoint for $\varepsilon\in(-\varepsilon_{0}(1),\varepsilon_{0}(1))$, $\beta\in(0,\infty)$, and $h>0$. Therefore, we regard $\mathcal{L}_{\varepsilon,\beta}$, defined above, as a bounded linear operator from $(H^{1}(\mathbb{T}))^{2}$ to $(L^{2}(\mathbb{T}))^{2}$. Moreover, $\mathcal{L}_{\varepsilon,\beta}$ can be written in Hamiltonian form
$\mathcal{L}_{\varepsilon,\beta}=J\mathcal{H}_{\varepsilon,\beta},\quad J=\begin{bmatrix}0&1\\ -1&0\end{bmatrix},\quad\mathcal{H}_{\varepsilon,\beta}=\begin{bmatrix}\frac{1+q(x)}{\zeta^{\prime}(x)}&-p(x)\partial_{x}\\ \partial_{x}(p(x)\cdot)&\mathcal{G}_{\varepsilon,\beta}\end{bmatrix},$ (2.86)
where $\mathcal{H}_{\varepsilon,\beta}$ is self-adjoint. All entries of $\mathcal{H}_{\varepsilon,\beta}$ depend on $\varepsilon$ and $h$ but only the lower right corner depends on $\beta$. For the two-dimensional modulational instability, $\beta=0$ and it was shown in [32] that $\mathcal{G}_{\varepsilon,0}=|D|\tanh(h|D|)$ (see also Proposition 4.1 below). However, when $\beta\neq 0$, $\mathcal{G}_{\varepsilon,\beta}$ is no longer a Fourier multiplier but a genuine pseudo-differential operator, which will be expanded up to $O(\varepsilon^{3})$ in Section 4.1.
Definition 2.9.
The Stokes wave $(\eta^{*},\psi^{*},c^{*})$ is said to be unstable with respect to transverse perturbations if there exists $\beta>0$ (and, hence, a transverse wave number $\alpha=\sqrt{\beta}$) such that $\mathcal{L}_{\varepsilon,\beta}$ has an eigenvalue with positive real part.
The expansions in powers of $\varepsilon$ for the conformal depth $h_{\varepsilon}$, Riemann stretch $\zeta$, and the coefficients in (2.82)–(2.83) are given in the next proposition.
Proposition 2.10.
The conformal depth $h_{\varepsilon}$ is real analytic in $\varepsilon$ with expansion
$h_{\varepsilon}=h+h_{2}\varepsilon^{2}+O\left(\varepsilon^{4}\right),$ (2.87)
where
$h_{2}=\frac{c_{0}^{4}-3}{4c_{0}^{2}}\quad\textrm{and}\quad c_{0}=\tanh^{\frac{1}{2}}(h).$ (2.88)
Moreover, for any $s>\frac{5}{2}$, the following functions are analytic in $\varepsilon$ with values in $H^{s}(\mathbb{T})$ provided $\varepsilon$ is sufficiently small. They have the expansions:
$\zeta(x)=x+\varepsilon\zeta_{1,1}\sin x+\varepsilon^{2}\zeta_{2,2}\sin(2x)+\varepsilon^{3}\Big{[}\zeta_{3,1}\sin x+\zeta_{3,3}\sin(3x)\Big{]}+O(\varepsilon^{4}),$ (2.89)
$p(x)=c_{0}+\varepsilon p_{1,1}\cos x+\varepsilon^{2}\Big{[}p_{2,0}+p_{2,2}\cos(2x)\Big{]}+\varepsilon^{3}\Big{[}p_{3,1}\cos x+p_{3,3}\cos(3x)\Big{]}+O(\varepsilon^{4}),$ (2.90)
$q(x)=\varepsilon q_{1,1}\cos x+\varepsilon^{2}\Big{[}q_{2,0}+q_{2,2}\cos(2x)\Big{]}+\varepsilon^{3}\Big{[}q_{3,1}\cos x+q_{3,3}\cos(3x)\Big{]}+O(\varepsilon^{4}),$ (2.91)
$\frac{1+q(x)}{\zeta^{\prime}(x)}=1+\varepsilon r_{1,1}\cos x+\varepsilon^{2}\Big{[}r_{2,0}+r_{2,2}\cos(2x)\Big{]}+\varepsilon^{3}\Big{[}r_{3,1}\cos x+r_{3,3}\cos(3x)\Big{]}+O(\varepsilon^{4}),$ (2.92)
where
$\zeta_{1,1}$ $=c_{0}^{-2},$ (2.93)
$\zeta_{2,2}$ $=\frac{c_{0}^{8}+4c_{0}^{4}+3}{8c_{0}^{8}},$ (2.94)
$\zeta_{3,1}$ $=\frac{4c_{0}^{14}+2c_{0}^{12}-17c_{0}^{10}-14c_{0}^{8}+10c_{0}^{6}+10c_{0}^{4}-15c_{0}^{2}-12}{16c_{0}^{10}(c_{0}^{2}+1)}$ (2.95)
$\zeta_{3,3}$ $=\frac{3c_{0}^{12}+43c_{0}^{8}+41c_{0}^{4}+9}{64c_{0}^{14}}$ (2.96)
$p_{1,1}$ $=-2c_{0}^{-1},$ (2.97)
$p_{2,0}$ $=\frac{-2c_{0}^{12}+5c_{0}^{8}+12c_{0}^{4}+9}{16c_{0}^{7}},$ (2.98)
$p_{2,2}$ $=-\frac{c_{0}^{4}+3}{2c_{0}^{7}},$ (2.99)
$p_{3,1}$ $=\frac{-2c_{0}^{14}+14c_{0}^{10}+11c_{0}^{8}-10c_{0}^{6}-10c_{0}^{4}+24c_{0}^{2}+21}{8c_{0}^{9}(c_{0}^{2}+1)},$ (2.100)
$p_{3,3}$ $=-\frac{c_{0}^{12}+17c_{0}^{8}+51c_{0}^{4}+27}{32c_{0}^{13}}$ (2.101)
$q_{1,1}$ $=-c_{0}^{2},$ (2.102)
$q_{2,0}$ $=1,$ (2.103)
$q_{2,2}$ $=2-3c_{0}^{-4},$ (2.104)
$q_{3,1}$ $=\frac{4c_{0}^{14}+6c_{0}^{12}-9c_{0}^{10}-12c_{0}^{8}-30c_{0}^{6}-30c_{0}^{4}+69c_{0}^{2}+66}{16c_{0}^{6}(c_{0}^{2}+1)},$ (2.105)
$q_{3,3}$ $=-\frac{3(3c_{0}^{12}+19c_{0}^{8}-71c_{0}^{4}+81)}{64c_{0}^{10}},$ (2.106)
$r_{1,1}$ $=-(c_{0}^{2}+c_{0}^{-2}),$ (2.107)
$r_{2,0}$ $=\frac{3}{2}+\frac{1}{2}c_{0}^{-4},$ (2.108)
$r_{2,2}$ $=\frac{9c_{0}^{8}-14c_{0}^{4}-3}{4c_{0}^{8}},$ (2.109)
$r_{3,1}$ $=\frac{4c_{0}^{18}+6c_{0}^{16}-11c_{0}^{14}-12c_{0}^{12}-45c_{0}^{10}-48c_{0}^{8}+93c_{0}^{6}+90c_{0}^{4}+27c_{0}^{2}+24}{16c_{0}^{10}(c_{0}^{2}+1)},$ (2.110)
$r_{3,3}$ $=\frac{-c_{0}^{16}-98c_{0}^{12}+252c_{0}^{8}-318c_{0}^{4}-27}{64c_{0}^{14}}.$ (2.111)
Proof.
The proof of Proposition 2.10 is given in Appendix A.2 of [7], noting that their $f|_{\varepsilon}$, $\mathfrak{p}_{\varepsilon}$, $p_{\varepsilon}$, and $a_{\varepsilon}$ correspond to $(h_{\varepsilon}-h)$, $(\zeta-x)$, $(p-c_{0})$, and $[(1+q)/\zeta^{\prime}-1]$, respectively, in this work. ∎
Combining Theorem 2.6 (with $s=1$) and Proposition 2.10, we conclude that the mapping
$(-\varepsilon_{0}(1),\varepsilon_{0}(1))\times(0,\infty)\ni(\varepsilon,\beta)\mapsto\mathcal{L}_{\varepsilon,\beta}\in\mathcal{L}\big{(}(H^{1}(\mathbb{T}))^{2},(L^{2}(T))^{2}\big{)}$ (2.112)
is analytic for each $h>0$.
3. Resonance condition and Kato’s perturbed basis
3.1. Resonance condition
We begin with the basic case $\varepsilon=0$. Using the expansions (2.90) and (2.91) for $p$ and $q$ and the formula (2.34) for $\mathcal{G}_{0,\beta}$, we find in view of (2.85) that
$\mathcal{L}_{0,\beta}=\begin{bmatrix}c_{0}\partial_{x}&(|D|^{2}+\beta)^{\frac{1}{2}}\tanh\left(h\left(|D|^{2}+\beta\right)^{\frac{1}{2}}\right)\\ -1&c_{0}\partial_{x}\end{bmatrix}.$ (3.1)
The spectrum of $\mathcal{L}_{0,\beta}$ consists of the purely imaginary eigenvalues
$\lambda^{0}_{\pm}(k,\beta)=i\left[c_{0}k\pm(k^{2}+\beta)^{\frac{1}{4}}\tanh^{\frac{1}{2}}\left(h\left(k^{2}+\beta\right)^{\frac{1}{2}}\right)\right],\quad k\in\mathbb{Z},$ (3.2)
which are the roots of the quadratic characteristic polynomial
$\Delta_{0}(\lambda;k,\beta)=(\lambda-ic_{0}k)^{2}+(k^{2}+\beta)^{\frac{1}{2}}\tanh\left(h\left(|D|^{2}+\beta\right)^{\frac{1}{2}}\right)=[\lambda-\lambda^{0}_{+}(k,\beta)][\lambda-\lambda^{0}_{-}(k,\beta)].$ (3.3)
We learn from [30] that the resonance condition is a double eigenvalue $\lambda^{0}_{+}(-(m+1),\beta)=\lambda^{0}_{-}(m,\beta)$. That is,
$-c_{0}(m+1)+((m+1)^{2}+\beta)^{\frac{1}{4}}\tanh^{\frac{1}{2}}\left(h\left((m+1)^{2}+\beta\right)^{\frac{1}{2}}\right)=c_{0}m-(m^{2}+\beta)^{\frac{1}{4}}\tanh^{\frac{1}{2}}\left(h\left(m^{2}+\beta\right)^{\frac{1}{2}}\right).$ (3.4)
For our purposes, we choose the simplest resonance $m=1$. Then, we have the following proposition on the existence, uniqueness and asymptotics of the resonant value of $\beta$.
Proposition 3.1.
For each $h>0$, there exists a unique $\beta_{*}=\beta_{*}(h)\in(0,\infty)$ such that (3.4) is satisfied and
(i) $30<\beta_{*}(h)<$ 3,
(ii) $\beta_{*}(h)$ is real analytic for $h>0$,
(iii) $\beta_{*}(h)=\frac{4}{3}h^{2}+\mathscr{o}(h^{2})$ as $h\rightarrow 0^{+}$, and
(iv) $\beta_{*}(h)=\beta_{*,\infty}-\frac{12e^{-2h}}{\left(1+\beta_{*,\infty}\right)^{-\frac{3}{4}}+\left(4+\beta_{*,\infty}\right)^{-\frac{3}{4}}}+\mathscr{o}\left(e^{-2h}\right)$ as $h\rightarrow\infty$, where $\beta_{*,\infty}=2.7275...$ is the root of (3.7).
Proof.
Define the function
$F(\beta,h)=3\tanh^{\frac{1}{2}}(h)-(1+\beta)^{\frac{1}{4}}\tanh^{\frac{1}{2}}\left(h\left(1+\beta\right)^{\frac{1}{2}}\right)-(4+\beta)^{\frac{1}{4}}\tanh^{\frac{1}{2}}\left(h\left(4+\beta\right)^{\frac{1}{2}}\right)$ (3.5)
for $\beta,\ h>0$. For fixed $h>0$, the zeros of $F(\cdot,h)$ correspond to the solution of the resonance condition (3.4) with $m=1$. We note that $F$ is real analytic for $\beta>0,\ h>0$. We have
$F(0,h)=2\tanh^{\frac{1}{2}}(h)-2^{\frac{1}{2}}\tanh^{\frac{1}{2}}\left(2h\right)>0,$
since $\tanh(2h)/\tanh(h)<2$ valid for $h>0$. On the other hand, for any $h>0$,
$F(3,h)$ $=3\tanh^{\frac{1}{2}}(h)-4^{\frac{1}{4}}\tanh^{\frac{1}{2}}(2h)-7^{\frac{1}{4}}\tanh^{\frac{1}{2}}\left(7^{\frac{1}{2}}h\right),$
$<3\tanh^{\frac{1}{2}}(h)-4^{\frac{1}{4}}\tanh^{\frac{1}{2}}(h)-7^{\frac{1}{4}}\tanh^{\frac{1}{2}}(h),$
$<\left(3-4^{\frac{1}{4}}-7^{\frac{1}{4}}\right)\tanh^{\frac{1}{2}}(h),$
$<0,$
where the first inequality follows from the monotonicity of $\tanh^{\frac{1}{2}}(\xi)$ for $\xi>0$. Consequently, for each $h>0$, there exists $30<\beta_{*}<$ 3 such that $F(\beta_{*},h)=0$. Because $\partial_{\beta}F<0$, we deduce that $\beta_{*}(h)$ is the unique positive zero of $F(\cdot,h)$ This proves (i). By the Analytic Implicit Function Theorem, $\beta_{*}(h)$ is real analytic for $h>0$, proving (ii).
We now prove (iii). By (i), $h(1+\beta_{*}(h))^{\frac{1}{2}}=O(h)$ and $h(4+\beta_{*}(h))^{\frac{1}{2}}=O(h)$ as $h\rightarrow 0^{+}$ or $h\to\infty$. Consequently, using the expansion
$\tanh^{\frac{1}{2}}(\xi)=\xi^{\frac{1}{2}}\left(1-\frac{\xi^{2}}{6}+O\left(\xi^{4}\right)\right)\quad\textrm{as}\quad\xi\rightarrow 0^{+},$
we obtain
$F(\beta_{*}(h),h)$ $=\left[3-\left(1+\beta_{*}(h)\right)^{\frac{1}{2}}-\left(4+\beta_{*}(h)\right)^{\frac{1}{2}}\right]h^{\frac{1}{2}}-\frac{1}{6}\left[3-\left(1+\beta_{*}(h)\right)^{\frac{3}{2}}-\left(4+\beta_{*}(h)\right)^{\frac{3}{2}}\right]h^{\frac{5}{2}}$
$\quad+O\left(h^{\frac{9}{2}}\right)\quad\textrm{as}\quad h\rightarrow 0^{+}.$ (3.6)
Given $F(\beta_{*}(h),h)=0$, it follows from (3.6) that
$3-\left(1+\beta_{*}(h)\right)^{\frac{1}{2}}-\left(4+\beta_{*}(h)\right)^{\frac{1}{2}}=O(h^{2})\quad\textrm{as}\quad h\rightarrow 0^{+}.$
Rearranging this result, we have
$3-(1+\beta_{*}(h))^{\frac{1}{2}}+O(h^{2})$ $=(4+\beta_{*}(h))^{\frac{1}{2}},$
and upon squaring both sides, while using the bound $30<\beta_{*}(h)<$ 3, we find
$\left(1+\beta_{*}(h)\right)^{\frac{1}{2}}=1+O(h^{2})\quad\textrm{as}\quad h\rightarrow 0^{+}.$
Squaring both sides a final time, we conclude
$\beta_{*}(h)=O(h^{2})\quad\textrm{as}\quad h\rightarrow 0^{+}.$
In light of this result, we now posit the expansion $\beta_{*}(h)=\beta_{2}h^{2}+\mathscr{o}(h^{2})$ as $h\rightarrow 0^{+}$ for some $\beta_{2}$ to be determined. Substituting this expansion into (3.6) and setting $F(\beta_{*}(h),h)=0$, we find at leading order that $\beta_{2}=4/3$, proving (iii).
In order to prove (iv), we claim that there exists $H_{0}>0$ such that $\beta_{*}^{\prime}(h)>0$ for all $h>H_{0}$. Indeed, we calculate
$\partial_{h}F(\beta,h)=\left(\frac{3\textrm{sech}^{2}(h)}{2\tanh^{\frac{1}{2}}(h)}\right)g(\beta,h),$
where
$g(\beta,h)$ $=1-\frac{(1+\beta)^{\frac{3}{4}}}{3}\left(\frac{\textrm{sech}\left(h\left(1+\beta\right)^{\frac{1}{2}}\right)}{\textrm{sech}(h)}\right)^{2}\left(\frac{\tanh(h)}{\tanh\left(h\left(1+\beta\right)^{\frac{1}{2}}\right)}\right)^{\frac{1}{2}}$
$\quad-\frac{(4+\beta)^{\frac{3}{4}}}{3}\left(\frac{\textrm{sech}\left(h\left(4+\beta\right)^{\frac{1}{2}}\right)}{\textrm{sech}(h)}\right)^{2}\left(\frac{\tanh(h)}{\tanh\left(h\left(4+\beta\right)^{\frac{1}{2}}\right)}\right)^{\frac{1}{2}}.$
From the monotonicity of $\tanh$, we have $\tanh(h)/\tanh(wh)<1$ for $h>0,\ w>1$. Thus
$g(\beta,h)>1-\frac{(1+\beta)^{\frac{3}{4}}}{3}\left(\frac{\textrm{sech}\left(h\left(1+\beta\right)^{\frac{1}{2}}\right)}{\textrm{sech}(h)}\right)^{2}-\frac{(4+\beta)^{\frac{3}{4}}}{3}\left(\frac{\textrm{sech}\left(h\left(4+\beta\right)^{\frac{1}{2}}\right)}{\textrm{sech}(h)}\right)^{2}.$
For $30<\beta<$ 3, we have the bounds
$\frac{(1+\beta)^{\frac{3}{4}}}{3}\left(\frac{\textrm{sech}\left(h\left(1+\beta\right)^{\frac{1}{2}}\right)}{\textrm{sech}(h)}\right)^{2}<\frac{4^{\frac{3}{4}}}{3}\quad\textrm{and}\quad\frac{(4+\beta)^{\frac{3}{4}}}{3}\left(\frac{\textrm{sech}\left(h\left(4+\beta\right)^{\frac{1}{2}}\right)}{\textrm{sech}(h)}\right)^{2}<\frac{7^{\frac{3}{4}}}{3}\left(\frac{\textrm{sech}(2h)}{\textrm{sech}(h)}\right)^{2},$
which follow from the monotonicty of $\textrm{sech}(\xi)$ for $\xi>0$. It follows that
$g(\beta,h)>1-\frac{4^{\frac{3}{4}}}{3}-\frac{7^{\frac{3}{4}}}{3}\left(\frac{\textrm{sech}(2h)}{\textrm{sech}(h)}\right)^{2}.$
Now $01-4^{\frac{3}{4}}/3>$ 0 and $\textrm{sech}(2h)/\textrm{sech}(h)\rightarrow 0$ as $h\rightarrow\infty$, so that there exists $H_{0}>0$ such that
$1-\frac{4^{\frac{3}{4}}}{3}-\frac{7^{\frac{3}{4}}}{3}\left(\frac{\textrm{sech}(2h)}{\textrm{sech}(h)}\right)^{2}>0\quad\forall h>H_{0}.$
Consequently, $\partial_{h}F(\beta,h)>0$ for all $30<\beta<$ 3 and $h>H_{0}$. In particular, $(\partial_{h}F)(\beta_{*}(h),h)>0$ for $h>H_{0}$. However, $\partial_{\beta}F(\beta,h)<0$ for all $\beta>0$ and $h>0$, so that $\beta_{*}^{\prime}(h)=-\frac{(\partial_{h}F)(\beta_{*}(h),h)}{(\partial_{\beta}F)(\beta_{*}(h),h)}>0$ for $h>H_{0}$, proving our claim.
Because $\beta_{*}(h)$ is bounded above and increasing for $h>H_{0}$, we know that $\beta_{*,\infty}:=\lim_{h\rightarrow\infty}\beta_{*}(h)$ exists and is finite. Taking the limit of $F(\beta_{*}(h),h)=0$ as $h\rightarrow\infty$, we arrive at the following equation for $\beta_{*,\infty}$:
$0\displaystyle 3-(1+\beta_{*,\infty})^{\frac{1}{4}}-(4+\beta_{*,\infty})^{\frac{1}{4}}=$ 0 . (3.7)
The solution of (3.7) is
$\beta_{*,\infty}=\frac{1}{48}\left[-687+\left(70498161-7872768\sqrt{7}\right)^{\frac{1}{3}}+3\left(2611043+291584\sqrt{7}\right)^{\frac{1}{3}}\right]=2.7275...,$
consistent with the infinite depth results presented in [17]. From the expansion
$\tanh^{\frac{1}{2}}(\xi)=1-e^{-2\xi}+O\left(e^{-4\xi}\right)\quad\textrm{as}\quad\xi\rightarrow\infty,$
the expression (3.5) has the asymptotic behavior
$F(\beta_{*}(h),h)$ $=3-\left(1+\beta_{*}(h)\right)^{\frac{1}{4}}-\left(4+\beta_{*}(h)\right)^{\frac{1}{4}}-3e^{-2h}+\mathscr{o}\left(e^{-2h}\right)\quad\textrm{as}\quad h\rightarrow\infty.$ (3.8)
Define $\underline{\beta}(h):=\beta_{*}(h)-\beta_{*,\infty}$. Then (3.8) becomes
$F(\beta_{*}(h),h)$ $=3-\left(1+\beta_{*,\infty}\right)^{\frac{1}{4}}\left(1+\frac{\underline{\beta}(h)}{1+\beta_{*,\infty}}\right)^{\frac{1}{4}}-\left(4+\beta_{*,\infty}\right)^{\frac{1}{4}}\left(1+\frac{\underline{\beta}(h)}{4+\beta_{*,\infty}}\right)^{\frac{1}{4}}-3e^{-2h}+\mathscr{o}\left(e^{-2h}\right),$
as $h\rightarrow\infty$. Since $\underline{\beta}(h)\rightarrow 0^{-}$ as $h\rightarrow\infty$, combining the expansion $\left(1+\xi\right)^{\frac{1}{4}}=1+\frac{1}{4}\xi+O\left(\xi^{2}\right)$ with (3.7) yields
$F(\beta_{*}(h),h))=-\frac{\underline{\beta}(h)}{4}\left(\left(1+\beta_{*,\infty}\right)^{-\frac{3}{4}}+\left(4+\beta_{*,\infty}\right)^{-\frac{3}{4}}\right)-3e^{-2h}+O\left(\underline{\beta}(h)^{2}\right)+\mathscr{o}\left(e^{-2h}\right)\quad\textrm{as}\quad h\rightarrow\infty.$
Because $F(\beta_{*}(h),h)=0$, we obtain the following implicit formula for $\underline{\beta}(h)$:
$\underline{\beta}(h)=\frac{-12e^{-2h}}{\left(1+\beta_{*,\infty}\right)^{-\frac{3}{4}}+\left(4+\beta_{*,\infty}\right)^{-\frac{3}{4}}}+O\left(\underline{\beta}(h)^{2}\right)+\mathscr{o}\left(e^{-2h}\right)\quad\textrm{as}\quad h\rightarrow\infty.$
Hence $\underline{\beta}(h)=O\left(e^{-2h}\right)$ as $h\rightarrow\infty$. Therefore, we obtain
$\underline{\beta}(h)=\frac{-12e^{-2h}}{\left(1+\beta_{*,\infty}\right)^{-\frac{3}{4}}+\left(4+\beta_{*,\infty}\right)^{-\frac{3}{4}}}+\mathscr{o}\left(e^{-2h}\right)\quad\textrm{as}\quad h\rightarrow\infty,$
from which (iv) follows. ∎
Figure 2. A plot of $\beta_{*}$ as a function of $h$ (solid blue) along with its asymptotic expansions as $h\rightarrow 0^{+}$ and $h\rightarrow\infty$ (dashed orange) according to Proposition  3.1.
Lemma 3.2.
For each $\beta>0$ and $h>0$, we have
$\frac{\partial}{\partial k}\textrm{Im}\left\{\lambda^{0}_{\pm}(k,\beta)\right\}>\begin{cases}c_{0}\left(1-(1+\beta)^{-\frac{1}{2}}\right)\quad\text{if}~{}|k|\geq 1,\\ 0\quad\text{if}~{}k\neq 0.\end{cases}$ (3.9)
Proof.
For ease of notation, we define the auxiliary functions
$\omega(\xi,h)=\left(\xi\tanh(h\xi)\right)^{\frac{1}{2}}\quad\text{and}\quad\xi(k,\beta)=(k^{2}+\beta)^{\frac{1}{2}},$
so that $\textrm{Im}\left\{\lambda^{0}_{\pm}(k,\beta)\right\}=c_{0}k\pm\omega(\xi(k,\beta),h)$. A direct calculation shows
$\frac{\partial}{\partial k}\omega(\xi(k,\beta),h)=\frac{\sqrt{h}}{2}\left[\sqrt{\frac{\tanh(h\xi)}{h\xi}}+\text{sech}^{\frac{3}{2}}(h\xi)\sqrt{\frac{h\xi}{\sinh(h\xi)}}\right]\frac{\partial}{\partial k}\xi(k,\beta)$
and $\left|\frac{\partial}{\partial k}\xi(k,\beta)\right|=\frac{|k|}{(k^{2}+\beta)^{\frac{1}{2}}}>0$ for $k\neq 0$. Using the inequalities $s/\sinh(s)<1$ and $\text{sech}^{\frac{3}{2}}(s)<\text{sech}(s)<\sqrt{\tanh(s)/s}$ for $s>0$, we have
$\left|\frac{\partial}{\partial k}\omega(\xi(k,\beta),h)\right|<\sqrt{h}\sqrt{\frac{\tanh(h\xi)}{h\xi}}\left|\frac{\partial}{\partial k}\xi(k,\beta)\right|\quad\forall k\neq 0.$
Since $\tanh(s)/s$ is decreasing for all $s>0$, we deduce
$\left|\frac{\partial}{\partial k}\omega(\xi(k,\beta),h)\right|<\sqrt{h}\sqrt{\frac{\tanh(h)}{h}}\left|\frac{\partial}{\partial k}\xi(k,\beta)\right|=c_{0}\left|\frac{\partial}{\partial k}\xi(k,\beta)\right|\quad\forall k\neq 0.$
Then, (3.9) follows from this and the inequalities
$0<\left|\frac{\partial}{\partial k}\xi(k,\beta)\right|\leq\begin{cases}(1+\beta)^{-\frac{1}{2}}\quad\text{if}~{}|k|\geq 1,\\ 1\quad\text{if}~{}k\neq 0.\end{cases}$
Given $\beta_{*}$, we define $\sigma$ by
$i\sigma:=\lambda^{0}_{+}(-2,\beta_{*},h)=\lambda^{0}_{-}(1,\beta_{*},h).$ (3.10)
The second inequality in (3.9) implies that the functions $k\mapsto\textrm{Im}\left\{\lambda^{0}_{\pm}(k,\beta)\right\}$ are strictly increasing on $\mathbb{R}$, and hence
$\Delta_{0}(i\sigma;k,\beta_{*})=0\iff k\in\{1,-2\}.$ (3.11)
Thus $i\sigma$ is a double eigenvalue of $\mathcal{L}_{0,\beta_{*}}$ for all $h>0$. Moreover, the first inequality in (3.9) implies that $i\sigma$ is isolated in the spectrum of $\mathcal{L}_{0,\beta_{*}}$. The eigenspace associated to $i\sigma$ is
$N(\mathcal{L}_{0,\beta_{*}}-i\sigma\operatorname{I})=\text{span}\{U_{1},U_{2}\},$ (3.12)
where $\operatorname{I}$ denotes the identity operator and
$U_{1}(x;h)$ $=\begin{bmatrix}i(1+\beta_{*})^{\frac{1}{4}}\tanh^{\frac{1}{2}}\left(h(1+\beta_{*})^{\frac{1}{2}}\right)\\ 1\end{bmatrix}e^{ix},$ (3.13)
$U_{2}(x;h)$ $=\begin{bmatrix}-i(4+\beta_{*})^{\frac{1}{4}}\tanh^{\frac{1}{2}}\left(h(4+\beta_{*})^{\frac{1}{2}}\right)\\ 1\end{bmatrix}e^{-2ix}.$ (3.14)
For ease of notation, we define
$\gamma_{j}:=(j^{2}+\beta_{*})^{\frac{1}{4}}\tanh^{\frac{1}{2}}\left(h\left(j^{2}+\beta_{*}\right)^{\frac{1}{2}}\right),$ (3.15)
so that
$\sigma=c_{0}-\gamma_{1}=-2c_{0}+\gamma_{2}.$ (3.16)
3.2. Kato’s perturbed basis
We fix any depth $h\in(0,\infty)$ and let $i\sigma$ be the isolated double eigenvalue (3.10) of $\mathcal{L}_{0,\beta_{*}}$. Then we fix $\Gamma$ a circle centered at $i\sigma$ with sufficiently small radius so that $\Gamma\subset\rho(\mathcal{L}_{\varepsilon,\beta})$ for $(\varepsilon,\beta)$ close to $(0,\beta_{*})$. We denote the spectral projection associated to $\Gamma$ by
[…]
The analyticity (2.112) of $\mathcal{L}_{\varepsilon,\beta}$ implies that $P_{\varepsilon,\delta}$ is analytic in $(\varepsilon,\delta)$ near $(0,0)$. Recall from (3.12) that the set $\{U_{1},U_{2}\}$ forms a basis of the range $\mathcal{V}_{0,0}:=R(P_{0,0})$. Our goal is to perturb the basis in a convenient way. Following Kato [23] and Berti et al [4], we define perturbing transformation operators $\mathcal{K}_{\varepsilon,\delta}$ as
$\mathcal{K}_{\varepsilon,\delta}:=\{1-(P_{\varepsilon,\delta}-P_{0,0})^{2}\}^{-\frac{1}{2}}\{P_{\varepsilon,\delta}P_{0,0}+(1-P_{\varepsilon,\delta})(1-P_{0,0})\}$ (3.18)
and a perturbed basis $\{U^{\varepsilon,\delta}_{1},U^{\varepsilon,\delta}_{2}\}$ by
$U^{\varepsilon,\delta}_{m}=\mathcal{K}_{\varepsilon,\delta}U_{m}\quad(m=1,2).$ (3.19)
Since $P_{\varepsilon,\delta}$ is analytic in $(\varepsilon,\delta)$, so are $\mathcal{K}_{\varepsilon,\delta}$ and $U^{\varepsilon,\delta}_{m}$. Moreover, when acting on $U_{m}$, the last term from (3.18) in $\mathcal{K}_{\varepsilon,\delta}U_{m}$ vanishes since $(1-P_{0,0})U_{m}=0$. By virtue of Lemma 3.1 in [4], the perturbed range $\mathcal{V}_{\varepsilon,\delta}:=R(P_{\varepsilon,\delta})$ is spanned by $\{U^{\varepsilon,\delta}_{1},U^{\varepsilon,\delta}_{2}\}$, and we have
$\mathcal{L}_{\varepsilon,\beta_{*}+\delta}:\mathcal{V}_{\varepsilon,\delta}\to\mathcal{V}_{\varepsilon,\delta}.$
By direct calculation, we obtain
$(JU_{1},U_{2})=(JU_{2},U_{1})=0,\quad(JU_{1},U_{1})=-i4\pi\gamma_{1},\quad(JU_{2},U_{2})=i4\pi\gamma_{2}.$ (3.20)
Since $\mathcal{K}_{\varepsilon,\delta}$ is symplectic, i.e. $\mathcal{K}_{\varepsilon,\delta}^{*}J\mathcal{K}_{\varepsilon,\delta}=J$ (see Lemma 3.2 in [4]), we have $(J\mathcal{K}_{\varepsilon,\delta}U,\mathcal{K}_{\varepsilon,\delta}V)=(JU,V)$. Consequently, (3.20) yields
$(JU_{1}^{\varepsilon,\delta},U_{2}^{\varepsilon,\delta})=(JU_{2}^{\varepsilon,\delta},U_{1}^{\varepsilon,\delta})=0,\quad(JU_{1}^{\varepsilon,\delta},U_{1}^{\varepsilon,\delta})=-i4\pi\gamma_{1},\quad(JU_{2}^{\varepsilon,\delta},U_{2}^{\varepsilon,\delta})=i4\pi\gamma_{2}.$ (3.21)
In view of (3.20) and (3.21), we normalize the eigenvectors according to
$V_{m}=\frac{1}{\sqrt{\gamma_{m}}}U_{m},\quad V_{m}^{\varepsilon,\delta}=\frac{1}{\sqrt{\gamma_{m}}}U^{\varepsilon,\delta}_{m}.$ (3.22)
Lemma 3.3.
Using our notation $\beta=\beta_{*}+\delta$, the $2\times 2$ matrix that represents the linear operator $\mathcal{L}_{\varepsilon,\beta}=J\mathcal{H}_{\varepsilon,\beta}:\mathcal{V}_{\varepsilon,\delta}\to\mathcal{V}_{\varepsilon,\delta}$ with respect to the basis $\{V_{1}^{\varepsilon,\delta},V_{2}^{\varepsilon,\delta}\}$ is
$\textrm{L}_{\varepsilon,\delta}=\begin{bmatrix}-\frac{i}{4\pi}(\mathcal{H}_{\varepsilon,\beta}V_{1}^{\varepsilon,\delta},V_{1}^{\varepsilon,\delta})&\frac{i}{4\pi}(\mathcal{H}_{\varepsilon,\beta}V_{1}^{\varepsilon,\delta},V_{2}^{\varepsilon,\delta})\\ -\frac{i}{4\pi}(\mathcal{H}_{\varepsilon,\beta}V_{2}^{\varepsilon,\delta},V_{1}^{\varepsilon,\delta})&\frac{i}{4\pi}(\mathcal{H}_{\varepsilon,\beta}U_{2}^{\varepsilon,\delta},V_{2}^{\varepsilon,\delta})\end{bmatrix}.$ (3.23)
$\textrm{L}_{\varepsilon,\delta}$ is purely imaginary and
$(\textrm{L}_{\varepsilon,\delta})_{12}=-(\textrm{L}_{\varepsilon,\delta})_{21}.$ (3.24)
Proof.
The claim (3.23) follows from (3.21) and the Hamiltonian form $\mathcal{L}_{\varepsilon,\beta}=J\mathcal{H}_{\varepsilon,\beta}$. We refer to the proof of Lemma 3.1 in [17] for details. The second claim follows from the property (2.33) of $\mathcal{G}_{\varepsilon,\beta}$ and the fact that the functions $p,q,\zeta^{\prime}$ are real and even; see the end of Section 3.2 in [17]. ∎
We have reduced the original transverse instability problem to the spectral analysis of the $2\times 2$ matrix $\textrm{L}_{\varepsilon,\delta}$. To capture the unstable eigenvalues of $\textrm{L}_{\varepsilon,\delta}$, we shall expand $\textrm{L}_{\varepsilon,\delta}$ to third order in $\varepsilon$ and $\delta$ in the next section.
4. Third-order expansions
4.1. Third-order expansions of $\mathcal{L}_{\varepsilon,\beta}$
We recall the Hamiltonian operator (2.86)
$\mathcal{H}_{\varepsilon,\beta}=\begin{bmatrix}\frac{1+q(x)}{\zeta^{\prime}(x)}&-p(x)\partial_{x}\\ \partial_{x}(p(x)\cdot)&\mathcal{G}_{\varepsilon,\beta}\end{bmatrix}.$
By virtue of Proposition 2.10 and Theorem 2.6 (ii), for each $h>0$, $\mathcal{H}_{\varepsilon,\beta}$ is analytic in $(\varepsilon,\beta)\in(-\varepsilon_{0}(1),\varepsilon_{0}(1))\times(0,\infty)$ with values in $\mathcal{L}(H^{1}(\mathbb{T}),L^{2}(\mathbb{T}))$. In particular, for fixed $\beta>0$, $\mathcal{G}_{\varepsilon,\beta}$ can be expanded in power series of $\varepsilon$ as
$\mathcal{G}_{\varepsilon,\beta}=\sum_{j=0}^{\infty}\varepsilon^{j}R_{j,\beta},\quad|\varepsilon|<\varepsilon_{0}(1).$ (4.1)
Then, invoking the expansions in Proposition 2.10, we obtain the third-order expansion in $\varepsilon$ of $\mathcal{H}_{\varepsilon,\beta}$:
$\mathcal{H}_{\varepsilon,\beta}=\sum_{j=0}^{3}\varepsilon^{j}\mathcal{H}^{j}_{\beta}+O(\varepsilon^{4}),$ (4.2)
where
$\mathcal{H}^{0}_{\beta}=\begin{bmatrix}1&-c_{0}\partial_{x}\\ c_{0}\partial_{x}&R_{0,\beta}\end{bmatrix},\quad c_{0}=\sqrt{\tanh(h)},$ (4.3)
$\mathcal{H}^{1}_{\beta}=\begin{bmatrix}r_{1,1}\cos x&-p_{1,1}\cos x\partial_{x}\\ \partial_{x}(p_{1,1}\cos x\cdot)&R_{1,\beta}\end{bmatrix},$
$\mathcal{H}^{2}_{\beta}=\begin{bmatrix}r_{2,0}+r_{2,2}\cos(2x)&-\big{[}p_{2,0}+p_{2,2}\cos(2x)\big{]}\partial_{x}\\ \partial_{x}\{[p_{2,0}+p_{2,2}\cos(2x)]\cdot)\}&R_{2,\beta}\end{bmatrix},$
$\mathcal{H}^{3}_{\beta}=\begin{bmatrix}r_{3,1}\cos x+r_{3,3}\cos(3x)&-\big{[}p_{3,1}\cos x+p_{3,3}\cos(3x)\big{]}\partial_{x}\\ \partial_{x}\left\{\big{[}p_{3,1}\cos x+p_{3,3}\cos(3x)\big{]}\cdot\right\}&R_{3,\beta}\end{bmatrix}.$
The coefficients $r_{1,1},\dots,p_{3,3}$ are given in Prop. 2.10. The nonlocal operators $R_{j,\beta}$ are Fourier multipliers. We explicitly compute them for $0\leq j\leq 3$ in the following proposition.
Proposition 4.1.
The Fourier multipliers $R_{j}$ for $0\leq j\leq 3$ take the form
$\widehat{R_{0,\beta}f}(k)={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}A^{0}_{k,\beta}}\widehat{f}(k)=\sqrt{|k|^{2}+\beta}\ \tanh(h\sqrt{|k|^{2}+\beta})\widehat{f}(k),$ (4.4)
$\widehat{R_{1,\beta}f}(k)={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}B^{-1}_{k,\beta}}\widehat{f}(k-1)+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}B^{1}_{k,\beta}}\widehat{f}(k+1),$ (4.5)
$\widehat{R_{2,\beta}f}(k)=C^{-2}_{k,\beta}\widehat{f}(k-2)+C^{0}_{k,\beta}\widehat{f}(k)+C^{2}_{k,\beta}\widehat{f}(k+2),$ (4.6)
$\widehat{R_{3,\beta}f}(k)=D^{-3}_{k,\beta}\widehat{f}(k-3)+D^{-1}_{k,\beta}\widehat{f}(k-1)+D^{1}_{k,\beta}\widehat{f}(k+1)+D^{3}_{k,\beta}\widehat{f}(k+3),$ (4.7)
where the coefficients $A^{s}_{k,\beta}$, $B^{s}_{k,\beta}$, $C^{s}_{k,\beta}$, and $D^{s}_{k,\beta}$ are explicit functions of $k$, $\beta$, and $h$ derived below.
Proof.
We write the expansions
$\zeta=x+\varepsilon\zeta^{1}+\varepsilon^{2}\zeta^{2}+\varepsilon^{3}\zeta^{3}+O(\varepsilon^{4}),\quad\eta^{*}=\varepsilon\eta^{1}+\varepsilon^{2}\eta^{2}+\varepsilon^{3}\eta^{3}+O(\varepsilon^{4}),$
where the coefficients are given in Prop. 2.10 and Prop. 2.2. A Taylor expansion leads to
$\eta^{*}(\zeta(x))$ $=\varepsilon\eta^{1}(x)+\varepsilon^{2}(\zeta^{1}\partial_{x}\eta^{1}(x)+\eta^{2}(x))$ (4.8)
$\quad+\varepsilon^{3}\left\{\zeta^{2}(x)\partial_{x}\eta^{1}(x)+\frac{1}{2}(\zeta^{1}(x))^{2}\partial^{2}_{x}\eta^{1}(x)+\zeta^{1}(x)\partial_{x}\eta^{2}(x)+\eta^{3}(x)\right\}+O(\varepsilon^{4}).$
Thereby we obtain
$\eta^{*}(\zeta(x))$ $=\varepsilon\cos x+\frac{1}{2}\varepsilon^{2}\big{[}2\eta_{2,0}-\zeta_{1,1}+(\zeta_{1,1}+2\eta_{2,2})\cos(2x)\big{]}+\frac{1}{8}\varepsilon^{3}\Big{[}\big{(}-\zeta_{1,1}^{2}-4\zeta_{2,2}-8\zeta_{1,1}\eta_{2,2}$
$\quad+8\eta_{3,1}\big{)}\cos x+\big{(}\zeta_{1,1}^{2}+4\zeta_{2,2}+8\zeta_{1,1}\eta_{2,2}+8\eta_{3,3}\big{)}\cos(3x)\Big{]}+O(\varepsilon^{4}).$ (4.9)
Taking the Fourier transform of this expression, we use the simple fact that $\widehat{\cos mx}(k)=\pi\delta(k-m)+\pi\delta(k+m)$. Inserting this expansion into the Fourier expansion of the Riemann stretch (2.25), we find that
$\partial_{x}X$ $=1+\varepsilon\frac{C_{h+z}}{C_{h}}\zeta_{1,1}\cos x+2\varepsilon^{2}\frac{C_{2(h+z)}}{C_{2h}}\zeta_{2,2}\cos(2x)+\varepsilon^{3}\biggr{[}\Big{(}\frac{d_{2}S_{z}}{C_{h}^{2}}\zeta_{1,1}+\frac{C_{h+z}}{C_{h}}\zeta_{3,1}\Big{)}\cos x$
$\quad+3\frac{C_{3(h+z)}}{C_{3h}}\zeta_{3,3}\cos(3x)\biggr{]}+O(\varepsilon^{4}),$
$\partial_{z}X$ $=\varepsilon\frac{S_{h+z}}{C_{h}}\zeta_{1,1}\sin x+2\varepsilon^{2}\frac{S_{2(h+z)}}{C_{2h}}\zeta_{2,2}\sin(2x)+\varepsilon^{3}\biggr{[}\Big{(}\frac{d_{2}C_{z}}{C_{h}^{2}}\zeta_{1,1}+\frac{S_{h+z}}{C_{h}}\zeta_{3,1}\Big{)}\sin x$
$\quad+3\frac{S_{3(h+z)}}{C_{3h}}\zeta_{3,3}\sin(3x)\biggr{]}+O(\varepsilon^{4}),$
where we have defined $C_{s}:=\cosh(s)\quad\textrm{and}\quad S_{s}:=\sinh(s).$ We have also used the expansion of $h_{\varepsilon}$ in (2.87) and substituted $\sinh(mh)=\cosh(mh)\tanh(mh)$ in the denominators,as well as expansions of $\tanh(mh)$ and standard trigonometric identities. It follows that the Jacobian is
$\mathcal{J}$ $=1+2\varepsilon\frac{C_{h+z}}{C_{h}}\zeta_{1,1}\cos x+\varepsilon^{2}\biggr{[}\frac{C_{2(h+z)}}{2C_{h}^{2}}\zeta_{1,1}^{2}+\Big{(}4\frac{C_{2(h+z)}}{C_{2h}}\zeta_{2,2}+\frac{1}{2C_{h}^{2}}\zeta_{1,1}^{2}\Big{)}\cos(2x)\biggr{]}$
$\quad+2\varepsilon^{3}\biggr{[}\Big{(}\frac{d_{2}S_{z}}{C_{h}^{2}}\zeta_{1,1}+\frac{C_{3(h+z)}}{C_{h}C_{2h}}\zeta_{1,1}\zeta_{2,2}+\frac{C_{h+z}}{C_{h}}\zeta_{3,1}\Big{)}\cos x+\Big{(}\frac{C_{h+z}}{C_{h}C_{2h}}\zeta_{1,1}\zeta_{2,2}+3\frac{C_{3(h+z)}}{C_{3h}}\zeta_{3,3}\Big{)}\cos(3x)\biggr{]}$
$\quad+O(\varepsilon^{4}).$
Given these various expansions, we return to $\Theta=\Theta_{f}$ which satisfies
$\begin{cases}\Delta_{x,z}\Theta-\beta\mathcal{J}\Theta=0\quad\text{in }\{(x,z):-h_{\varepsilon}<z<0\},\\ \Theta(x,0)=f(x),\quad\partial_{z}\Theta(x,-h_{\varepsilon})=0,\end{cases}$ (4.10)
as defined in (2.37). We recall from the proof of Theorem 2.6 that $\mathcal{G}_{\varepsilon,\beta}f=\partial_{z}\Theta(\cdot,0)$ and that $\Theta$ is analytic in $(\varepsilon,\beta)$ for each $h>0$. In particular, for fixed $\beta>0$ and $h>0$, we can expand $\Theta=\Theta^{0}+\varepsilon\Theta^{1}+\varepsilon^{2}\Theta^{2}+O(\varepsilon^{3})$, where $\Theta^{0}$, $\Theta^{1}$, $\Theta^{2}$, and $\Theta^{3}$ satisfy, by use of the expansion of the Jacobian, the equations
$\begin{cases}\Delta_{x,z}\Theta^{0}-\beta\Theta^{0}=0\quad\text{in }\{(x,z):-h<z<0\},\\ \Theta^{0}(x,0)=f(x),\\ \partial_{z}\Theta^{0}(x,-h)=0,\end{cases}$ (4.11)
$\begin{cases}\Delta_{x,z}\Theta^{1}-\beta\Theta^{1}=2\beta\frac{C_{h+z}}{C_{h}}\zeta_{1,1}(\cos x)\Theta^{0}\quad\text{in }\{(x,z):-h<z<0\},\\ \Theta^{1}(x,0)=0,\\ \partial_{z}\Theta^{1}(x,-h)=0,\end{cases}$ (4.12)
$&\Delta_{x,z}\Theta^{2}-\beta\Theta^{2}=\beta\biggr{[}\frac{C_{2(h+z)}}{2C_{h}^{2}}\zeta_{1,1}^{2}+\Big{(}4\frac{C_{2(h+z)}}{C_{2h}}\zeta_{2,2}+\frac{1}{2C_{h}^{2}}\zeta_{1,1}^{2}\Big{)}\cos(2x)\biggr{]}\Theta^{0}\\ &\hskip 88.2037pt+2\beta\frac{C_{h+z}}{C_{h}}\zeta_{1,1}\cos x\Theta^{1}\quad\text{in }\{(x,z):-h<z<0\},\\ &\Theta^{2}(x,0)=0,\\ &\partial_{z}\Theta^{2}(x,-h)=h_{2}\partial_{z}^{2}\Theta^{0}(x,-h),\end{aligned}\end{cases}$ (4.13)
$\Delta_{x,z}\Theta^{3}-\beta\Theta^{3}&=2\beta\biggr{[}\Big{(}\frac{d_{2}S_{z}}{C_{h}^{2}}\zeta_{1,1}+\frac{C_{3(h+z)}}{C_{h}C_{2h}}\zeta_{1,1}\zeta_{2,2}+\frac{C_{h+z}}{C_{h}}\zeta_{3,1}\Big{)}\cos x+\Big{(}\frac{C_{h+z}}{C_{h}C_{2h}}\zeta_{1,1}\zeta_{2,2}\\ &\hskip 24.18501pt+3\frac{C_{3(h+z)}}{C_{3h}}\zeta_{3,3}\Big{)}\cos(3x)\biggr{]}\Theta^{0}+\beta\biggr{[}\frac{C_{2(h+z)}}{2C_{h}^{2}}\zeta_{1,1}^{2}+\Big{(}4\frac{C_{2(h+z)}}{C_{2h}}\zeta_{2,2}\\ &\hskip 24.18501pt+\frac{1}{2C_{h}^{2}}\zeta_{1,1}^{2}\Big{)}\cos(2x)\biggr{]}\Theta^{1}+2\beta\frac{C_{h+z}}{C_{h}}\zeta_{1,1}\cos x\Theta^{2}\quad\text{in}\{(x,z):-h<z<0\},\end{aligned}\\ \Theta^{3}(x,0)=0,\\ \partial_{z}\Theta^{3}(x,-h)=h_{2}\partial_{z}^{2}\Theta^{1}(x,-h).\end{cases}$ (4.14)
Comparing with the expansion (4.1), we find that $R_{j,\beta}f=\partial_{z}\Theta^{j}(\cdot,0)$.
We consider the functions $\Theta^{j}$, one at a time. For $j=0$, we take the Fourier transform of (4.11) in $x$, leading to the new boundary value problem
$\begin{cases}\partial_{z}^{2}\widehat{\Theta^{0}}-(k^{2}+\beta)\widehat{\Theta^{0}}=0\quad\text{in }\{(k,z):-h<z<0\},\\ \widehat{\Theta^{0}}(x,0)=\widehat{f}(k),\\ \partial_{z}\widehat{\Theta^{0}}(k,-h)=0,\end{cases}$ (4.15)
For ease of notation, we define the functions
$\mathcal{C}_{i,j}(z):=\cosh\left\{z\left[i+\left((j+k)^{2}+\beta\right)^{\frac{1}{2}}\right]\right\},$ (4.16)
$\mathcal{S}_{i,j}(z):=\sinh\left\{z\left[i+\left((j+k)^{2}+\beta\right)^{\frac{1}{2}}\right]\right\}.$ (4.17)
The general solution of (4.15) may be expressed in terms of these functions as
$\widehat{\Theta^{0}}(k,z)=\mathcal{A}^{(0)}_{1}\mathcal{C}_{0,0}(z)+\mathcal{A}^{(0)}_{2}\mathcal{S}_{0,0}(z),$ (4.18)
where $\mathcal{A}^{(0)}_{1}$ and $\mathcal{A}^{(0)}_{2}$ are constants. Enforcing the requisite boundary conditions on $\widehat{\Theta^{0}}$, we find
$\mathcal{A}^{(0)}_{1}$ $=1,$ (4.19)
$\mathcal{A}^{(0)}_{2}$ $=\tanh\left(h(k^{2}+\beta)^{\frac{1}{2}}\right),$ (4.20)
so that
$\widehat{\Theta^{0}}(k,z)=\Big{[}\mathcal{C}_{0,0}(z)+\tanh\left(h(k^{2}+\beta)^{\frac{1}{2}}\right)\mathcal{S}_{0,0}(z)\Big{]}\widehat{f}(k).$ (4.21)
Consequently,
$\partial_{z}\widehat{\Theta^{0}}(k,0)=\widehat{R_{0,\beta}f}(k)=A^{0}_{k,\beta}\widehat{f}(k),\quad\text{where}\quad A^{0}_{k,\beta}:=(k^{2}+\beta)^{\frac{1}{2}}\tanh(h(k^{2}+\beta)^{\frac{1}{2}})$ (4.22)
with the notation (4.1). In other words, $R_{0,\beta}$ is the Fourier multiplier operator,
$R_{0,\beta}f=(|D|^{2}+\beta)^{\frac{1}{2}}\tanh(h(|D|^{2}+\beta)^{\frac{1}{2}})f,\quad D=-i\partial_{x},$ (4.23)
which is equivalent to $\mathcal{G}_{0,\beta}f$ given in (2.34).
Remark 4.2.
If $\beta=0$, we find $R_{0,0}=|D|\tanh(h|D|)$, in agreement with [6]. If instead $\beta\neq 0$ and we let $h\rightarrow\infty$, we find $R_{0,\beta}\rightarrow(|D|^{2}+\beta)^{\frac{1}{2}}$, in agreement with [17].
We can follow a similar series of calculations to determine $R_{1,\beta}$. The Fourier transform in $x$ of the equation for $\Theta^{1}$ given in (4.12) is
$\partial_{z}^{2}\widehat{\Theta^{1}}-(k^{2}+\beta)\widehat{\Theta^{1}}$ $=\beta\frac{C_{h+z}}{C_{h}}\zeta_{1,1}\Big{[}\widehat{\Theta^{0}}(-1+k,z)+\widehat{\Theta^{0}}(1+k,z)\Big{]}.$ (4.24)
Substituting (4.21) into the right side of (4.24) and expanding, the equation for $\Theta^{1}$ simplifies to
$\partial_{z}^{2}\widehat{\Theta^{1}}-(k^{2}+\beta)\widehat{\Theta^{1}}$ $=\sum_{i,j\in\{-1,1\}}\Big{(}\mathfrak{c}^{(1)}_{i,j}\mathcal{C}_{i,j}(z)+\mathfrak{s}^{(1)}_{i,j}\mathcal{S}_{i,j}(z)\Big{)},$ (4.25)
where
$\mathfrak{c}^{(1)}_{-1,-1}$ $=-\frac{1}{2}\beta\zeta_{1,1}\Big{\{}-1+\tanh(h)\tanh\left[h((-1+k)^{2}+\beta)^{\frac{1}{2}}\right]\Big{\}}\widehat{f}(-1+k),$ (4.26)
$\mathfrak{c}^{(1)}_{1,-1}$ $=\frac{1}{2}\beta\zeta_{1,1}\Big{\{}1+\tanh(h)\tanh\left[h((-1+k)^{2}+\beta)^{\frac{1}{2}}\right]\Big{\}}\widehat{f}(-1+k),$ (4.27)
$\mathfrak{c}^{(1)}_{-1,1}$ $=-\frac{1}{2}\beta\zeta_{1,1}\Big{\{}-1+\tanh(h)\tanh\left[h((1+k)^{2}+\beta)^{\frac{1}{2}}\right]\Big{\}}\widehat{f}(1+k),$ (4.28)
$\mathfrak{c}^{(1)}_{1,1}$ $=\frac{1}{2}\beta\zeta_{1,1}\Big{\{}1+\tanh(h)\tanh\left[h((1+k)^{2}+\beta)^{\frac{1}{2}}\right]\Big{\}}\widehat{f}(1+k),$ (4.29)
$\mathfrak{s}^{(1)}_{-1,-1}$ $=\frac{1}{2}\beta\zeta_{1,1}\Big{\{}-\tanh(h)+\tanh\left[h((-1+k)^{2}+\beta)^{\frac{1}{2}}\right]\Big{\}}\widehat{f}(-1+k),$ (4.30)
$\mathfrak{s}^{(1)}_{1,-1}$ $=\frac{1}{2}\beta\zeta_{1,1}\Big{\{}\tanh(h)+\tanh\left[h((-1+k)^{2}+\beta)^{\frac{1}{2}}\right]\Big{\}}\widehat{f}(-1+k),$ (4.31)
$\mathfrak{s}^{(1)}_{-1,1}$ $=\frac{1}{2}\beta\zeta_{1,1}\Big{\{}-\tanh(h)+\tanh\left[h((1+k)^{2}+\beta)^{\frac{1}{2}}\right]\Big{\}}\widehat{f}(1+k),$ (4.32)
$\mathfrak{s}^{(1)}_{1,1}$ $=\frac{1}{2}\beta\zeta_{1,1}\Big{\{}\tanh(h)+\tanh\left[h((1+k)^{2}+\beta)^{\frac{1}{2}}\right]\Big{\}}\widehat{f}(1+k).$ (4.33)
The general solution of (4.24) is
$\widehat{\Theta^{1}}(k,z)=\mathcal{A}^{(1)}_{1}\mathcal{C}_{0,0}(z)+\mathcal{A}^{(1)}_{2}\mathcal{S}_{0,0}(z)+\sum_{i,j\in\{-1,1\}}\Big{(}\tilde{\mathfrak{c}}^{(1)}_{i,j}\mathcal{C}_{i,j}(z)+\tilde{\mathfrak{s}}^{(1)}_{i,j}\mathcal{S}_{i,j}(z)\Big{)},$ (4.34)
where $\mathcal{A}^{(1)}_{1}$ and $\mathcal{A}^{(1)}_{2}$ are arbitrary constants and
$\tilde{\mathfrak{c}}^{(1)}_{i,j}=\frac{\mathfrak{c}^{(1)}_{i,j}}{\Big{[}i+\left((j+k)^{2}+\beta\right)^{\frac{1}{2}}\Big{]}^{2}-(k^{2}+\beta)},$ (4.35)
$\tilde{\mathfrak{s}}^{(1)}_{i,j}=\frac{\mathfrak{s}^{(1)}_{i,j}}{\Big{[}i+\left((j+k)^{2}+\beta\right)^{\frac{1}{2}}\Big{]}^{2}-(k^{2}+\beta)}.$ (4.36)
Enforcing the requisite boundary conditions on $\widehat{\Theta^{1}}$, we determine
$\mathcal{A}^{(1)}_{1}$ $=-\sum_{i,j\in\{-1,1\}}\tilde{\mathfrak{c}}^{(1)}_{i,j},$ (4.37)
$\mathcal{A}^{(1)}_{2}$ $=-\sum_{i,j\in\{-1,1\}}\biggr{\{}\Big{[}\frac{i+\left((j+k)^{2}+\beta\right)^{\frac{1}{2}}}{\mathcal{C}_{0,0}(h)(k^{2}+\beta)^{\frac{1}{2}}}\Big{]}\Big{[}-\tilde{\mathfrak{c}}^{(1)}_{i,j}\mathcal{S}_{i,j}(h)+\tilde{\mathfrak{s}}^{(1)}_{i,j}\mathcal{C}_{i,j}(h)\Big{]}+\tilde{\mathfrak{c}}^{(1)}_{i,j}\tanh\left(h\left(k^{2}+\beta\right)^{\frac{1}{2}}\right)\biggr{\}}.$ (4.38)
A direct calculation from (4.11) yields
$\partial_{z}\widehat{\Theta^{1}}(0,z)=\mathcal{A}_{2}^{(1)}(k^{2}+\beta)^{\frac{1}{2}}+\sum_{i,j\in\{-1,1\}}\tilde{\mathfrak{s}}^{(1)}_{i,j}\left(i+\left((j+k)^{2}+\beta\right)^{\frac{1}{2}}\right).$ (4.39)
Substituting $\mathcal{A}^{(1)}_{2}$, $\tilde{\mathfrak{s}}^{(1)}_{i,j}$, and $\tilde{\mathfrak{c}}^{(1)}_{i,j}$ into (4.39) and expanding, we find that
$\partial_{z}\widehat{\Theta^{1}}(0,z)=B_{k,\beta}^{-1}\widehat{f}(-1+k)+B_{k,\beta}^{1}\widehat{f}(1+k),$ (4.40)
where
$B_{k,\beta}^{-1}$ $:=\frac{1}{2}\biggr{\{}\beta-\sqrt{\beta+(k-1)^{2}}\sqrt{\beta+k^{2}}\tanh\left(h\sqrt{\beta+(k-1)^{2}}\right)\tanh\left(h\sqrt{\beta+k^{2}}\right)$
$\quad+\coth(h)\left[k\sqrt{\beta+(k-1)^{2}}\tanh\left(h\sqrt{\beta+(k-1)^{2}}\right)-(k-1)\sqrt{\beta+k^{2}}\tanh\left(h\sqrt{\beta+k^{2}}\right)\right]$
$\quad+k^{2}-k\biggr{\}},$ (4.41)
$B_{k,\beta}^{1}$ $:=\frac{1}{2}\biggr{\{}\beta-\sqrt{\beta+k^{2}}\sqrt{\beta+(k+1)^{2}}\tanh\left(h\sqrt{\beta+k^{2}}\right)\tanh\left(h\sqrt{\beta+(k+1)^{2}}\right)$
$\quad+\coth(h)\left[(k+1)\sqrt{\beta+k^{2}}\tanh\left(h\sqrt{\beta+k^{2}}\right)-k\sqrt{\beta+(k+1)^{2}}\tanh\left(h\sqrt{\beta+(k+1)^{2}}\right)\right]$
$\quad+k^{2}+k\biggr{\}}.$ (4.42)
It follows that $R_{1,\beta}$ is the operator given by
$R_{1,\beta}=B_{k,\beta}^{-1}\widehat{f}(-1+k)+B_{k,\beta}^{1}\widehat{f}(1+k).$ (4.43)
Remark 4.3.
Because $\mathcal{G}_{\varepsilon,\beta}$ is self-adjoint, we have $B^{-1}_{-k,\beta}=B^{1}_{k,\beta}$ for all $k\in\mathbb{R}$, in agreement with (4.41) and (4.42). Moreover, we easily verify the following limits hold at infinite depth:
$\lim_{h\rightarrow+\infty}B^{-1}_{k,\beta}$ $=\frac{1}{2}\left(\beta-(k-1)\sqrt{\beta+k^{2}}-\sqrt{\left(\beta+(k-1)^{2}\right)\left(\beta+k^{2}\right)}+k^{2}+k\sqrt{\beta+(k-1)^{2}}-k\right),$
$\lim_{h\rightarrow+\infty}B^{1}_{k,\beta}$ $=\frac{1}{2}\left(\beta+(k+1)\sqrt{\beta+k^{2}}-\sqrt{\left(\beta+k^{2}\right)\left(\beta+(k+1)^{2}\right)}+k^{2}-k\sqrt{\beta+(k+1)^{2}}+k\right),$
which are consistent with the coefficients obtained in [17].
The calculations leading to $R_{2,\beta}$ and $R_{3,\beta}$ follow the same procedure as those for $R_{0,\beta}$ and $R_{1,\beta}$ shown above. However, the expressions involved in these calculations become extremely lengthy and arduous to work with by hand. We instead appeal to Mathematica to help us finish these calculations. The interested reader will find the explicit expressions for the coefficients of $R_{2,\beta}$ and $R_{3,\beta}$ in our companion Mathematica file CompanionToTransverseInstabilityFiniteDepth.m. ∎
Next, we expand the operators $\mathcal{H}^{j}_{\beta}$ in (4.2) to third order in $\delta=\beta-\beta_{*}$. By virtue of Theorem 2.6 (ii), for each $h>0$, the operators $R_{j,\beta}$ depend analytically on $\beta\in(0,\infty)$. In particular, we have
$R_{j,\beta}=\sum_{\ell=0}^{3}\delta^{\ell}R^{j,\ell}+O(\delta^{4}),$ (4.44)
where of course the $R^{j,\ell}$ depend on $\beta_{*}$ and $h$. Noting that $\beta$ only appears in the lower right corner of $\mathcal{H}^{j}_{\beta}$, we set
$\mathcal{H}^{j,0}=\mathcal{H}^{j}_{\beta}|_{\delta=0},\quad\mathcal{H}^{j,\ell}=R^{j,\ell}\begin{bmatrix}0&0\\ 0&1\end{bmatrix},\quad\ell\in\{1,2,3\},$ (4.45)
so that
$\mathcal{H}^{j}_{\beta}=\sum_{\ell=0}^{3}\delta^{\ell}\mathcal{H}^{j,\ell}+O(\delta^{4}).$ (4.46)
Combining (LABEL:Hexpansion:eps) and (4.46), we obtain the third-order expansion of the Hamiltonian $\mathcal{H}_{\varepsilon,\beta_{*}+\delta}$ in both $\varepsilon$ and $\delta$:
$\mathcal{H}_{\varepsilon,\beta_{*}+\delta}=\sum_{j=0}^{3}\sum_{\ell=0}^{3}\varepsilon^{j}\delta^{\ell}\mathcal{H}^{j,\ell}+O(\varepsilon^{4}+\delta^{4}).$ (4.47)
4.2. Third-order expansions of $\textrm{L}_{\varepsilon,\delta}$
We recall the basis $\{U_{1},U_{2}\}$ from (3.13) and the projections $P_{\varepsilon,\delta}$ from (3.17). Since $P_{0,0}U_{j}=U_{j}$ for $j=1,2$, we have
$U^{\varepsilon,\delta}_{j}$ =$=\{1-(P_{\varepsilon,\delta}-P_{0,0})^{2}\}^{-\frac{1}{2}}P_{\varepsilon,\delta}U_{j}.$
Next, we record the following third-order expansions of $U^{\varepsilon,\delta}_{j}$ and $\textrm{L}_{\varepsilon,\delta}$. We refer to [17] for the derivation of these formulas. Denoting
$P^{m,n}=\partial_{\varepsilon}^{m}\partial_{\delta}^{n}P_{\varepsilon,\delta}|_{(\varepsilon,\delta)=(0,0)},$ (4.48)
we have
$U^{\varepsilon,\delta}_{j}=U_{j}+\sum_{m+n=1}^{3}\varepsilon^{m}\delta^{n}U^{(m,n)}_{j}+O((\varepsilon+\delta)^{4}),$ (4.49)
where the coefficients $U_{j}^{(m,n)}$ are given by
$U^{(1,0)}_{j}=P^{1,0}U_{j},\quad U^{(0,1)}_{j}=P^{0,1}U_{j},$ (4.50)
$U^{(2,0)}_{j}=\frac{1}{2}(P^{2,0}+P^{1,0}P^{1,0})U_{j},\quad U^{(0,2)}_{j}=\frac{1}{2}(P^{0,2}+P^{0,1}P^{0,1})U_{j},$ (4.51)
$U^{(1,1)}_{j}=(P^{1,1}+\frac{1}{2}P^{0,1}P^{1,0}+\frac{1}{2}P^{1,0}P^{0,1})U_{j},$ (4.52)
$U_{j}^{(3,0)}=\frac{1}{6}\Big{(}P^{3,0}+\frac{3}{2}\left(P^{2,0}P^{1,0}+P^{1,0}P^{2,0}\right)+3P^{1,0}P^{1,0}P^{1,0}\Big{)}U_{j},$ (4.53)
$U_{j}^{(0,3)}=\frac{1}{6}\Big{(}P^{0,3}+\frac{3}{2}\left(P^{0,1}P^{0,2}+P^{0,2}P^{0,1}\right)+3P^{0,1}P^{0,1}P^{0,1}\Big{)}U_{j},$ (4.54)
$U_{j}^{(2,1)}=\frac{1}{2}\Big{(}P^{2,1}+\frac{1}{2}\left(P^{0,1}P^{2,0}+2P^{1,0}P^{1,1}+2P^{1,1}P^{1,0}+P^{2,0}P^{0,1}\right)$
$\hskip 28.45274pt+\left(P^{1,0}P^{0,1}+P^{0,1}P^{1,0}\right)P^{1,0}+P^{1,0}P^{1,0}P^{0,1}\Big{)}U_{j},$ (4.55)
$U_{j}^{(1,2)}=\frac{1}{2}\Big{(}P^{1,2}+\frac{1}{2}\left(P^{1,0}P^{0,2}+2P^{1,1}P^{0,1}+2P^{0,1}P^{1,1}+P^{0,2}P^{1,0}\right)$
$\hskip 28.45274pt+\left(P^{1,0}P^{0,1}+P^{0,1}P^{1,0}\right)P^{0,1}+P^{0,1}P^{0,1}P^{1,0}\Big{)}U_{j}.$ (4.56)
Here $P^{m,n}U_{j}$ are computed from the following contour integrals:
$P^{1,0}U_{j}$ $=\frac{1}{2\pi i}\int_{\Gamma}-S_{\lambda}J\mathcal{H}^{1,0}U_{j}\frac{d\lambda}{\lambda-i\sigma},$ (4.57)
$P^{0,1}U_{j}$ $=\frac{1}{2\pi i}\int_{\Gamma}-S_{\lambda}J\mathcal{H}^{0,1}U_{j}\frac{d\lambda}{\lambda-i\sigma},$ (4.58)
$P^{2,0}U_{j}$ $=\frac{1}{2\pi i}\int_{\Gamma}S_{\lambda}\Big{(}-2J\mathcal{H}^{2,0}+2J\mathcal{H}^{1,0}J\mathcal{H}^{1,0}\Big{)}U_{j}\frac{d\lambda}{\lambda-i\sigma},$ (4.59)
$P^{0,2}U_{j}$ $=\frac{1}{2\pi i}\int_{\Gamma}S_{\lambda}\Big{(}-2J\mathcal{H}^{0,2}+2J\mathcal{H}^{0,1}S_{\lambda}J\mathcal{H}^{0,1}\Big{)}U_{j}\frac{d\lambda}{\lambda-i\sigma},$ (4.60)
$P^{1,1}U_{j}$ $=\frac{1}{2\pi i}\int_{\Gamma}S_{\lambda}\Big{(}-J\mathcal{H}^{1,1}+J\mathcal{H}^{1,0}S_{\lambda}J\mathcal{H}^{0,1}+J\mathcal{H}^{0,1}S_{\lambda}J\mathcal{H}^{1,0}\Big{)}U_{j}\frac{d\lambda}{\lambda-i\sigma},$ (4.61)
$P^{3,0}U_{j}$ $=\frac{1}{2\pi i}\int_{\Gamma}6S_{\lambda}\Big{[}-JH^{3,0}+\big{(}JH^{1,0}S_{\lambda}JH^{2,0}+JH^{2,0}S_{\lambda}JH^{1,0}\big{)}$
$\hskip 28.45274pt-JH^{1,0}S_{\lambda}JH^{1,0}S_{\lambda}JH^{1,0}\Big{]}U_{j}\frac{d\lambda}{\lambda-i\sigma},$ (4.62)
$P^{0,3}U_{j}$ $=\frac{1}{2\pi i}\int_{\Gamma}6S_{\lambda}\Big{[}-JH^{0,3}+\big{(}JH^{0,1}S_{\lambda}JH^{0,2}+JH^{0,2}S_{\lambda}JH^{0,1}\big{)}$
$\hskip 28.45274pt-JH^{0,1}S_{\lambda}JH^{0,1}S_{\lambda}JH^{0,1}\Big{]}U_{j}\frac{d\lambda}{\lambda-i\sigma},$ (4.63)
$P^{2,1}U_{j}$ $=\frac{1}{2\pi i}\int_{\Gamma}2S_{\lambda}\Big{[}-JH^{2,1}+JH^{1,0}S_{\lambda}JH^{1,1}+JH^{0,1}S_{\lambda}JH^{2,0}$
$\hskip 28.45274pt+JH^{2,0}S_{\lambda}JH^{0,1}+JH^{1,1}S_{\lambda}JH^{1,0}$
$\hskip 28.45274pt-JH^{1,0}S_{\lambda}\big{(}JH^{1,0}S_{\lambda}JH^{0,1}+JH^{0,1}S_{\lambda}JH^{1,0}\big{)}\Big{]}U_{j}\frac{d\lambda}{\lambda-i\sigma},$ (4.64)
$P^{1,2}U_{j}$ $=\frac{1}{2\pi i}\int_{\Gamma}2S_{\lambda}\Big{[}-JH^{1,2}+JH^{0,1}S_{\lambda}JH^{1,1}+JH^{1,0}S_{\lambda}JH^{0,2}$
$\hskip 28.45274pt+JH^{0,2}S_{\lambda}JH^{1,0}+JH^{1,1}S_{\lambda}JH^{0,1}$
$\hskip 28.45274pt-JH^{0,1}S_{\lambda}\big{(}JH^{0,1}S_{\lambda}JH^{1,0}+JH^{0,1}S_{\lambda}JH^{0,1}\big{)}\Big{]}U_{j}\frac{d\lambda}{\lambda-i\sigma}.$ (4.65)
We recall that $\textrm{L}_{\varepsilon,\delta}$ given by (3.23) is the matrix representation of the linearized operator $\mathcal{L}_{\varepsilon,\beta_{*}+\delta}$ with respect to the basis $\{V_{j}^{\varepsilon,\delta}:j=1,2\}$, where $V_{j}^{\varepsilon,\delta}=\frac{1}{\sqrt{\gamma_{j}}}U_{j}^{\varepsilon,\delta}$. It follows from the expansion (4.49) of $U_{j}^{\varepsilon,\delta}$ that
$V^{\varepsilon,\delta}_{j}=V_{j}+\sum_{m+n=1}^{3}\varepsilon^{m}\delta^{n}V_{j}^{(m,n)}+O((\varepsilon+\delta)^{4}),\quad V_{j}=\frac{1}{\sqrt{\gamma_{j}}}U_{j},\quad V_{j}^{(m,n)}=\frac{1}{\sqrt{\gamma_{j}}}U_{j}^{(m,n)}.$ (4.66)
Combining (4.66) with the expansion (4.47) for $\mathcal{H}_{\varepsilon,\beta_{*}+\delta}$, we expand the inner products appearing in $L_{\varepsilon,\delta}$ (3.23) as
$\left(\mathcal{H}_{\varepsilon,\beta_{*}+\delta}V_{j}^{\varepsilon,\delta},V_{k}^{\varepsilon,\delta}\right)=(\mathcal{H}^{0,0}V_{j},V_{k})+\sum_{m+n=1}^{3}\left(\mathcal{H}_{\varepsilon,\beta_{*}+\delta}V_{j}^{\varepsilon,\delta},V_{k}^{\varepsilon,\delta}\right)_{m,n}\varepsilon^{m}\delta^{n}+O((\varepsilon+\delta)^{4}),$ (4.67)
where
$\left(\mathcal{H}_{\varepsilon,\beta_{*}+\delta}V_{j}^{\varepsilon,\delta},V_{k}^{\varepsilon,\delta}\right)_{1,0}$ $=(\mathcal{H}^{0,0}V_{j},V^{(1,0)}_{k})+(\mathcal{H}^{1,0}V_{j},V_{k})+(\mathcal{H}^{0,0}V_{j}^{(1,0)},V_{k}),$ (4.68)
$\left(\mathcal{H}_{\varepsilon,\beta_{*}+\delta}V_{i}^{\varepsilon,\delta},V_{j}^{\varepsilon,\delta}\right)_{0,1}$ $=(\mathcal{H}^{0,0}V_{j},V^{(0,1)}_{k})+(\mathcal{H}^{0,1}V_{j},V_{k})+(\mathcal{H}^{0,0}V_{j}^{(0,1)},V_{k}),$ (4.69)
$\left(\mathcal{H}_{\varepsilon,\beta_{*}+\delta}V_{i}^{\varepsilon,\delta},V_{j}^{\varepsilon,\delta}\right)_{2,0}$ $=(\mathcal{H}^{0,0}V_{j},V^{(2,0)}_{k})+(\mathcal{H}^{2,0}V_{j},V_{k})+(\mathcal{H}^{1,0}V^{(1,0)}_{j},V_{k})$
$\hskip 28.45274pt+(\mathcal{H}^{0,0}V^{(2,0)}_{j},V_{k})+(\mathcal{H}^{1,0}V_{j},V^{(1,0)}_{k})+(\mathcal{H}^{0,0}V_{j}^{(1,0)},V_{k}^{(1,0)}),$ (4.70)
$\left(\mathcal{H}_{\varepsilon,\beta_{*}+\delta}V_{i}^{\varepsilon,\delta},V_{j}^{\varepsilon,\delta}\right)_{0,2}$ $=(\mathcal{H}^{0,0}V_{j},V^{(0,2)}_{k})+(\mathcal{H}^{0,2}V_{j},V_{k})+(\mathcal{H}^{0,1}V^{(0,1)}_{j},V_{k})$
$\hskip 28.45274pt+(\mathcal{H}^{0,0}V^{(0,2)}_{j},V_{k})+(\mathcal{H}^{0,1}V_{j},V^{(0,1)}_{k})+(\mathcal{H}^{0,0}V_{j}^{(0,1)},V_{k}^{(0,1)}),$ (4.71)
$\left(\mathcal{H}_{\varepsilon,\beta_{*}+\delta}V_{i}^{\varepsilon,\delta},V_{j}^{\varepsilon,\delta}\right)_{1,1}$ $=(\mathcal{H}^{1,1}V_{j},V_{k})+(\mathcal{H}^{1,0}V_{j}^{(0,1)},V_{k})+(\mathcal{H}^{0,1}V_{j}^{(1,0)},V_{k})$
$\hskip 28.45274pt+(\mathcal{H}^{0,0}V_{j},V_{k}^{(1,1)})+(\mathcal{H}^{0,0}V_{j}^{(1,1)},V_{k})+(\mathcal{H}^{1,0}V_{j},V_{k}^{(0,1)})$
$\quad+(\mathcal{H}^{0,0}V_{j}^{(1,0)},V_{k}^{(0,1)})+(\mathcal{H}^{0,1}V_{j},V_{k}^{(1,0)})+(\mathcal{H}^{0,0}V^{(0,1)}_{j},V_{k}^{(1,0)}),$ (4.72)
$\left(\mathcal{H}_{\varepsilon,\beta_{*}+\delta}V^{\varepsilon,\delta}_{j},V^{\varepsilon,\delta}_{k}\right)_{3,0}$ $=\Big{(}\mathcal{H}^{3,0}V_{j}+\mathcal{H}^{2,0}V_{j}^{(1,0)}+\mathcal{H}^{1,0}V_{j}^{(2,0)}+\mathcal{H}^{0,0}V_{j}^{(3,0)},V_{k}\Big{)}$
$\hskip 28.45274pt+\Big{(}\mathcal{H}^{2,0}V_{j}+\mathcal{H}^{1,0}V_{j}^{(1,0)}+\mathcal{H}^{0,0}V_{j}^{(2,0)},V_{k}^{(1,0)}\Big{)}$
$\hskip 28.45274pt+\Big{(}\mathcal{H}^{1,0}V_{j}+\mathcal{H}^{0,0}V_{j}^{(1,0)},V_{k}^{(2,0)}\Big{)}+\Big{(}\mathcal{H}^{0,0}V_{j},V_{k}^{(3,0)}\Big{)},$ (4.73)
$\left(\mathcal{H}_{\varepsilon,\beta_{*}+\delta}V^{\varepsilon,\delta}_{j},V^{\varepsilon,\delta}_{k}\right)_{0,3}$ $=\Big{(}\mathcal{H}^{0,3}V_{j}+\mathcal{H}^{0,2}V_{j}^{(0,1)}+\mathcal{H}^{0,1}V_{j}^{(0,2)}+\mathcal{H}^{0,0}V_{j}^{(0,3)},V_{k}\Big{)}$
$\hskip 28.45274pt+\Big{(}\mathcal{H}^{0,2}V_{j}+\mathcal{H}^{0,1}V_{j}^{(0,1)}+\mathcal{H}^{0,0}V_{j}^{(0,2)},V_{k}^{(0,1)}\Big{)}$
$\hskip 28.45274pt+\Big{(}\mathcal{H}^{0,1}V_{j}+\mathcal{H}^{0,0}V_{j}^{(0,1)},V_{k}^{(0,2)}\Big{)}+\Big{(}\mathcal{H}^{0,0}V_{j},V_{k}^{(0,3)}\Big{)},$ (4.74)
$\left(\mathcal{H}_{\varepsilon,\beta_{*}+\delta}V^{\varepsilon,\delta}_{j},V^{\varepsilon,\delta}_{k}\right)_{2,1}$ $=\Big{(}\mathcal{H}^{2,1}V_{j}+\mathcal{H}^{2,0}V_{j}^{(0,1)}+\mathcal{H}^{1,1}V_{j}^{(1,0)}+\mathcal{H}^{1,0}V_{j}^{(1,1)}+\mathcal{H}^{0,1}V_{j}^{(2,0)}+\mathcal{H}^{0,0}V_{j}^{(2,1)},V_{k}\Big{)}$
$\hskip 28.45274pt+\Big{(}\mathcal{H}^{2,0}V_{j}+\mathcal{H}^{1,0}V_{j}^{(1,0)}+\mathcal{H}^{0,0}V_{j}^{(2,0)},V_{k}^{(0,1)}\Big{)}$
$\hskip 28.45274pt+\Big{(}\mathcal{H}^{1,1}V_{j}+\mathcal{H}^{1,0}V_{j}^{(0,1)}+\mathcal{H}^{0,1}V_{j}^{(1,0)}+\mathcal{H}^{0,0}V_{j}^{(1,1)},V_{n}^{(1,0)}\Big{)}$
$\hskip 28.45274pt+\Big{(}\mathcal{H}^{1,0}V_{j}+\mathcal{H}^{0,0}V_{j}^{(1,0)},V_{k}^{(1,1)}\Big{)}+\Big{(}\mathcal{H}^{0,1}V_{j}+\mathcal{H}^{0,0}V_{j}^{(0,1)},V_{k}^{(2,0)}\Big{)}$
$\hskip 28.45274pt+\Big{(}\mathcal{H}^{0,0}V_{j},V_{k}^{(2,1)}\Big{)},$ (4.75)
$\left(\mathcal{H}_{\varepsilon,\beta_{*}+\delta}V^{\varepsilon,\delta}_{j},V^{\varepsilon,\delta}_{k}\right)_{1,2}$ $=\Big{(}\mathcal{H}^{1,2}V_{j}+\mathcal{H}^{1,1}V_{j}^{(0,1)}+\mathcal{H}^{1,0}V_{j}^{(0,2)}+\mathcal{H}^{0,2}V_{j}^{(1,0)}+\mathcal{H}^{0,1}V_{j}^{(1,1)}+\mathcal{H}^{0,0}V_{j}^{(1,2)},V_{k}\Big{)}$
$\hskip 28.45274pt+\Big{(}\mathcal{H}^{0,2}V_{j}+\mathcal{H}^{0,1}V_{j}^{(0,1)}+\mathcal{H}^{0,0}V_{j}^{(0,2)},V_{k}^{(1,0)}\Big{)}$
$\hskip 28.45274pt+\Big{(}\mathcal{H}^{1,1}V_{j}+\mathcal{H}^{1,0}V_{j}^{(0,1)}+\mathcal{H}^{0,1}V_{j}^{(1,0)}+\mathcal{H}^{0,0}V_{j}^{(1,1)},V_{n}^{(0,1)}\Big{)}$
$\hskip 28.45274pt+\Big{(}\mathcal{H}^{0,1}V_{j}+\mathcal{H}^{0,0}V_{j}^{(0,1)},V_{k}^{(1,1)}\Big{)}+\Big{(}\mathcal{H}^{1,0}V_{j}+\mathcal{H}^{0,0}V_{j}^{(1,0)},V_{k}^{(0,2)}\Big{)}$
$\hskip 28.45274pt+\Big{(}\mathcal{H}^{0,0}V_{j},V_{k}^{(1,2)}\Big{)}.$ (4.76)
5. Proof of Theorem 1.1
Lemma 5.1.
The purely imaginary matrix $\textrm{L}_{\varepsilon,\delta}$ can be written as
$\textrm{L}_{\varepsilon,\delta}=\begin{pmatrix}i\sigma&0\\ 0&i\sigma\end{pmatrix}+i\begin{pmatrix}A&B\\ -B&C\end{pmatrix},$ (5.1)
where $i\sigma$ is the repeated eigenvalue of the unperturbed operator $\mathcal{L}_{0,\beta_{*}}$ and where $A$, $B$, and $C$ are real analytic functions of $(\varepsilon,\delta)$ in a neighborhood of $(0,0)$ that have the expansions
$A$ $=a_{0,1}\delta+a_{2,0}\varepsilon^{2}+a_{0,2}\delta^{2}+a_{2,1}\varepsilon^{2}\delta+a_{0,3}\delta^{3}+a_{4,0}\varepsilon^{4}+O(\delta(|\varepsilon|^{3}+|\delta|^{3})),$ (5.2)
$B$ $=b_{3,0}\varepsilon^{3}+O(|\varepsilon|^{4}+|\delta|^{4}),$
$C$ $=c_{0,1}\delta+c_{2,0}\varepsilon^{2}+c_{0,2}\delta^{2}+c_{2,1}\varepsilon^{2}\delta+c_{0,3}\delta^{3}+c_{4,0}\varepsilon^{4}+O(\delta(|\varepsilon|^{3}+|\delta|^{3})).$
We emphasize that all these coefficients depend analytically on the depth $h\in(0,\infty)$.
Proof.
The proof is identical to that in the infinite depth case [17]. In particular, since $\textrm{L}_{\varepsilon,\delta}$ is purely imaginary and real analytic in $\varepsilon$ and $\delta$, we have that $A$, $B$, and $C$ are real-valued, real-analytic functions of $\varepsilon$ and $\delta$. The anti-symmetry of the off-diagonal entries of $\textrm{L}_{\varepsilon,\delta}$ follows from (3.24). Expansions (5.2) come from substituting the expansions (4.68)-(4.2) of the inner products $(\mathcal{H}_{\varepsilon,\beta_{*}+\delta}V_{j}^{\varepsilon,\delta},V_{k}^{\varepsilon,\delta})$ into the matrix $L_{\varepsilon,\delta}$ (3.23). As in the infinite depth case, each term in the expansion of $V_{j}^{\varepsilon,\delta}$ consists of a finite Fourier series spanning a small set of wave numbers given in the table below. Terms in the expansion of $\mathcal{H}_{\varepsilon,\beta_{*}+\delta}$ act to modulate these wave numbers in such a way that many terms in the expansion of $(\mathcal{H}_{\varepsilon,\beta_{*}+\delta}V_{j}^{\varepsilon,\delta},V_{k}^{\varepsilon,\delta})$ vanish. Those that do not vanish yield the coefficients $a_{i,j}$, $b_{i,j}$, and $c_{i,j}$ that appear above. These coefficients are real analytic functions of the resonant value $\beta_{*}$ and the depth parameter $h$. Their explicit expressions, except for $a_{4,0}$ and $c_{4,0}$, can be found in our Mathematica file CompanionToTransverseInstabilityFiniteDepth.nb. It will turn out that the fourth-order coefficients $a_{4,0}$ and $c_{4,0}$ are not needed for the proof of Theorem 1.1.
$V_{1}^{(k,\ell)}$ Wave Numbers $V_{2}^{(k,\ell)}$ Wave Numbers
$V_{1}$$\{1\}$ $V_{2}$$\{-2\}$
$V_{1}^{(1,0)}$$\{0,2\}$ $V_{2}^{(1,0)}$$\{-3,-1\}$
$V_{1}^{(0,1)}$$\{1\}$ $V_{2}^{(0,1)}$$\{-2\}$
$V_{1}^{(2,0)}$$\{-1,1,3\}$ $V_{2}^{(2,0)}$$\{-4,-2,0\}$
$V_{1}^{(1,1)}$$\{0,2\}$ $V_{2}^{(1,1)}$$\{-3,-1\}$
$V_{1}^{(0,2)}$$\{1\}$ $V_{2}^{(0,2)}$$\{-2\}$
$V_{1}^{(3,0)}$$\{-2,0,2,4\}$ $V_{2}^{(3,0)}$$\{-5,-3,-1,1\}$
$V_{1}^{(2,1)}$$\{-1,1,3\}$ $V_{2}^{(2,1)}$$\{-4,-2,0\}$
$V_{1}^{(1,2)}$$\{0,2\}$ $V_{2}^{(1,2)}$$\{-3,-1\}$
$V_{1}^{(0,3)}$$\{1\}$ $V_{2}^{(0,3)}$$\{-2\}$
As it turns out, most of the coefficients $a_{i,j}$, $b_{i,j}$, and $c_{i,j}$ in expansions (5.2) do not play a central role in the stability calculations that follow. However, we will need both that $a_{0,1}\neq c_{0,1}$ and that $b_{3,0}$ is not identically zero for $h>0$, which we prove in the following lemmata.
Lemma 5.2.
For all $h>0$, $a_{0,1}<0<c_{0,1}$.
Proof.
From CompanionToTransverseInstabilityFiniteDepth.nb, we have
$a_{0,1}=-\frac{\tau_{1}}{2\gamma_{1}}\quad\textrm{and}\quad c_{0,1}=\frac{\tau_{2}}{2\gamma_{2}},$
where $\gamma_{j}>0$ is given in (3.15) and
$\tau_{j}:=\frac{1}{2}\left(h\ \textrm{sech}^{2}\left(h\left(j^{2}+\beta_{*}\right)^{\frac{1}{2}}\right)+\frac{\tanh(h\left(j^{2}+\beta_{*}\right)^{\frac{1}{2}})}{\left(j^{2}+\beta_{*}\right)^{\frac{1}{2}}}\right)>0.$
Thus $a_{0,1}<0<c_{0,1}$. ∎
Lemma 5.3.
The coefficient $b_{3,0}$ exhibits the following limiting behaviors:
$b_{3,0}(h)=\frac{9}{1024\sqrt{2}}h^{-9/2}+\mathscr{o}\left(h^{-9/2}\right)~{}\textrm{as}~{}h\rightarrow 0^{+}\quad\textrm{and}\quad\displaystyle\lim_{h\rightarrow\infty}b_{3,0}(h)=b_{3,0,\infty}<0.$
Consequently, $b_{3,0}$ has only a finite number of zeros over $h>0$. In fact, $b_{3,0}$ has at least one zero at $h_{\textrm{crit}}=0.25065...$.
Proof.
We require computer assistance to prove the limiting behavior of $b_{3,0}$ as $h\rightarrow 0^{+}$ and $h\rightarrow\infty$. However, at no point in our computer-assisted proof do we rely on numerical approximations. We work directly with the exact expression for $b_{3,0}$ given in CompanionToTransverseInstabilityFiniteDepth.nb.
For the limit $h\rightarrow 0^{+}$, we substitute the expansion $\beta_{*}(h)\sim 4/3h^{2}$ from Proposition  3.1 into the exact expression of $b_{3,0}$. What remains is a function only of $h$. With the help of Mathematica’s Series function, we then perform a Puiseux series expansion of this function of $h$ about $h=0$ and, in particular, obtain the leading-order term of this expansion, giving us
$b_{3,0}(h)=\frac{9}{1024\sqrt{2}}h^{-9/2}+\mathscr{o}\left(h^{-9/2}\right)~{}\textrm{as}~{}h\rightarrow 0^{+}.$
The interested reader can verify this result in CompanionToTransverseInstabilityFiniteDepth.nb.
For the limit $h\rightarrow\infty$, we refer to $\beta_{*}(h)\rightarrow\beta_{*,\infty}$ from Proposition  3.1 in combination with Mathematica’s Limit function to calculate $b_{3,0,\infty}$ explicitly. The resulting expression we obtain for $b_{3,0,\infty}$ is initially quite complicated. However, we can compare this expression with the $b_{3,0}$ obtained in our infinite depth paper [17]. In particular, we show using Mathematica that the expression we obtain for $b_{3,0,\infty}$ in fact equals the $b_{3,0}$ of the infinite depth case. Using the expression for $b_{3,0}$ from our work in infinite depth in place of what we initially calculate for $b_{3,0,\infty}$ in this work, we find
$b_{3,0,\infty}=-\frac{(1+\xi^{2})\left(p(\xi)+q(\xi)\sqrt{\xi^{4}-1}\right)}{r(\xi)},$ (5.3)
where
$\xi$ $=\frac{1}{2}\left(3-3\left(2\sqrt{7}-1\right)^{-\frac{1}{3}}+\left(2\sqrt{7}-1\right)^{\frac{1}{3}}\right),$ (5.4a)
$p(\xi)$ $=87016-379693\xi+739050\xi^{2}-974082\xi^{3}+1042548\xi^{4}-898707\xi^{5}+559890\xi^{6}-214548\xi^{7}$
$\hskip 42.67912pt+22208\xi^{8}+25957\xi^{9}-15538\xi^{10}+3670\xi^{11}+100\xi^{12}-229\xi^{13}+54\xi^{14},$ (5.4b)
$q(\xi)$ $=-3760-127707\xi+448476\xi^{2}-659689\xi^{3}+526682\xi^{4}-230882\xi^{5}+32188\xi^{6}+23322\xi^{7}$
$\hskip 42.67912pt-14544\xi^{8}+3149\xi^{9}+280\xi^{10}-257\xi^{11}+54\xi^{12},$ (5.4c)
$r(\xi)$ $=64\left(\xi^{2}-3\xi+5\right)\left(\xi^{2}-3\xi+6\right)\left(\xi^{2}-6\xi+11+\sqrt{\xi^{4}-1}\right)\Big{(}\xi^{3}-\xi^{2}+\xi-1$
$\hskip 42.67912pt+\xi\sqrt{\xi^{4}-1}\Big{)}^{2}\sqrt{\xi(3-\xi)}.$ (5.4d)
Note that the variable $\xi$ in this paper is equivalent to the variable $\gamma_{1}$ in our infinite depth paper.
Directly substituting (5.4a) into (5.4b), we find
$p(\xi)=\frac{-97739871069+p_{1}+p_{2}+p_{3}}{4096\left(2\sqrt{7}-1\right)^{\frac{14}{3}}},$
where
$p_{1}$ $=78384770484\sqrt{7},$
$p_{2}$ $=-37428273\left(2\sqrt{7}-1\right)^{\frac{2}{3}}\left(953-232\sqrt{7}\right),\quad\textrm{and}$
$p_{3}$ $=-3\left(2\sqrt{7}-1\right)^{\frac{1}{3}}\left(64260927809-20320120798\sqrt{7}\right).$
Now we have the simple bounds
$2.6<\sqrt{7},\quad\left(2\sqrt{7}-1\right)^{\frac{1}{3}}<1.6,\quad\textrm{and}\quad\left(2\sqrt{7}-1\right)^{\frac{2}{3}}<2.7,$
so that we obtain the following rational lower bounds for $p_{1}$, $p_{2}$, and $p_{3}$:
$(78384770484)\cdot(2.6)<p_{1},$
$-(37428273)\cdot(2.7)\cdot(953-232\cdot 2.6)<p_{2},$
$-(3)\cdot(1.6)\cdot(64260927809-20320120798\cdot 2.6)<p_{3},$
from which it follows that
$p(\xi)>0.$
A similar argument shows that $q(\xi)>0$, and in [17] we show that $r(\xi)>0$ and $21<\xi<$ 2. It follows that
$b_{3,0,\infty}<0.$
In fact, if we numerically evaluate $b_{3,0,\infty}$, we find $b_{3,0,\infty}=-0.49476...$.
Given the behavior of $b_{3,0}$ as $h\rightarrow 0^{+}$ and $h\rightarrow\infty$, $b_{3,0}$ must have at least one zero in $(0,\infty)$. Since $b_{3,0}$ is analytic on $(0,\infty)$, its zero set $\mathcal{Z}$ cannot have an accumulation point in $(0,\infty)$. $\mathcal{Z}$ is bounded because $\lim_{h\to\infty}b_{3,0}(h)<0$. Moreover, 00 cannot be an accumulation point of $\mathcal{Z}$ because $\lim_{h\to 0^{+}}b_{3,0}(h)=\infty$. Thus $\mathcal{Z}$ is finite.
[…]
Figure 3. A plot of $b_{3,0}$ as a function of $h$. As $h\rightarrow 0^{+}$, we have $b_{3,0}\rightarrow+\infty$. As $h\rightarrow\infty$, we have $b_{3,0}\rightarrow-0.49476...$, in agreement with [17]. At $h_{crit}=0.25065....$, we have $b_{3,0}=0$. Thus, for this depth, we observe no transverse instability at $O(\varepsilon^{3})$.
Proof of Theorem 1.1
The characteristic polynomial of $\textrm{L}_{\varepsilon,\delta}-i\sigma I$ is
$\det(\textrm{L}_{\varepsilon,\delta}-i\sigma I-\lambda I)=\lambda^{2}-i(A+C)\lambda-AC-B^{2},$ (5.5)
whose discriminant is the real-valued function
$\Delta(\varepsilon,\delta)=-(A-C)^{2}+4B^{2}.$ (5.6)
The eigenvalues of $\textrm{L}_{\varepsilon,\delta}$ are
$\lambda_{\pm}=i\Big{(}\sigma+\frac{1}{2}(A+C)\Big{)}\pm\frac{1}{2}\sqrt{\Delta(\varepsilon,\delta)},$ (5.7)
where $A+C$ is the trace of $\text{L}_{\varepsilon,\delta}-i\sigma I$ and so is real-valued. Since $A$, $B$, and $C$ are real analytic in $(\varepsilon,\delta)$ in a neighborhood of $(0,0)$, so are $A+C$ and $\Delta(\varepsilon,\delta)$. Moreover, it follows from the expansions of $A$, $B$, and $C$ that $\Delta(\varepsilon,\delta)=O(\delta^{2})$ and $A+C=O(\delta)$ as $(\varepsilon,\delta)\rightarrow(0,0)$. We will complete the proof of Theorem 1.1 by showing that the characteristic polynomial (5.5) has a root with positive real part, that is, $\Delta(\varepsilon,\delta)>0$ for suitably small $\varepsilon$ and $\delta$.
Taking inspiration from our work in infinite depth [17], we map $\delta$ to a new variable $\theta$ as follows:
$\delta=\kappa_{0}\varepsilon^{2}+\theta\varepsilon^{3},\quad\textrm{where}\quad\kappa_{0}:=\frac{a_{2,0}-c_{2,0}}{a_{0,1}-c_{0,1}}.$ (5.8)
This transformation is well-defined by Lemma 5.2. Upon substituting (5.8) into the series expansions of $A$, $B$, and $C$ (5.2), we obtain the following expansion for $\Delta$:
$\Delta(\varepsilon,\delta)=\Delta(\varepsilon,\kappa_{0}\varepsilon^{2}+\theta\varepsilon^{3})=\varepsilon^{6}\Big{[}4b_{3,0}^{2}-(a_{0,1}-c_{0,1})^{2}\theta^{2}\Big{]}+O\left(\varepsilon^{7}\right).$ (5.9)
We refer to Section 6 in [17] for the detailed derivation of (5.9). We stress that the choice of $\kappa_{0}$ eliminates the contribution of the fourth-order coefficients $a_{4,0}$ and $c_{4,0}$ to the leading term in (5.9). Choosing $\theta$ such that
$|\theta|<\kappa_{1},\quad\textrm{where}\quad\kappa_{1}:=\frac{2|b_{3,0}|}{|a_{0,1}-c_{0,1}|},$ (5.10)
we appear to achieve $\Delta(\varepsilon,\delta)>0$ for all sufficiently small $\varepsilon>0$, as desired. This choice assumes, however, that $b_{3,0}$ does not vanish for all depths $h$. Lemma  5.3 proves this is not the case, unless $h=h_{crit}=0.25065...$. Thus, provided $h\neq h_{crit}$, we have $\Delta(\varepsilon,\delta)>0$ for all sufficiently small $\varepsilon>0$.
Returning to the expression for the eigenvalues (5.7), we find that $\text{Re}\lambda_{+}=\frac{1}{2}\sqrt{\Delta(\varepsilon,\delta)}>0$ for $h\neq h_{crit}$. Therefore, small-amplitude Stokes waves are unstable with respect to transverse perturbations with wave numbers $\alpha=\sqrt{\beta_{*}+\delta}$ for any $\delta\in(\varepsilon^{2}\kappa_{0}-\varepsilon^{3}\kappa_{1},\varepsilon^{2}\kappa_{0}+\varepsilon^{3}\kappa_{1})$, provided $\varepsilon$ is sufficiently small. Finally, by (5.9), we have $\text{Re}\lambda_{+}=O(\varepsilon^{3})$ as $\varepsilon\rightarrow 0$, completing the proof. ∎
Substituting (5.9) into (5.7) yields
$\lambda_{\pm}\equiv\lambda_{\pm}(\varepsilon,\theta)$ $=i\left[\sigma+\frac{a_{0,1}c_{2,0}-a_{2,0}c_{0,1}}{(a_{0,1}-c_{0,1})}\varepsilon^{2}+\frac{a_{0,1}+c_{0,1}}{2}\theta\varepsilon^{3}\right]$ (5.11)
±$\qquad\pm\frac{1}{2}\left[4b_{3,0}^{2}-(a_{0,1}-c_{0,1})^{2}\theta^{2}\right]^{\frac{1}{2}}|\varepsilon|^{3}+O(\varepsilon^{4}).$
By eliminating $\theta$ in from the real and imaginary parts of $\lambda_{\pm}$, we obtain
Corollary 5.4.
For sufficiently small $\varepsilon$, the curve $(-\kappa_{1},\kappa_{1})\ni\theta\mapsto\lambda_{+}(\varepsilon,\theta)$ (resp. $(-\kappa_{1},\kappa_{1})\ni\theta\mapsto\lambda_{-}(\varepsilon,\theta)$) is within $O(\varepsilon^{4})$ distance to the entire right (resp. left) half of the ellipse
$\frac{\lambda_{r}^{2}}{\left(b_{3,0}\varepsilon^{3}\right)^{2}}+\frac{\left(\lambda_{i}-\sigma-\left(\frac{a_{0,1}c_{2,0}-a_{2,0}c_{0,1}}{(a_{0,1}-c_{0,1})}\right)\varepsilon^{2}\right)^{2}}{\left(\frac{b_{3,0}(a_{0,1}+c_{0,1})}{a_{0,1}-c_{0,1}}\varepsilon^{3}\right)^{2}}=1$ (5.12)
in the complex plane. See Figure 1.
Appendix A Normal trace of vector fields
We recall the following classical trace operator.
Proposition A.1 (Section 3.2, Chapter IV, [9]).
Let $U\subset\mathbb{R}^{n}$ be Lipschitz domain with compact boundary and denote
$H_{\operatorname{div}}(U)=\{u\in L^{2}(U;\mathbb{R}^{n}):\operatorname{div}u\in L^{2}(U)\}.$ (A.1)
There exists a bounded linear operator $\gamma:H_{\operatorname{div}}(U)\to H^{\frac{1}{2}}(\partial U)$ such that $\gamma(u)=u\cdot\nu$ for $u\in C^{\infty}_{0}(\overline{U})$, where $\nu$ is the unit outward normal to $\partial U$. In particular, there exists a constant $C>0$ such that
$\|\gamma(u)\|_{H^{-\frac{1}{2}}(\partial U)}\leq C\Big{(}\|u\|_{L^{2}(U)}+\|\operatorname{div}u\|_{L^{2}(U)}\Big{)}\quad\forall u\in H_{\operatorname{div}}(U).$ (A.2)
Moreover, if $u\in H_{\operatorname{div}}(U)$ and $w\in H^{1}(U)$, then the Stokes formula
$\langle\gamma(u),\gamma_{0}(w)\rangle_{H^{-\frac{1}{2}}(\partial U),H^{\frac{1}{2}}(\partial U)}=\int_{U}u\cdot\nabla w+\int_{U}w\operatorname{div}u$ (A.3)
holds, where $\gamma_{0}(w)$ is the trace of $w$.
The following lemma provides an extension of the preceding trace operator for a cylindrical shell that is sufficient for our purposes.
Lemma A.2.
Let $U=\mathbb{T}\times(a,b)$, where $-\infty<a<b<\infty$. Since $\partial U=(\mathbb{T}\times\{a\})\cup(\mathbb{T}\times\{b\})$, we identify any function $g$ on $\partial U$ with the pair $(g_{1},g_{2})\equiv(g(\cdot,a),g(\cdot,b)):\mathbb{T}\to\mathbb{R}^{2}$. We denote a point in $U$ by $(x,z)$ and set
$\tilde{H}_{\operatorname{div}}(U)=\{u\in L^{2}(U;\mathbb{R}^{2}):\operatorname{div}u=\partial_{x}^{2}F\quad\text{for~{}some~{}}F\in H^{1}(U)\}.$ (A.4)
(i) There exists a linear operator $\tilde{\gamma}:\tilde{H}_{\operatorname{div}}(U)\to(H^{-\frac{1}{2}}(\mathbb{T}))^{2}$ with the following properties
  • (a)
    $\tilde{\gamma}(u)=\gamma(u)$ if $u\in\tilde{H}_{\operatorname{div}}(U)\cap H_{\operatorname{div}}(U)$.
  • (b)
    For some constant $C=C(a,b)>0$, we have
    $\|\tilde{\gamma}(u)\|_{(H^{-\frac{1}{2}}(\mathbb{T}))^{2}}\leq C(\|u\|_{L^{2}(U)}+\|\partial_{x}F\|_{L^{2}(U)}),$ (A.5)
    where the right-hand side of (A.5) is independent of the choice of $F$.
  • (c)
    If $u=\begin{bmatrix}u_{1}\\ 0\end{bmatrix}\in\tilde{H}_{\operatorname{div}}(U)$, then $\tilde{\gamma}(u)=0$.
(ii) We extend $\tilde{\gamma}$ to $\tilde{H}_{\operatorname{div}}(U)+H_{\operatorname{div}}(U)$ by defining
$\tilde{\gamma}(u)=\tilde{\gamma}(\tilde{v})+\gamma(v)$ (A.6)
if $u=\tilde{v}+v$, $\tilde{v}\in\tilde{H}_{\operatorname{div}}(U)$, $v\in H_{\operatorname{div}}(U)$. This definition is independent of the choice of the pair $(\tilde{v},v)$. As a consequence, if $u\in L^{2}(U;\mathbb{R}^{2})$ and $\operatorname{div}u=G+\partial_{x}^{2}F$ for some $G\in L^{2}(U)$ and $F\in H^{1}(U)$, then
$\|\tilde{\gamma}(u)\|_{(H^{-\frac{1}{2}}(\mathbb{T}))^{2}}\leq C(\|u\|_{L^{2}(U)}+\|G\|_{L^{2}(U)}+\|\partial_{x}F\|_{L^{2}(U)}).$ (A.7)
Proof.
(i) For any $u\in\tilde{H}_{\operatorname{div}}(U)$ with $\operatorname{div}u=\partial_{x}^{2}F$, using a lifting operator $\ell_{U}:(H^{\frac{1}{2}}(\mathbb{T}))^{2}\to H^{1}(U)$, we define $\tilde{\gamma}(u)\in(H^{-\frac{1}{2}}(\mathbb{T}))^{2}$ by
$\langle\tilde{\gamma}(u),g\rangle_{(H^{-\frac{1}{2}}(\mathbb{T}))^{2},(H^{\frac{1}{2}}(\mathbb{T}))^{2}}:=(u,\nabla\ell_{U}g)_{L^{2}(U)}-(\partial_{x}F,\partial_{x}\ell_{U}g)_{L^{2}(U)}\quad\forall g\in(H^{\frac{1}{2}}(\mathbb{T}))^{2}.$ (A.8)
Clearly (A.8) defines $\tilde{\gamma}(u)$ as a bounded linear functional on $(H^{\frac{1}{2}}(\mathbb{T}))^{2}$ with norm bounded by $C(\|u\|_{L^{2}(U)}+\|\partial_{x}F\|_{L^{2}(U)})$. Thus (A.5) is valid. This norm bound is independent of the choice of $F$. Indeed, if $\operatorname{div}u=\partial_{x}^{2}F=\partial_{x}^{2}F_{1}$, then $F(x,z)-F_{1}(x,z)$ is independent of $x$ due to the periodicity in $x$, whence $\partial_{x}F=\partial_{x}F_{1}$.
Now suppose that $u\in\tilde{H}_{\operatorname{div}}(U)\cap H_{\operatorname{div}}(U)$, so that $L^{2}(U)\ni\operatorname{div}u=\partial_{x}^{2}F$ for some $F\in H^{1}(U)$. The Stokes formula (A.3) implies
$\langle\gamma(u),g\rangle_{H^{-\frac{1}{2}}(\partial U),H^{\frac{1}{2}}(\partial U)}$ $=(u,\nabla\ell_{U}g)_{L^{2}(U)}+(\partial^{2}_{x}F,\ell_{U}g)_{L^{2}(U)}.$
Since $\partial_{x}^{2}F=\operatorname{div}u\in L^{2}(U)$, $F\in H^{1}(U)$, and $\ell_{U}g\in H^{1}(U)$, we have $(\partial^{2}_{x}F,\ell_{U}g)_{L^{2}(U)}=-(\partial_{x}F,\partial_{x}\ell_{U}g)_{L^{2}(U)}$ by the periodicity in $x$, and hence for all $g\in H^{\frac{1}{2}}(\mathbb{T}))^{2}$,
$\langle\gamma(u),g\rangle_{H^{-\frac{1}{2}}(\partial U),H^{\frac{1}{2}}(\partial U)}=(u,\nabla\ell_{U}g)_{L^{2}(U)}-(\partial_{x}F,\partial_{x}\ell_{U}g)_{L^{2}(U)}=\langle\tilde{\gamma}(u),g\rangle_{(H^{-\frac{1}{2}}(\mathbb{T}))^{2},(H^{\frac{1}{2}}(\mathbb{T}))^{2}}.$
Therefore, $\tilde{\gamma}=\gamma$ on $\tilde{H}_{\operatorname{div}}(U)\cap H_{\operatorname{div}}(U)$.
Assume that $u=\begin{bmatrix}u_{1}\\ 0\end{bmatrix}\in\tilde{H}_{\operatorname{div}}(U)$. Then $\partial_{x}^{2}F=\operatorname{div}u=\partial_{x}u_{1}$ for some $F\in H^{1}(U)$, and hence $u_{1}(x,z)=\partial_{x}F(x,z)+f(z)$ for some $f\in L^{2}((a,b))$. The definition (A.8) implies that for all $g\in(H^{\frac{1}{2}}(\mathbb{T}))^{2}$ we have
$\langle\tilde{\gamma}(u),g\rangle_{(H^{-\frac{1}{2}}(\mathbb{T}))^{2},(H^{\frac{1}{2}}(\mathbb{T}))^{2}}$ $=(u_{1},\partial_{x}\ell_{U}g)_{L^{2}(U)}-(\partial_{x}F,\partial_{x}\ell_{U}g)_{L^{2}(U)}$
$=(f,\partial_{x}\ell_{U}g)_{L^{2}(U)}=\int_{a}^{b}f(z)\int_{\mathbb{T}}\partial_{x}\ell_{U}g(x,z)dxdz=0$
due to the periodicity in $x$. Thus $\tilde{\gamma}(u)=0$.
(ii) Next, we show that the definition (A.6) of the extended $\tilde{\gamma}$ is well defined. Suppose that $u=\tilde{v}+v=\tilde{w}+w$, where $\tilde{v}$, $\tilde{w}\in\tilde{H}_{\operatorname{div}}(U)$, and $v$, $w\in H_{\operatorname{div}}(U)$. Then $\tilde{v}-\tilde{w}=w-v\in\tilde{H}_{\operatorname{div}}(U)\cap H_{\operatorname{div}}(U)$, and thus $\tilde{\gamma}(\tilde{v}-\tilde{w})=\gamma(w-v)$. It follows that $\tilde{\gamma}(\tilde{v})+\gamma(v)=\tilde{\gamma}(\tilde{w})+\gamma(w)$.
Suppose that $u\in L^{2}(U;\mathbb{R}^{2})$ and $\operatorname{div}u=G+\partial_{x}^{2}F$ for some $G\in L^{2}(U)$ and $F\in H^{1}(U)$. Let $\phi:U\to\mathbb{R}$ be the $H^{2}(U)\cap H^{1}_{0}(U)$ solution to the Poisson equation $\Delta\phi=G$, and set $v=\nabla\phi$, $\tilde{v}=u-v$. Then $v\in H_{\operatorname{div}}(U)$, $\tilde{v}\in\tilde{H}_{\operatorname{div}}(U)$ with $\operatorname{div}\tilde{v}=\partial_{x}^{2}F$, and $\|\phi\|_{H^{2}(U)}\leq C\|G\|_{L^{2}(U)}$. Therefore it follows from (A.6), (A.5) and (A.2) that
$\|\tilde{\gamma}(u)\|_{(H^{-\frac{1}{2}}(\mathbb{T}))^{2}}$ $\leq\|\tilde{\gamma}(\tilde{v})\|_{(H^{-\frac{1}{2}}(\mathbb{T}))^{2}}+\|\gamma(v)\|_{(H^{-\frac{1}{2}}(\mathbb{T}))^{2}}$
$\leq C\left(\|\tilde{v}\|_{L^{2}(U)}+\|\partial_{x}F\|_{L^{2}(U)}+\|v\|_{L^{2}(U)}+\|\operatorname{div}v\|_{L^{2}(U)}\right)$
$\leq C\left(\|u\|_{L^{2}(U)}+\|v\|_{L^{2}(U)}+\|\partial_{x}F\|_{L^{2}(U)}+\|v\|_{L^{2}(U)}+\|\Delta\phi\|_{L^{2}(U)}\right),$
which implies (A.7). ∎
Acknowledgment. The work of HQN was partially supported by NSF grant DMS-2205710. The work of RPC was partially supported by NSF grant DMS-2402044.

### v2

@@
 In this paper we consider the case of finite depth and prove that the same spectral instability result holds for all but finitely many values of the depth.
-The computations and some aspects of the theory are considerably more complicated in the finite depth case.
+The computations are considerably more complicated in the finite depth case.
 Key words and phrases:
@@
 It has been known for a century that a curve of small-amplitude Stokes waves exists [35, 31, 42].
-Several decades ago it was proven that the curve extends to large amplitudes as well [24].
+Several decades ago it was proven that the curve extends to large amplitudes as well [28, 27, 43, 3, 39].
 A brief history of longitudinal (non-transverse) instabilities was given in [18], to which the reader is referred.
@@
 To bring the references up to date, we mention some very recent work on longitudinal instability by Berti et al [9] and on near-extreme Stokes waves by Deconinck et al [21].
-It is important to note that there are several other models of water waves for which the transverse instability has been studied rigorously, for instance for gravity-capillary waves.
+It is important to note that there are several other models of water waves for which the transverse instability has been studied rigorously, for instance for gravity-capillary waves [40, 24].
 As in [18] we now specify the parameters of our problem.
@@
 This curve is parametrized by a small parameter $\varepsilon$ which represents the wave amplitude of the Stokes waves.
-Such a steady wave can be described in the moving $(x,z)$ plane (where $x-ct$ is replaced by $x$) by its free surface $S=\{(x,y,z)\ |\ z=\eta^{*}(x;\varepsilon)\}$ and by its velocity potential $\psi^{*}(x;\varepsilon)$ restricted to $S$.
+Such a steady wave can be described in the moving $(x,z)$ plane (where xctx-ct is replaced by $x$) by its free surface $S=\{(x,y,z)\ |\ z=\eta^{*}(x;\varepsilon)\}$ and by its velocity potential $\psi^{*}(x;\varepsilon)$ restricted to $S$.
 We take the perturbation of $\eta^{*}$ to have the form $\overline{\eta}(x)e^{\lambda t+i\alpha y}$, where $\overline{\eta}$ has the same period $2\pi$ as the Stokes wave, $\lambda\in\mathbb{C}$ is the growth rate of the perturbation, and $\alpha\in\mathbb{R}$ is the transverse wave number of the perturbation.
+In other words, we consider purely transverse perturbations, leaving out any effect of modulational instabilities.
 The problem is to find at least one value of $\alpha$ that leads to instability, that is, $\text{Re}\lambda>0$.
@@
 Motivated by [34], we determine a resonant transverse wave number $\alpha_{*}$ so that the unperturbed operator $\mathcal{L}_{h,0,\beta_{*}}$ with $\beta_{*}=\alpha_{*}^{2}$ has an imaginary double eigenvalue $i\sigma$.
-This eigenvalue corresponds to the lowest possible resonance that generates a Type II transverse instability according to McLean [30], of which there are infinitely many higher-order resonances that potentially generate higher-order transverse instabilities.
+This eigenvalue corresponds to the lowest possible resonance ($m=1$ in (3.4)) that generates a Type II transverse instability according to McLean [34], of which there are infinitely many higher-order resonances that potentially generate higher-order transverse instabilities.
 We expect however that higher-order transverse instabilities have smaller growth rates for small Stokes waves.
+More precisely, we conjecture that the $m^{th}$ resonance in (3.4) would lead to instabilities at order $\varepsilon^{2m+1}$.
 In order to capture the transverse instabilities we introduce a new small parameter $\delta$ for the perturbation of $\beta$ about $\beta_{*}$.
@@
 Remark 1.3.
+Shortly after the appearance of the preprint version of this paper, the preprint [23] of Jiao-Rodrigues-Yang proved the oblique instability of small Stokes waves when additional localized perturbations in the longitudinal direction are included.
+They thereby obtained an $\varepsilon^{2}$-order instability, which is expected from the dominant modulational instability [36, 5].
+Remark 1.4.
 Figure 1 compares our transverse instability isolae with the numerically computed eigenvalues for a Stokes wave with amplitude $\varepsilon=0.01$ at the three depths $h=1$, $3/2$, and $2$.
@@
 The main result that is required here is Theorem 2.6, which is devoted to the three-dimensional Dirichlet-Neumann operator $\mathcal{G}_{\varepsilon,\beta}$ that occurs in the linearized operator subject to the two-dimensional conformal mapping, and the proof of its analyticity in $\varepsilon$ and $\delta$.
-The lack of a trace theorem from $H^{1/2}$ of a domain to its boundary necessitates a careful alternative analysis.
-Compared to the infinite depth case [17], the proof of analyticity is considerably more difficult due to the dependence on $\varepsilon$ of the strip $\Omega_{\varepsilon}$.
+The analyticity is proven by an approach inspired by Groves [22] and is significantly shorter than our direct approach in [18].
 Theorem 1.1 is then reduced to studying the eigenvalues of the linearized operator $\mathcal{L}_{h,\varepsilon,\beta}$, which has a Hamiltonian form and is reversible.
 Section 3 is devoted to a discussion of the first resonance on the imaginary axis when $\varepsilon=0$, the introduction of Kato’s perturbed basis, and the reduction to the study of the eigenvalues of a reduced $2\times 2$ matrix $\textrm{L}_{h,\varepsilon,\delta}$.
-In Section 4 we perform extremely tedious expansions of $\mathcal{L}_{h,\varepsilon,\beta_{*}+\delta}$ out to third order in both $\varepsilon$ and $\delta$, which are considerably more tedious than the corresponding expansions in [17, 32, 4].
+In Section 4 we perform expansions of $\mathcal{L}_{h,\varepsilon,\beta_{*}+\delta}$ out to third order in both $\varepsilon$ and $\delta$, which are considerably more tedious than the corresponding expansions in [18, 36, 5].
 We use the expansions of $\mathcal{L}_{h,\varepsilon,\beta_{*}+\delta}$ to compute the expansions of the Kato basis vectors and of the matrix $\textrm{L}_{h,\varepsilon,\delta}$.
@@
 The full equation that defines the isola of eigenvalues up to $O(\varepsilon^{4})$ is given in Corollary 5.4.
-Finally, we prove a trace result in Appendix A.1, which is used in the proof of the analyticity of the operator $\mathcal{G}_{\varepsilon,\beta}$.
 2. Transverse perturbations of Stokes waves
@@
 The second equation is Bernoulli’s, which follows from the pressure being constant along the free surface; the third equation expresses the kinematic boundary condition that particles on the surface remain there; the last condition asserts that there is no flux of water out of the bottom boundary $\{z=-h\}$.
-For convenience we will take $P=0$.
+Without loss of generality we will take $P=0$.
 In order to reduce the system to the free surface $S$, we introduce the Dirichlet-Neumann operator $G(\eta)$ associated to $D$, namely,
@@
 $\begin{cases}\partial_{t}\eta=c\partial_{x}\eta+G(\eta)\psi=:F_{1}(\eta,\psi,c),\\ \partial_{t}\psi=c\partial_{x}\psi-\frac{1}{2}|\nabla_{x,y}\psi|^{2}+\frac{1}{2}\frac{\big(G(\eta)\psi+\nabla_{x,y}\psi\cdot\nabla_{x,y}\eta\big)^{2}}{1+|\nabla_{x,y}\eta|^{2}}-g\eta=:F_{2}(\eta,\psi,c),\end{cases}$ (2.5)
-By the change of variables
-$c\to\sqrt{g}{c},\quad\eta(x,y,t)\to\eta(x,y,\sqrt{g}t),\quad\psi(x,y,t)\to\sqrt{g}\psi(,x,y,\sqrt{g}t),$
-we henceforth assume without loss of generality that $g=1$.
+For any $a>0$, if we make the change of variables
+$\eta(x,y,t)=\frac{1}{a}\tilde{\eta}(ax,ay,\sqrt{g}t),\quad\psi(x,y)=\frac{\sqrt{g}}{a\sqrt{a}}\tilde{\psi}(ax,ay,\sqrt{g}t),$
+[…]
 By a Stokes wave we mean a periodic steady solution of (2.5) that is independent of $y$.
-Without loss of generality, we henceforth consider $2\pi$-periodic Stokes waves.
+In view of the aforementioned scaling, without loss of generality, we henceforth consider $2\pi$-periodic Stokes waves.
 As has been known for over a century, there exists a curve of small Stokes waves parametrized analytically by the small amplitude $\varepsilon$.
@@
   • (i) the mapping $(-\varepsilon_{St}(s),\varepsilon_{St}(s))\ni\varepsilon\to\big(\eta^{*}(\cdot;\varepsilon),\psi^{*}(\cdot;\varepsilon),c^{*}(\varepsilon)\big)\in H^{s}(\mathbb{T})\times H^{s}(\mathbb{T})\times\mathbb{R}$ is analytic;
-  • (ii) $\eta^{*}(\cdot;\varepsilon)$ is even and has average zero, and $\psi^{*}(\cdot;\varepsilon)$ is odd.
+  • (ii) $\eta^{*}(\cdot;\varepsilon)$ is even, and $\psi^{*}(\cdot;\varepsilon)$ is odd.
 Proof.
@@
 where
-$B(\eta)\psi:=\frac{G(\eta)\psi+\nabla\eta\cdot\nabla\psi}{1+|\nabla\eta|^{2}},\quad V(\eta)\psi:=\nabla\psi-B\nabla\psi.$ (2.15)
+$B(\eta)\psi:=\frac{G(\eta)\psi+\nabla\eta\cdot\nabla\psi}{1+|\nabla\eta|^{2}},\quad V(\eta)\psi:=\nabla\psi-\nabla\eta B(\eta)\psi.$ (2.15)
 Proof.
@@
     The “conformal depth” $h_{\varepsilon}$ is given by
-[…]
+    $h_{\varepsilon}=h+\frac{1}{2\pi}\int_{0}^{2\pi}\eta^{*}(\zeta(x))\ dx.$ (2.27)
 Here we have defined the Fourier transform as
-[…]
+$\widehat{f}(k)=\int_{\mathbb{T}}f(x)e^{-ikx}dx;$
 Proof.
@@
 We will prove (iii) and (iv).
-For brevity, denote $f(x)=X(x,0)-x$.
+For brevity, denote f(x)=X(x,0)xf(x)=X(x,0)-x.
 We have $\Delta X=0$ and $\partial_{z}X(x,-h_{\varepsilon})=0$ as well as $\Delta Z=0$ and $Z(x,-h_{\varepsilon})=-h$.
@@
 $\hat{f}(k)=-i\frac{\mathrm{sign}(k)}{\tanh(|k|h_{\varepsilon})}\ \widehat{\eta^{*}\circ\zeta}(k).$ (2.31)
-Substituting this formula for $\hat{f}$ into (2.28) and (2.29), and noting that $X(x,z)-x$ is odd in $x$, we obtain (2.25) and (2.26).
+Substituting this formula for $\hat{f}$ into (2.28) and (2.29), and noting that X(x,z)xX(x,z)-x is odd in $x$, we obtain (2.25) and (2.26).
 Thus we have derived the formulas (2.25), (2.26), and (2.27). Moreover, by evaluating (2.25) at $z=0$ and invoking (2.27), we obtain
@@
 We emphasize that the operator $\mathcal{G}_{\varepsilon,\beta}$ and the constants $\iota$ and $\varepsilon_{0}(s)$ all depend on the depth $h$.
-Remark 2.7.
 In the infinite depth case, Theorem 2.6 was proven in Theorem 2.7, [18].
 The assertions (i) and (iii) can be obtained similarly to the infinite depth case.
-On the other hand, the analyticity of $\mathcal{G}_{\varepsilon,\beta}$ in $(\varepsilon,\beta)$ is more difficult to establish due to the dependence on $\varepsilon$ of the flat domain $\Omega_{\varepsilon}$.
-We will be led to proving the analyticity of solutions to the second-order elliptic problem (2.44) where the highest order term has an $\varepsilon$-dependent coefficient.
-Moreover, the case $s\in[-\frac{1}{2},1]$ will be deduced by interpolation from the cases $s=-\frac{1}{2}$ and $s>1$, where $s=-\frac{1}{2}$ requires the nonstandard trace result in Lemma A.2.
-Proof.
+Regarding the analyticity of $G_{\varepsilon,\beta}$ in $(\varepsilon,\beta)$, the direct proof in [18] can be adapted to the finite depth case.
+However, the low regularity case $s=\frac{1}{2}$ would require a careful treatment, due to the $\varepsilon$-dependence of the conformal domain $\{-h_{\varepsilon}<z<0\}$ and the trace operator for vector fields not being in $H_{\operatorname{div}}$ (see (2.37)) which arise if one changes $z\mapsto z\frac{h}{h_{\varepsilon}}$ to cope with $h_{\varepsilon}$.
+We will provide an alternative, shorter proof using the analytic implicit function theorem, in the spirit of the proof of Lemma 2.2 in [22].
+We will need the following classical trace operator.
+Proposition 2.7 (Section 3.2, Chapter IV, [10]).
+Let $U\subset\mathbb{R}^{n}$ be Lipschitz domain with compact boundary and denote
+$H_{\operatorname{div}}(U)=\{u\in L^{2}(U;\mathbb{R}^{n}):\operatorname{div}u\in L^{2}(U)\}.$ (2.37)
+There exists a bounded linear operator $\gamma:H_{\operatorname{div}}(U)\to H^{-\frac{1}{2}}(\partial U)$ such that $\gamma(u)=u\cdot\nu$ for $u\in C^{\infty}_{0}(\overline{U})$, where $\nu$ is the unit outward normal to $\partial U$.
+In particular, there exists a constant $C>0$ such that
+$\|\gamma(u)\|_{H^{-\frac{1}{2}}(\partial U)}\leq C\Big(\|u\|_{L^{2}(U)}+\|\operatorname{div}u\|_{L^{2}(U)}\Big)\quad\forall u\in H_{\operatorname{div}}(U).$ (2.38)
+Moreover, if $u\in H_{\operatorname{div}}(U)$ and $w\in H^{1}(U)$, then the Stokes formula
+$\langle\gamma(u),\gamma_{0}(w)\rangle_{H^{-\frac{1}{2}}(\partial U),H^{\frac{1}{2}}(\partial U)}=\int_{U}u\cdot\nabla w+\int_{U}w\operatorname{div}u$ (2.39)
+holds, where $\gamma_{0}(w)$ is the trace of $w$.
+Proof of Theorem 2.6
 We assume without loss of generality that $\alpha>0$ and denote $\mathbb{T}^{2}_{\alpha}=\mathbb{T}\times(\mathbb{R}/2\pi\alpha\mathbb{Z})$.
 Let $(X,Z):\Omega_{\varepsilon}\to\Omega$ be the Riemann mapping in Theorem 2.4.
-It has the Jacobian $\mathcal{J}=\mathcal{J}(x,z,\varepsilon)=|\partial_{x}X|^{2}+|\partial_{z}X|^{2}$.
+It has the Jacobian $\mathcal{J}=\mathcal{J}(x,z,\varepsilon)=|\partial_{x}X|^{2}+|\partial_{z}X|^{2}$, which is analytic in $\varepsilon$ near 0.
 For an arbitrary function $f\in H^{\frac{1}{2}}(\mathbb{T})$, we will define the function $\Theta_{f}$ by solving
@@
 $\mathcal{G}_{\varepsilon,\beta}f=\partial_{z}\Theta_{f}(\cdot,0),$ (2.41)
-where $\partial_{z}\Theta_{f}(\cdot,0)$ is the trace $\gamma(\nabla\Theta_{f})$ given in Proposition A.1.
+where $\partial_{z}\Theta_{f}(\cdot,0)$ is the trace $\gamma(\nabla\Theta_{f})$ given in Proposition 2.7.
 As shown in the proof of Theorem 2.7 in [18], for any $s\geq\frac{1}{2}$ there exists $\varepsilon_{0}(s)>0$ such that
@@
 This proves (2.34).
-The remainder of this proof is devoted to the analyticity (2.35).
-We fix $s\geq\frac{1}{2}$ unless stated otherwise.
-We first note that since $\eta^{*}$ is analytic in $\varepsilon\in(-\varepsilon_{0}(s),\varepsilon_{0}(s))$ with values in Sobolev spaces, so are $\zeta$, $(X,Z)-\operatorname{I}$, $\mathcal{J}$, and $h_{\varepsilon}$.
-For fixed $f\in H^{s}(\mathbb{T})$, we let $\Theta(\varepsilon,\beta)\equiv\Theta(x,z;\varepsilon,\beta)$ be the solution of (2.37), which is posed on $\Omega_{\varepsilon}=\{(x,z):x\in\mathbb{T},~{}z\in(-h_{\varepsilon},0)\}$.
-Compared to the infinite depth case considered in [17], the extra complication comes from the $\varepsilon$-dependence of the domain $\Omega_{\varepsilon}$.
-To cope with this we make the simple rescaling
+The remainder of this proof is devoted to the analyticity (2.35) of $\mathcal{G}_{\varepsilon,\beta}$. We denote
+$\Omega_{0}:=\{(x,z):x\in\mathbb{T},~z\in(-h,0)\}.$
+We fix $\beta_{0}>0$ and $f\in H^{s}(\mathbb{T})$ with $s\geq\frac{1}{2}$.
+We first prove that for any $F_{0}\in H^{s-\frac{1}{2}}(\Omega_{0})$ and $G\in(H^{s-\frac{1}{2}}(\Omega_{0}))^{2}$, the problem
+$\begin{cases}\Delta_{x,z}v-\beta_{0}v=F_{0}+\operatorname{div}G\quad\text{ in }\Omega_{0},\\ v(x,0)=f(x),\\ \partial_{z}v(x,-h)=G_{2}(x,-h)\end{cases}$ (2.47)
+has a unique solution $v\in H^{s+\frac{1}{2}}(\Omega_{0})$ in the following sense: $v|_{\{z=0\}}=f$ in the trace sense and
+$\int_{\Omega_{0}}\nabla_{x,z}v\cdot\nabla_{x,z}\varphi+\beta_{0}v\varphi\,dxdz=\int_{\Omega_{0}}-F_{0}\varphi+G\cdot\nabla_{x,z}\varphi\,dxdz\quad\forall\varphi\in H^{1,0}(\Omega_{0}),$ (2.48)
+where $H^{1,0}(\Omega_{0})=\{u\in H^{1}(\Omega_{0}):~u|_{\{z=0\}}=0\}$.
+Note that (2.48) implies $\Delta_{x,z}v-\operatorname{div}G=\operatorname{div}(\nabla_{x,z}v-G)=\beta_{0}v+F_{0}\in L^{2}(\Omega_{0})$ in the sense of distributions, and hence
+$\int_{\Omega_{0}}(\nabla_{x,z}v-G)\cdot\nabla_{x,z}\varphi+\operatorname{div}(\nabla_{x,z}v-G)\varphi\,dxdz=0\quad\forall\varphi\in H^{1,0}(\Omega_{0}).$
+In view of the Stokes formula (2.39), we deduce that the normal trace of $\nabla_{x,z}v-G$ at $\{z=-h\}$ vanishes in $H^{-\frac{1}{2}}(\mathbb{T})$, justifying the Neumann condition $\partial_{z}v(x,-h)=G_{2}(x,-h)$.
+For $s=\frac{1}{2}$, the existence and uniqueness of $v\in H^{1}(\Omega_{0})$ follow from the Lax-Milgram theorem and the lifting operator from $H^{s}(\mathbb{T}^{d})$ to $H^{s+\frac{1}{2}}(\Omega_{0})$.
+Moreover, we have
+$\|v\|_{H^{1}(\Omega_{0})}\leq C(\|F_{0}\|_{L^{2}(\Omega_{0})}+\|G\|_{(L^{2}(\Omega_{0}))^{2}}+\|f\|_{H^{\frac{1}{2}}(\mathbb{T})}).$
+For $s>\frac{1}{2}$, it remains to obtain the regularity $v\in H^{s+\frac{1}{2}}(\Omega_{0})$. We decompose $v=v_{1}+v_{2}$, where $v_{1}\in H^{1}(\Omega_{0})$ solves
+$\begin{cases}\Delta_{x,z}v_{1}-\beta_{0}v_{1}=F_{0}\quad\text{ in }\Omega_{0},\\ v_{1}(x,0)=f(x),\\ \partial_{z}v_{1}(x,-h)=0,\end{cases}$ (2.49)
+and $v_{2}\in H^{1}(\Omega_{0})$ solves
+$\begin{cases}\Delta_{x,z}v_{2}-\beta_{0}v_{2}=\operatorname{div}G\quad\text{ in }\Omega_{0},\\ v_{2}(x,0)=0,\\ \partial_{z}v_{2}(x,-h)=G_{2}(x,-h).\end{cases}$ (2.50)
+Since $F_{0}\in H^{s+\frac{1}{2}}(\Omega_{0})$ has nonnegative regularity, combining standard elliptic regularity with linear interpolation yields $v_{1}\in H^{s+\frac{1}{2}}(\Omega_{0})$.
+This is also the case for $v_{2}$ provided $s\geq\frac{3}{2}$, so that $\operatorname{div}G\in H^{s-\frac{3}{2}}(\Omega_{0})$ has nonnegative regularity.
+Since the mapping $H^{s-\frac{1}{2}}(\Omega_{0})\ni G\mapsto v_{2}\in H^{s+\frac{1}{2}}(\Omega_{0})$ is linear, the remaining case $s\in(\frac{1}{2},\frac{3}{2})$ follows by linear interpolation between $s=\frac{1}{2}$ and $s=\frac{3}{2}$.
+We denote the mapping $(F_{0},G,f)\mapsto v$ by
+$v=\mathcal{S}(F_{0},G,f).$ (2.51)
+Since
+$\mathcal{S}(F_{0},G,f)=\mathcal{S}(0,0,f)+\mathcal{S}(F_{0},0,0)+\mathcal{S}(0,G,0)$
+and each operator on the right-hand side is linear in its respective argument, it follows that
+$\mathcal{S}:H^{s-\frac{1}{2}}(\Omega_{0})\times(H^{s-\frac{1}{2}}(\Omega_{0}))^{2}\times H^{s}(\mathbb{T}^{d})\to H^{s+\frac{1}{2}}(\Omega_{0})\quad\text{is~ analytic}.$ (2.52)
+Returning to the problem (2.40), in order to cope with the $\varepsilon$-dependence of $\Omega_{\varepsilon}$, we make the simple rescaling
 $\Theta(x,z;\varepsilon,\beta)=\Psi\left(x,\frac{h}{h_{\varepsilon}}z;\varepsilon,\beta\right),\quad\mathcal{J}(x,z;\varepsilon)=\tilde{J}\left(x,\frac{h}{h_{\varepsilon}}z;\varepsilon\right),$
 so that $\Psi\equiv\Psi(\varepsilon,\beta):\Omega_{0}\to\mathbb{R}$ satisfies
-$\begin{cases}\left(\partial_{x}^{2}+\big{(}\frac{h}{h_{\varepsilon}}\big{)}^{2}\partial_{z}^{2}\right)\Psi-\beta\tilde{J}\Psi=0\quad\text{ in }\Omega_{0}\ ,\\ \Psi(x,0)=f(x),\quad\partial_{z}\Psi(x,-h)=0,\end{cases}$ (2.44)
-where $\Omega_{0}:=\{(x,z):x\in\mathbb{T},~{}z\in(-h,0)\}$.
-Thus the $\varepsilon$-dependence of $\Omega_{\varepsilon}$ is now encoded in the coefficient $\frac{h}{h_{\varepsilon}}$ of the highest order term $\partial_{z}^{2}\Psi$.
-Recalling (2.38) and using the chain rule, we obtain
-$\mathcal{G}_{\varepsilon,\beta}f=\frac{h}{h_{\varepsilon}}\partial_{z}\Psi(\varepsilon,\beta)|_{z=0}=\frac{h}{h_{\varepsilon}}\tilde{\gamma}(\nabla\Psi(\varepsilon,\beta)),$ (2.45)
-where $\tilde{\gamma}$ is defined in Lemma A.2. We divide the rest of the proof into several steps.
-1. We claim that the mapping
-$(-\varepsilon_{0}(s),\varepsilon_{0}(s))\times(0,\infty)\ni(\varepsilon,\beta)\mapsto\Psi(\varepsilon,\beta)\in H^{s+\frac{1}{2}}(\Omega_{0})$ (2.46)
-is $C^{\infty}$ for $s\geq\frac{1}{2}$.
-For sufficiently small $\varepsilon_{0}(s)$, we have $h_{\varepsilon}\geq\frac{1}{2}h$ and $\tilde{J}\geq\frac{1}{2}$. Then, elliptic estimates yield the uniform (in $\varepsilon$) bound
-$\|\Psi(\varepsilon,\beta)\|_{H^{s+\frac{1}{2}}(\Omega_{0})}\leq C_{1}(\beta,s)\|f\|_{H^{s}(\mathbb{T})}\quad\forall\varepsilon\in(-\varepsilon_{0}(s),\varepsilon_{0}(s)),$ (2.47)
-where $C_{1}(\beta,s)$ is bounded for $\beta$ in compact subsets of $(0,\infty)$.
-For fixed $\varepsilon\in(-\varepsilon_{0}(s),\varepsilon_{0}(s))$, the differentiability of $\Psi(\varepsilon,\beta)$ in $\beta$ can be proven analogously to the proof of Theorem 2.7, [17].
-On the other hand, for fixed $\beta>0$, the differentiability of $\Psi(\varepsilon,\beta)$ in $\varepsilon$ requires more attention since $\varepsilon$ appears in the highest order term $\frac{h}{h_{\varepsilon}}\partial_{z}^{2}\Psi$.
-We let $u_{\varepsilon,\beta}$ be the solution of the inhomogeneous problem
-$\begin{cases}(\partial_{x}^{2}+\big{(}\frac{h}{h_{\varepsilon}}\big{)}^{2}\partial_{z}^{2})u_{\varepsilon,\beta}-\beta\tilde{J}(\varepsilon)u_{\varepsilon,\beta}=\beta[\frac{d}{d\varepsilon}\tilde{J}(\varepsilon)]\Psi(\varepsilon,\beta)-\partial_{z}^{2}\Psi(\varepsilon,\beta)\frac{d}{d\varepsilon}\big{(}\frac{h}{h_{\varepsilon}}\big{)}^{2}:=H_{\varepsilon,\beta}\quad\text{in }\Omega_{0},\\ u_{\varepsilon,\beta}(\cdot,0)=0,\quad\partial_{z}u_{\varepsilon,\beta}(\cdot,-h)=0,\end{cases}$ (2.48)
-which is obtained by formally differentiating (2.44) in $\varepsilon$.
-We will prove that
-$\|u_{\varepsilon,\beta}\|_{H^{s+\frac{1}{2}}}\leq C_{2}(\beta,s)\|f\|_{H^{s}(\mathbb{T})}\quad\forall\varepsilon\in(-\varepsilon_{0}(s),\varepsilon_{0}(s)),$ (2.49)
-where $C_{2}(\beta,s)$ is independent of $\varepsilon$ and is bounded for $\beta$ in compact subsets of $(0,\infty)$.
-Consider first the smoother case $s\geq\frac{3}{2}$.
-Then, $s-\frac{3}{2}\geq 0$ and (2.47) implies that $\|H_{\varepsilon,\beta}\|_{H^{s-\frac{3}{2}}(\Omega_{0})}\leq C(\beta,s)\|f\|_{H^{s}(\mathbb{T})}$ uniformly in $\varepsilon$.
-Thus (2.49) follows from the standard elliptic estimates.
-It remains to prove (2.49) for $s=\frac{1}{2}$, since the mapping $f\mapsto u_{\varepsilon,\beta}$ is linear so that we can use interpolation.
-We do know that $\|\Psi(\varepsilon,\beta)\|_{H^{1}(\Omega_{0})}\leq C_{1}\|f\|_{H^{\frac{1}{2}}(\mathbb{T})}$ as in (2.47).
-Suppose for the moment that $u=u_{\varepsilon,\beta}$ is a smooth solution of (2.48).
-Let us denote
-$H^{1,0}(\Omega_{0})=\{\varphi\in H^{1}(\Omega_{0}):~{}u(\cdot,0)=0\},$
-which is a Hilbert space with the inner product $(\varphi,\tilde{\varphi})_{H^{1,0}}:=\int_{\Omega_{0}}\nabla\varphi\cdot\nabla\tilde{\varphi}$.
-Multiplying (2.48) by $\varphi\in H^{1,0}(\Omega_{0})$ and integrating by parts using the homogeneous Neumann boundary condition in (2.48), we find that
-[…]
-$\quad+\left\langle\partial_{z}^{2}\Psi(\varepsilon,\beta)\frac{d}{d\varepsilon}\big{(}\frac{h}{h_{\varepsilon}}\big{)}^{2},\varphi\right\rangle_{(H^{1,0})^{\prime},H^{1,0}},$
-where, since $\partial_{z}\Psi(\varepsilon,\beta)|_{z=-h}=0$, we denote
-$\left\langle\partial_{z}^{2}\Psi(\varepsilon,\beta)\frac{d}{d\varepsilon}\big{(}\frac{h}{h_{\varepsilon}}\big{)}^{2},\varphi\right\rangle_{(H^{1,0})^{\prime},H^{1,0}}:=-\int_{\Omega_{0}}\partial_{z}\Psi(\varepsilon,\beta)\frac{d}{d\varepsilon}\big{(}\frac{h}{h_{\varepsilon}}\big{)}^{2}\partial_{z}\varphi dxdz.$
-From the above considerations, we define $u\in H^{1,0}(\Omega_{0})$ to be a weak solution of (2.48) if
-[…]
-$\quad-\int_{\Omega_{0}}\partial_{z}\Psi(\varepsilon,\beta)\frac{d}{d\varepsilon}\big{(}\frac{h}{h_{\varepsilon}}\big{)}^{2}\partial_{z}\varphi dxdz\quad\forall\varphi\in H^{1,0}(\Omega_{0}).$
-Since $\beta>0$, $h_{\varepsilon}\geq\frac{1}{2}h$, and $\tilde{J}\geq\frac{1}{2}$, the left-hand side of (2.50) is a coercive continuous bilinear form on $H^{1,0}(\Omega_{0})$.
-On the other hand, the right-hand side of (2.50) defines a continuous linear functional $T$ on $H^{1,0}(\Omega_{0})$ with norm bounded by
-$\|T\|\leq C\|\Psi(\varepsilon,\beta)\|_{H^{1}(\Omega_{0})}\leq C^{\prime}\|f\|_{H^{\frac{1}{2}}}\ ,$
-where $C$ and $C^{\prime}$ are independent of $\varepsilon$.
-The Lax-Milgram theorem then yields the existence and uniqueness of a weak solution $u_{\varepsilon,\beta}$ to (2.48).
-Moreover, it satisfies $\|u_{\varepsilon,\beta}\|_{H^{1}(\Omega)}\leq C\|T\|$, and thus (2.49) holds for $s=\frac{1}{2}$.
-Now we fix $(\varepsilon,\beta)\in(-\varepsilon_{0}(s),\varepsilon_{0}(s))\times(0,\infty)$.
-For small $\mu$, the expression $v_{\mu}:=\frac{\Psi(\varepsilon+\mu,\beta)-\Psi(\varepsilon,\beta)}{\mu}-u_{\varepsilon,\beta}\in H^{1,0}(\Omega_{0})$ satisfies, in the weak sense analogously to (2.50),
-$\begin{cases}(\partial_{x}^{2}+\big{(}\frac{h}{h_{\varepsilon}}\big{)}^{2}\partial_{z}^{2})v_{\mu}-\beta\tilde{J}(\varepsilon+\mu)v_{\mu}=F_{\mu}\quad\text{in }\Omega_{0},\\ v_{\mu}(\cdot,0)=0,\quad\partial_{z}v_{\mu}(\cdot,-h)=0,\end{cases}$ (2.51)
-where
-$F_{\mu}$ $=\beta\left[\frac{\tilde{J}(\varepsilon+\mu)-\tilde{J}(\varepsilon)}{\mu}-\frac{d}{d\varepsilon}\tilde{J}(\varepsilon)\right]\Psi(\varepsilon,\beta)+\beta\left[\tilde{J}(\varepsilon+\mu)-\tilde{J}(\varepsilon)\right]u_{\varepsilon,\beta}$ (2.52)
-$\quad-\left[\frac{1}{\mu}\left(\Big{(}\frac{h}{h_{\varepsilon+\mu}}\Big{)}^{2}-\Big{(}\frac{h}{h_{\varepsilon}}\Big{)}^{2}\right)-\frac{d}{d\varepsilon}\big{(}\frac{h}{h_{\varepsilon}}\big{)}^{2}\right]\partial_{z}^{2}\Psi(\varepsilon,\beta)-\left(\Big{(}\frac{h}{h_{\varepsilon+\mu}}\Big{)}^{2}-\Big{(}\frac{h}{h_{\varepsilon}}\Big{)}^{2}\right)\partial_{z}^{2}u_{\varepsilon,\beta}.$
-As above, we first consider the case $s\geq\frac{3}{2}$. Then the uniform estimates (2.47) and (2.49) imply
-$\|F_{\mu}\|_{H^{s-\frac{3}{2}}(\Omega_{0})}\leq\mu C_{3}(\beta,s)\|f\|_{H^{s}(\mathbb{T})},\quad s\geq\frac{3}{2}.$ (2.53)
-Consequently, elliptic regularity theory yields
-$\|v_{\mu}\|_{H^{s+\frac{1}{2}}(\Omega_{0})}\leq\mu C_{4}(\beta,s)\|f\|_{H^{s}(\mathbb{T})},\quad s\geq\frac{3}{2}.$ (2.54)
-Now consider the case $s=\frac{1}{2}$.
-Since $\partial_{z}\Psi(\varepsilon,\beta)|_{z=0}=\partial_{z}u_{\varepsilon,\beta}|_{z=0}=0$, we can argue as in the proof of (2.49) for $s=\frac{1}{2}$ to obtain (2.54) for $s=\frac{1}{2}$.
-Of course, because the mappings
-$f\mapsto\Psi(\varepsilon,\beta)\mapsto u_{\varepsilon,\delta}\quad\text{and}\quad(\Psi(\varepsilon,\beta),u_{\varepsilon,\beta})\mapsto F_{\mu}\mapsto v_{\mu}$
-are linear, the mapping $f\mapsto v_{\mu}$ is also linear.
-Thus by interpolation, (2.54) also holds for all $s\geq\frac{1}{2}$.
-Letting $\mu\to 0$ in (2.54), we conclude that the mapping (2.46) is differentiable in $\varepsilon\in(-\varepsilon_{0}(s),\varepsilon_{0}(s))$ for fixed $\beta\in(0,\infty)$.
-An induction argument shows that this mapping is $C^{\infty}$ in $(\varepsilon,\beta)$.
-2.
-This step is devoted to the analyticity of $\Theta(\varepsilon,\beta)$.
-Fix an arbitrary number $\beta^{0}\in(0,\infty)$.
-We will prove the analyticity in a neighborhood of $\varepsilon=0$ and $\beta=\beta^{0}$.
-If we choose $\varepsilon_{0}(s)$ small enough, then $\tilde{J}(\varepsilon)$ is analytic in $\varepsilon$ with values in $H^{s+10}(\mathbb{T})$, say.
-Consequently, for any subinterval $I=[-a,a]\subset(-\varepsilon_{0}(s),\varepsilon_{0}(s))$, there exists a constant $M=M(I)>1$ such that for all $\varepsilon\in I$ we have the $\varepsilon$-expansions
-$\tilde{J}(x,z;\varepsilon)=\sum_{j=0}^{\infty}\varepsilon^{j}\chi_{j}(x,z),\quad\big{(}\frac{h}{h_{\varepsilon}}\big{)}^{2}=\sum_{j=0}^{\infty}\varepsilon^{j}\nu_{j}$
-with
-$\|\chi_{j}\|_{H^{s+10}(\Omega_{\varepsilon})}\leq M^{j+1},\quad|\nu_{j}|\leq M^{j+1}.$ (2.55)
-Moreover, $\chi_{0}=\nu_{0}=1$ and $\frac{h}{h_{\varepsilon}}\geq\frac{1}{2}$ if $\varepsilon_{0}(s)$ is small enough.
-Our goal in this part is to prove that $\Psi(\varepsilon,\beta)$ can be expanded in powers of $\varepsilon$ and $\beta-\beta^{0}$ as a convergent series
-$\Psi(\varepsilon,\beta)=\sum_{j=0}^{\infty}\sum_{\ell=0}^{\infty}\varepsilon^{j}(\beta-\beta^{0})^{\ell}\Psi_{j,\ell}\quad\text{in }H^{s+\frac{1}{2}}(\Omega_{0})$ (2.56)
-near $(\varepsilon,\beta)=(0,\beta^{0})$.
-We begin formally. From (2.44) the coefficients $\Psi_{j,\ell}$ must formally satisfy the following problems. First, $\Psi_{0,0}$ satisfies
-$\begin{cases}\Delta_{x,z}\Psi_{0,0}-\beta^{0}\Psi_{0,0}=0\quad\text{in~{}}\Omega_{0},\\ \Psi_{0,0}(x,0)=f(x),\quad\partial_{z}\Psi_{0,0}(x,-h)=0.\end{cases}$ (2.57)
-Next, for $\ell=0$ and $j\geq 1$, $\Psi_{j,0}$ satisfies
-$\begin{cases}\Delta_{x,z}\Psi_{j,0}-\beta^{0}\Psi_{j,0}=\beta^{0}\sum_{k=0}^{j-1}\chi_{j-k}\Psi_{k,0}-\sum_{k=0}^{j-1}\nu_{j-k}\partial_{z}^{2}\Psi_{k,0}\quad\text{in~{}}\Omega_{0},\\ \Psi_{j,0}(\cdot,0)=0,\quad\partial_{z}\Psi_{j,0}(\cdot,-h)=0.\end{cases}$ (2.58)
-Finally, for $\ell\geq 1$ and $j\geq 1$, $\Psi_{j,\ell}$ satisfies
-$\begin{cases}\Delta_{x,z}\Psi_{j,\ell}-\beta^{0}\Psi_{j,\ell}=\sum_{k=0}^{j}\chi_{j-k}\Psi_{k,\ell-1}+\beta^{0}\sum_{k=0}^{j-1}\chi_{j-k}\Psi_{k,\ell}-\sum_{k=0}^{j-1}\nu_{j-k}\partial_{z}^{2}\Psi_{k,\ell}\quad\text{in~{}}\Omega_{0},\\ \Psi_{j,\ell}(\cdot,0)=0,\quad\partial_{z}\Psi_{j,\ell}(\cdot,-h)=0\end{cases}$ (2.59)
-with the convention that $\sum_{k=0}^{j-1}=0$ when $j=0$.
-Now let us estimate $\Psi_{j,\ell}$ by elliptic theory. First, from (2.57) we have
-$\|\Psi_{0,0}\|_{H^{s+\frac{1}{2}}(\Omega_{0})}\leq M_{0}\|f\|_{H^{s}(\mathbb{T})},\quad M_{0}=M_{0}(\beta^{0},s,h).$ (2.60)
-Next, for any $s\geq\frac{1}{2}$ the problems (2.57), (2.58) and (2.59) should be understood in the weak sense analogous to (2.50).
-Precisely, $\Psi_{0,0}$ is the unique function in $H^{1}(\Omega)$ satisfying $\Psi_{0,0}|_{z=0}=f$ and
-$-\int_{\Omega_{0}}\partial_{x}\Psi_{0,0}\cdot\partial_{x}\varphi+\partial_{z}\Psi_{0,0}\partial_{z}\varphi+\beta^{0}\Psi_{0,0}\varphi=0\quad\forall\varphi\in H^{1,0}(\Omega_{0}),$ (2.61)
-$\Psi_{j,0}$ ($j\geq 1$) is the unique function in $H^{1,0}(\Omega_{0})$ satisfying
-$-\int_{\Omega_{0}}\partial_{x}\Psi_{j,0}\cdot\partial_{x}\varphi+\partial_{z}\Psi_{j,0}\partial_{z}\varphi+\beta^{0}\Psi_{j,0}\varphi$ (2.62)
-$\qquad=\int_{\Omega_{0}}\beta^{0}\sum_{k=0}^{j-1}\chi_{j-k}\Psi_{k,0}\varphi+\int_{\Omega_{0}}\sum_{k=0}^{j-1}\nu_{j-k}\partial_{z}\Psi_{k,0}\partial_{z}\varphi\quad\forall\varphi\in H^{1,0}(\Omega_{0}),$
-and $\Psi_{j,\ell}$ ($\ell\geq 1$) is the unique function in $H^{1,0}(\Omega_{0})$ satisfying
-$-\int_{\Omega_{0}}\partial_{x}\Psi_{j,\ell}\cdot\partial_{x}\varphi+\partial_{z}\Psi_{j,\ell}\partial_{z}\varphi+\beta^{0}\Psi_{j,\ell}\varphi$ (2.63)
-$\qquad=\int_{\Omega_{0}}\left(\sum_{k=0}^{j}\chi_{j-k}\Psi_{k,\ell-1}+\beta^{0}\sum_{k=0}^{j-1}\chi_{j-k}\Psi_{k,\ell}\right)\varphi+\int_{\Omega_{0}}\sum_{k=0}^{j-1}\nu_{j-k}\partial_{z}\Psi_{k,\ell}\partial_{z}\varphi\quad\forall\varphi\in H^{1,0}(\Omega_{0}).$
-By product estimates in Sobolev spaces, there exists a constant $M_{1}=M_{1}(\beta^{0},s,h)>1$ such that
-$\|\Psi_{j,0}\|_{H^{s+\frac{1}{2}}(\Omega_{0})}\leq M_{1}\sum_{k=0}^{j-1}\big{(}\beta^{0}\|\chi_{j-k}\|_{H^{s+10}(\Omega_{0})}+|\nu_{j-k}|\big{)}\|\Psi_{k,0}\|_{H^{s+\frac{1}{2}}(\Omega_{0})}\quad\text{for~{}}j\geq 1$ (2.64)
-and
-$\|\Psi_{j,\ell}\|_{H^{s+\frac{1}{2}}(\Omega_{0})}$ $\leq M_{1}\sum_{k=0}^{j}\|\chi_{j-k}\|_{H^{s+10}(\Omega_{0})}\|\Psi_{k,\ell-1}\|_{H^{s+\frac{1}{2}}(\Omega_{0})}$ (2.65)
-$\quad+M_{1}\sum_{k=0}^{j-1}\big{(}\beta^{0}\|\chi_{j-k}\|_{H^{s+10}(\Omega_{0})}+|\nu_{j-k}|\big{)}\|\Psi_{k,\ell}\|_{{}_{H^{s+\frac{1}{2}}(\Omega_{0})}}\quad\text{for~{}}\ell\geq 1.$
-Lemma 2.8.
-Choosing any number $A$ that satisfies
-$A>2+M+M_{1}+(\beta^{0}+1)M_{1}M\ ,$ (2.66)
-we have
-$\|\Psi_{j,\ell}\|_{H^{s+\frac{1}{2}}(\Omega_{0})}\leq\|\Psi_{0,0}\|_{H^{s+\frac{1}{2}}(\Omega_{0})}(AM)^{j+\ell}\quad\text{if}\quad j+\ell\geq 1.$ (2.67)
-Proof.
-We will prove the claim (2.67) by induction on $\ell$ and $j$. We begin with $\ell=0$ and $j=1$. Combining (2.64), (2.55), and (2.66) yields
-$\|\Psi_{1,0}\|_{H^{s+\frac{1}{2}}(\Omega_{0})}$ $\leq M_{1}(\beta^{0}\|\chi_{0}\|_{H^{s+10}(\Omega_{0})}+|\nu_{0}|)\|\Psi_{0,0}\|_{H^{s+\frac{1}{2}}(\Omega_{0})}$
-$\leq\|\Psi_{0,0}\|_{H^{s+\frac{1}{2}}(\Omega_{0})}(\beta^{0}+1)M_{1}M\leq\|\Psi_{0,0}\|_{H^{s+\frac{1}{2}}(\Omega_{0})}(AM)\quad(M>1).$
-This proves (2.67) for $(j,\ell)=(1,0)$. Letting $\ell=0$ and $j\geq 1$, we assume by induction on $j$ that (2.67) is valid up to $j-1$. Then we deduce from (2.64), (2.55), and (2.66) that
-$\|\Psi_{j,0}\|_{H^{s+\frac{1}{2}}(\Omega_{0})}$ $\leq M_{1}\sum_{k=0}^{j-1}\big{(}\beta^{0}M^{j-k+1}+M^{j-k+1}\big{)}\|\Psi_{0,0}\|_{H^{s+\frac{1}{2}}(\Omega_{0})}(AM)^{k}$
-$=\|\Psi_{0,0}\|_{H^{s+\frac{1}{2}}(\Omega_{0})}(AM)^{j}(\beta^{0}+1)M_{1}M\frac{1-A^{-j}}{A-1}$
-$\leq\|\Psi_{0,0}\|_{H^{s+\frac{1}{2}}(\Omega_{0})}(AM)^{j}.$
-Thus (2.67) holds for $\ell=0$ and $j\geq 1$.
-Next, we consider $\ell=1$.
-The case $j=0$ follows directly from (2.65) and the fact that $A>M_{1}$.
-Assume by induction on $j$ that (2.67) holds up to $j-1$, for $j\geq 2$.
-In view of (2.65), (2.55), (2.66), and the case $\ell=0$, we have
-$\|\Psi_{j,1}\|_{H^{s+\frac{1}{2}}(\Omega_{0})}$ $\leq M_{1}\sum_{k=0}^{j}\|\chi_{j-k}\|_{H^{s+10}(\Omega_{0})}\|\Psi_{k,0}\|_{H^{s+\frac{1}{2}}(\Omega_{0})}$
-$\quad+M_{1}\sum_{k=0}^{j-1}\big{(}\beta^{0}\|\chi_{j-k}\|_{H^{s+10}(\Omega_{0})}+|\nu_{j-k}|\big{)}\|\Psi_{k,1}\|_{{}_{H^{s+\frac{1}{2}}(\Omega_{0})}}$
-$\leq M_{1}\sum_{k=0}^{j}M^{j-k+1}(AM)^{k}\|\Psi_{0,0}\|_{H^{s+\frac{1}{2}}(\Omega_{0})}$
-$\quad+M_{1}\sum_{k=0}^{j-1}(\beta^{0}+1)M^{j-k+1}(AM)^{k+1}\|\Psi_{0,0}\|_{H^{s+\frac{1}{2}}(\Omega_{0})}\quad(\text{by induction})$
-$=M_{1}M^{j+1}\|\Psi_{0,0}\|_{H^{s+\frac{1}{2}}(\Omega_{0})}\frac{A^{j+1}-1}{A-1}+(\beta^{0}+1)M_{1}M^{j+2}\frac{A^{j+1}-A}{A-1}$
-$\leq(AM)^{j+1}\|\Psi_{0,0}\|_{H^{s+\frac{1}{2}}(\Omega_{0})}\big{(}M_{1}+(\beta^{0}+1)M_{1}M\big{)}\frac{1}{A-1}$
-$\leq(AM)^{j+1}\|\Psi_{0,0}\|_{H^{s+\frac{1}{2}}(\Omega_{0})}.$
-This proves (2.67) for $\ell=1$ and $j\geq 0$. The same reasoning allows us to prove by induction on $\ell$ that (2.67) holds for all $\ell\geq 1$. ∎
-Due to (2.67) and (2.60), it follows that the series (2.56) converges in $H^{s+\frac{1}{2}}(\Omega_{0})$ provided $|\varepsilon|,~{}|\beta-\beta^{0}|<(AM)^{-1}$.
-Denote the sum of this series by $\tilde{\Psi}\in H^{s+\frac{1}{2}}(\Omega_{0})$.
-To conclude the proof of (2.56), it suffices show that $\tilde{\Psi}=\Psi$.
-For this purpose it suffices to consider the case $s=\frac{1}{2}$.
-Define the partial sums
-$\Psi^{(m,n)}=\sum_{j=0}^{m}\sum_{\ell=0}^{n}\varepsilon^{j}(\beta-\beta^{0})^{\ell}\Psi_{j,\ell},\quad\chi^{(m)}=\sum_{j=0}^{m}\varepsilon^{j}\chi_{j},\quad H^{(m)}=\sum_{j=0}^{m}\varepsilon^{j}\nu_{j}.$ (2.68)
-Using the fact that $\chi_{0}=\nu_{0}=1$, we deduce from the weak equations (2.61), (2.62) and (2.63) that
-$-\int_{\Omega_{0}}\partial_{x}\Psi^{(m,n)}\partial_{x}\varphi=\int_{\Omega_{0}}(\beta-\beta^{0})\chi^{(m)}\Psi^{(m,n-1)}\varphi+\beta^{0}\chi^{(m)}\Psi^{(m,n)}\varphi+H^{(m)}\Psi^{(m,n)}\partial_{z}\varphi\quad\ \forall\varphi\in H^{1,0}(\Omega_{0}).$ (2.69)
-Indeed, using the convention that $\sum_{k=0}^{j-1}=0$ for $j=0$, we have for all $\varphi\in H^{1,0}(\Omega_{0})$ that
-$-\int_{\Omega}\partial_{x}\Psi^{(m,n)}\partial_{x}\varphi$ $=\int_{\Omega}\partial_{z}\Psi^{(m,n)}\partial_{z}\varphi+\int_{\Omega}\beta^{0}\Psi^{(m,n)}\varphi$
-$\quad+\beta^{0}\int_{\Omega}\sum_{j=0}^{m}\varepsilon^{j}\sum_{k=0}^{j-1}\chi_{j-k}\Psi_{k,0}\varphi+\int_{\Omega}\sum_{j=0}^{m}\varepsilon^{j}\sum_{k=0}^{j-1}\nu_{j-k}\partial_{z}\Psi_{k,0}\partial_{z}\varphi$
-$\quad+\int_{\Omega}\sum_{j=0}^{m}\sum_{\ell=1}^{n}\varepsilon^{j}(\beta-\beta^{0})^{\ell}\sum_{k=0}^{j}\chi_{j-k}\Psi_{k,\ell-1}\varphi$
-$\quad+\beta^{0}\int_{\Omega}\sum_{j=0}^{m}\sum_{\ell=1}^{n}\varepsilon^{j}(\beta-\beta^{0})^{\ell}\sum_{k=0}^{j-1}\chi_{j-k}\Psi_{k,\ell}\varphi+\int_{\Omega}\sum_{j=0}^{m}\sum_{\ell=1}^{n}\varepsilon^{j}(\beta-\beta^{0})^{\ell}\sum_{k=0}^{j-1}\nu_{j-k}\partial_{z}\Psi_{k,\ell}\partial_{z}\varphi$
-$=\int_{\Omega}\partial_{z}\Psi^{(m,n)}\partial_{z}\varphi+\int_{\Omega}\beta^{0}\Psi^{(m,n)}\varphi$
-$\quad+\int_{\Omega}\sum_{j=0}^{m}\sum_{\ell=1}^{n}\varepsilon^{j}(\beta-\beta^{0})^{\ell}\sum_{k=0}^{j}\chi_{j-k}\Psi_{k,\ell-1}\varphi$
-$\quad+\beta^{0}\int_{\Omega}\sum_{j=0}^{m}\sum_{\ell=0}^{n}\varepsilon^{j}(\beta-\beta^{0})^{\ell}\sum_{k=0}^{j-1}\chi_{j-k}\Psi_{k,\ell}\varphi+\int_{\Omega}\sum_{j=0}^{m}\sum_{\ell=0}^{n}\varepsilon^{j}(\beta-\beta^{0})^{\ell}\sum_{k=0}^{j-1}\nu_{j-k}\partial_{z}\Psi_{k,\ell}\partial_{z}\varphi$
-$=(\beta-\beta^{0})\quad\int_{\Omega}\sum_{j=0}^{m}\sum_{\ell^{\prime}=1}^{n-1}\varepsilon^{j}(\beta-\beta^{0})^{\ell^{\prime}}\sum_{k=0}^{j}\chi_{j-k}\Psi_{k,\ell^{\prime}}\varphi$
-$\quad+\beta^{0}\int_{\Omega}\sum_{j=0}^{m}\sum_{\ell=0}^{n}\varepsilon^{j}(\beta-\beta^{0})^{\ell}\sum_{k=0}^{j}\chi_{j-k}\Psi_{k,\ell}\varphi+\int_{\Omega}\sum_{j=0}^{m}\sum_{\ell=0}^{n}\varepsilon^{j}(\beta-\beta^{0})^{\ell}\sum_{k=0}^{j}\nu_{j-k}\partial_{z}\Psi_{k,\ell}\partial_{z}\varphi$
-$=\int_{\Omega_{0}}(\beta-\beta^{0})\chi^{(m)}\Psi^{(m,n-1)}\varphi+\beta^{0}\chi^{(m)}\Psi^{(m,n)}\varphi+H^{(m)}\Psi^{(m,n)}\partial_{z}\varphi,$
-which is (2.69). Letting $m,~{}n\to\infty$ in (2.69) yields
-$-\int_{\Omega_{0}}\partial_{x}\tilde{\Psi}\partial_{x}\varphi=\int_{\Omega_{0}}(\beta-\beta^{0})\tilde{J}\tilde{\Psi}\varphi+\beta^{0}\tilde{J}\tilde{\Psi}\varphi+\big{(}\frac{h}{h_{\varepsilon}}\big{)}^{2}\partial_{z}\Psi\partial_{z}\varphi\quad\forall\varphi\in H^{1,0}(\Omega_{0}),$
-and hence
-$\int_{\Omega_{0}}\partial_{x}\tilde{\Psi}\partial_{x}\varphi+\big{(}\frac{h}{h_{\varepsilon}}\big{)}^{2}\partial_{z}\Psi\partial_{z}\varphi=\int_{\Omega_{0}}\beta\tilde{J}\tilde{\Psi}\varphi\quad\forall\varphi\in H^{1,0}(\Omega_{0}).$ (2.70)
-By the continuity from $H^{1}(\Omega_{0})$ to $H^{\frac{1}{2}}(\mathbb{T})$ of the trace $\Psi_{j,\ell}\mapsto\Psi_{j,\ell}|_{z=0}$, we also have $\Psi|_{z=0}=\Psi_{0,0}|_{z=0}=f$.
-This and (2.70) imply that $\tilde{\Psi}$ is a weak solution of (2.44).
-Thus $\tilde{\Psi}=\Psi$ by the uniqueness of (2.44).
-This completes the proof of (2.56).
-3.
-Now we make the connection to $\mathcal{G}_{\varepsilon,\beta}$, which is given by (2.45).
-By abuse of notation we may regard $\Psi(\varepsilon,\beta)$ and $\Psi_{j,\ell}$ as linear operators acting on $f$.
-It follows from (2.67) and (2.60) that
-$\forall s\geq\frac{1}{2},\quad\|\Psi_{j,\ell}\|_{H^{s+\frac{1}{2}}(\Omega_{0})}\leq M_{0}(AM)^{j+\ell}\|f\|_{H^{s}(\mathbb{T})},$ (2.71)
-where $M_{0}$, $M$ and $A$ depend only on $(\beta^{0},s,h)$.
-In the smoother case $s>1$, we combine the analyticity (2.56) of $\Psi(\varepsilon,\beta)$ in $H^{s+\frac{1}{2}}(\Omega_{0})$ with the standard continuity from $H^{s-\frac{1}{2}}(\Omega_{0})$ to $H^{s-1}(\mathbb{T})$ of the trace of $\partial_{z}\Psi(\varepsilon,\beta)$ and $\partial_{z}\Psi_{j,\ell}$ onto $\{z=0\}$, to obtain
-$\tfrac{h_{\varepsilon}}{h}\mathcal{G}_{\varepsilon,\beta}=\partial_{z}\Psi(\varepsilon,\beta)|_{z=0}=\sum_{j=0}^{\infty}\sum_{\ell=0}^{\infty}\varepsilon^{j}(\beta-\beta^{0})^{\ell}\partial_{z}\Psi_{j,\ell}|_{z=0}\quad\text{in }H^{s-1}(\mathbb{T}).$
-Thus the mapping (2.35) is analytic. Because the trace theorem fails for $s\leq 1$, we shall provide a different argument for $s=\frac{1}{2}$ and then conclude the proof by interpolation.
-Now we fix $s=\frac{1}{2}$.
-Then the coefficient $\Psi_{j,\ell}$ belongs to $H^{1}(\Omega_{0})$ and satisfies the bound (2.67).
-First, since $\Psi_{0,0}=\Theta|_{(\varepsilon,\beta)=(0,\beta^{0})}$, in view of the explicit formula (2.43), we have $\partial_{z}\Psi_{0,0}|_{z=0}\in H^{-\frac{1}{2}}(\mathbb{T})$ and $\|\partial_{z}\Psi_{0,0}|_{z=0}\|_{H^{-\frac{1}{2}}}\leq C\|f\|_{H^{\frac{1}{2}}}$.
-Next, for general $j$ and $k$ we will appeal to Lemma A.2 to make sense of $\partial_{z}\Psi_{j,k}|_{z=0}$.
-For brevity, we define
-$w_{j,\ell}:=\sum_{k=0}^{j}\nu_{j-k}\Psi_{k,\ell}.$
-In case $\ell=0$ and $j\geq 1$, we add $\Delta_{x,z}\sum_{k=0}^{j-1}\nu_{j-k}\Psi_{k,0}$ to both sides of (2.58) and use $\nu_{0}=\chi_{0}=1$ to obtain
-$\Delta_{x,z}w_{j,0}=\beta^{0}\sum_{k=0}^{j}\chi_{j-k}\Psi_{k,0}+\partial_{x}^{2}\sum_{k=0}^{j-1}\nu_{j-k}\Psi_{k,0}\quad\text{in~{}}\Omega_{0}.$ (2.72)
-More generally, for $\ell\geq 1$ we add $\Delta_{x,z}\sum_{k=0}^{j-1}\nu_{j-k}\Psi_{k,\ell}$ to both sides of (2.59) to obtain
-$\Delta_{x,z}w_{j,\ell}=\sum_{k=0}^{j}\chi_{j-k}\Psi_{k,\ell-1}+\beta^{0}\sum_{k=0}^{j}\chi_{j-k}\Psi_{k,\ell}+\partial_{x}^{2}\sum_{k=0}^{j-1}\nu_{j-k}\Psi_{k,\ell}\quad\text{in~{}}\Omega_{0}.$ (2.73)
-Hence, applying the estimate (A.7) in Lemma A.2 to $u=\nabla w_{j,\ell}$, we have $\tilde{\gamma}(\nabla w_{j,\ell})\in(H^{-\frac{1}{2}}(\mathbb{T}))^{2}$ and the estimate
-$\|\tilde{\gamma}(\nabla w_{j,\ell})\|_{(H^{-\frac{1}{2}}(\mathbb{T}))^{2}}$ $\leq C\|\nabla w_{j,\ell}\|_{L^{2}(\Omega_{0})}+C\left\|\sum_{k=0}^{j}\chi_{j-k}\Psi_{k,\ell-1}+\beta^{0}\sum_{k=0}^{j}\chi_{j-k}\Psi_{k,\ell}\right\|_{L^{2}(\Omega_{0})}$
-$\qquad+C\left\|\partial_{x}\sum_{k=0}^{j-1}\nu_{j-k}\Psi_{k,\ell}\right\|_{L^{2}(\Omega_{0})}$
-with the convention that $\Psi_{j,-1}=0$, where $C$ depends only on $h$.
-Then, using the definition of $w_{j,\ell}$ and invoking the estimates (2.67) and (2.60) for $\|\Psi_{j,\ell}\|_{H^{1}(\Omega_{0})}$, as well as the bounds (2.55) for $\chi_{j-k}$ and $\nu_{j-k}$, we find
-$\|\tilde{\gamma}(\nabla w_{j,\ell})\|_{(H^{-\frac{1}{2}}(\mathbb{T}))^{2}}\leq\frac{C_{1}}{A-1}\|f\|_{H^{\frac{1}{2}}(\mathbb{T})}(AM)^{j+\ell},\quad j,~{}\ell\geq 0,$ (2.74)
-recalling that $A>1$ and $M>1$, where $C_{1}$ depends only $(\beta^{0},h)$.
-From (2.60) we have $\|\Psi_{0,0}\|_{H^{1}(\Omega_{0})}\leq M_{0}\|f\|_{H^{\frac{1}{2}}(\mathbb{T})}$, $M_{0}=M_{0}(\beta^{0},h)$.
-Then, the first equation in (2.57) implies $\Delta\Psi_{0,0}\in H^{1}(\Omega_{0})\subset L^{2}(\Omega_{0})$, so that $\nabla\Psi_{0,0}\in H_{\operatorname{div}}(\Omega_{0})$.
-Therefore, we can apply Proposition A.1 and Lemma A.2 to have $\tilde{\gamma}(\nabla\Psi_{0,0})={\gamma}(\nabla\Psi_{0,0})$ and $\|\tilde{\gamma}(\nabla\Psi_{0,0})\|_{H^{-\frac{1}{2}}(\mathbb{T})}\leq C_{2}\|f\|_{H^{\frac{1}{2}}(\mathbb{T})}$ for some constant $C_{2}=C_{2}(\beta^{0},h)\geq C_{1}$.
-Now we prove by induction on $j+\ell$ that
-$\|\tilde{\gamma}(\nabla\Psi_{j,\ell})\|_{(H^{-\frac{1}{2}}(\mathbb{T}))^{2}}\leq C_{2}\|f\|_{H^{\frac{1}{2}}}(AM)^{j+\ell},\quad j,~{}\ell\geq 0.$ (2.75)
-Indeed, the case $j+\ell=0$ is true.
-Suppose by induction that (2.75) holds for all pairs $(j,\ell)$ with $j+\ell\leq L-1$, and consider $j+\ell=L$.
-By the definition of $w_{j,\ell}$ we have $\Psi_{j,\ell}=w_{j,\ell}-\sum_{k=0}^{j-1}\nu_{j-k}\Psi_{k,\ell}$, so that (2.74) and the induction hypothesis yield
-$\|\tilde{\gamma}(\nabla\Psi_{j,\ell})\|_{(H^{-\frac{1}{2}}(\mathbb{T}))^{2}}$ $\leq\|\tilde{\gamma}(\nabla w_{j,\ell})\|_{(H^{-\frac{1}{2}}(\mathbb{T}))^{2}}+\sum_{k=0}^{j-1}|\nu_{j-k}\|\tilde{\gamma}(\nabla\Psi_{k,\ell})\|_{(H^{-\frac{1}{2}}(\mathbb{T}))^{2}}$
-$\leq\frac{C_{2}}{A-1}\|f\|_{H^{\frac{1}{2}}(\mathbb{T})}(AM)^{j+\ell}+\sum_{k=0}^{j-1}M^{j-k+1}C_{2}\|f\|_{H^{\frac{1}{2}}}(AM)^{k+\ell}$
-$\leq\frac{C_{2}}{A-1}\|f\|_{H^{\frac{1}{2}}(\mathbb{T})}(AM)^{j+\ell}+\frac{C_{2}M}{A-1}\|f\|_{H^{\frac{1}{2}}}(AM)^{j+\ell}$
-$\leq C_{2}\|f\|_{H^{\frac{1}{2}}}(AM)^{j+\ell},$
-where in the last inequality we have used the fact from (2.66) that $A-1>1+M$. The proof of (2.75) is complete.
-Combining (2.56) with (2.75), we deduce that
-the series $\text{the series ~{}}\sum_{j=0}^{\infty}\sum_{\ell=0}^{\infty}\varepsilon^{j}(\beta-\beta^{0})^{\ell}\tilde{\gamma}(\nabla\Psi_{j,\ell})\text{~{}converges to some~{}}g~{}\text{in~{}}(H^{-\frac{1}{2}}(\mathbb{T}))^{2}$ (2.76)
-provided $|\varepsilon|<(AM)^{-1}$ and $|\beta-\beta^{0}|<(AM)^{-1}$. We will show that $g=\tilde{\gamma}(\nabla\Psi(\varepsilon,\beta))$. To that end we define
-$u_{m,n}:=\begin{bmatrix}\partial_{x}\Psi^{(m,n)}\\ H^{(m)}\partial_{z}\Psi^{(m,n)}\end{bmatrix}=H^{(m)}\nabla\Psi^{(m,n)}+\begin{bmatrix}(1-H^{(m)})\partial_{x}\Psi^{(m,n)}\\ 0\end{bmatrix},$ (2.77)
-where $H^{(m)}$ is given in (2.68). Combining Lemma A.2 (i)(c) with (2.76), we obtain
-[…]
-On the other hand, by the first equations in (2.57), (2.58) and (2.59), we find
-$\operatorname{div}u_{m,n}=\partial_{x}^{2}\Psi^{(m,n)}+H^{(m)}\partial_{z}^{2}\Psi^{(m,n)}=\beta^{0}\chi^{(m)}\Psi^{(m,n)}+(\beta-\beta^{0})\chi^{(m)}\Psi^{(m,n-1)}\in H^{1}(\Omega_{0}),$
-which can be proven similarly to (2.69). In addition,
-$u:=\begin{bmatrix}\partial_{x}\Psi(\varepsilon,\beta)\\ \big{(}\frac{h}{h_{\varepsilon}}\big{)}^{2}\partial_{z}\Psi(\varepsilon,\beta)\end{bmatrix}$
-satisfies $\operatorname{div}u=\beta\tilde{J}\Psi(\varepsilon,\beta)\in H^{1}(\Omega_{0})$ in light of (2.44).
-It follows that $u_{m,n},\ u\in H_{\operatorname{div}}(\Omega_{0})$, and hence (A.2) implies
-$\|\gamma(u_{m,n}-u)\|_{(H^{-\frac{1}{2}}(\mathbb{T}))^{2}}\leq C\|u_{m,n}-u\|_{L^{2}(\Omega_{0})}+C\|\operatorname{div}(u_{m,n}-u)\|_{L^{2}(\Omega_{0})}\to 0\quad\text{as~{}}m,n\to\infty,$
-where we have used the fact from (2.56) that $\Psi^{(m,n)}\to\Psi(\varepsilon,\beta)$ in $H^{1}(\Omega_{0})$.
-Recalling our definition (A.6) of $\tilde{\gamma}$ on $H_{\operatorname{div}}+\tilde{H}_{\operatorname{div}}$, we obtain
-$\lim_{m,n\to\infty}\tilde{\gamma}(u_{m,n})=\tilde{\gamma}(u)=\big{(}\frac{h}{h_{\varepsilon}}\big{)}^{2}\tilde{\gamma}(\nabla\Psi(\varepsilon,\beta))\quad\text{in~{}}(H^{-\frac{1}{2}}(\mathbb{T}))^{2},$
-where the second equality follows from a decomposition analogous to (2.77).
-In conjunction with (2.78) this yields $g=\tilde{\gamma}(\nabla\Psi(\varepsilon,\beta))$ as claimed. It means that
-$\tilde{\gamma}(\nabla\Psi(\varepsilon,\beta))=\sum_{j=0}^{\infty}\sum_{\ell=0}^{\infty}\varepsilon^{j}(\beta-\beta^{0})^{\ell}\tilde{\gamma}(\nabla\Psi_{j,\ell})\quad\text{in~{}}(H^{-\frac{1}{2}}(\mathbb{T}))^{2}.$ (2.79)
-Due to (2.45), this yields the analyticity of the mapping (2.35) when $s=\frac{1}{2}$.
-Finally, in order to generalize to $s\geq\frac{1}{2}$, we shall interpolate.
-We recall (2.71) and view $M_{0}$, $A$ and $M$ as functions of $s\geq\frac{1}{2}$.
-We fix any $s_{0}>1$ and use the standard trace theorem to obtain
-$\|\gamma(\nabla\Psi_{j,\ell})\|_{(H^{s_{0}-1}(\mathbb{T}))^{2}}\leq N(s_{0})\|\Psi_{j,\ell}\|_{H^{s_{0}+\frac{1}{2}}(\Omega_{0})}\leq N(s_{0})M_{0}(s_{0})\big{(}A(s_{0})M(s_{0})\big{)}^{j+\ell}\|f\|_{H^{s_{0}}(\mathbb{T})}.$
-Thus we can view $\gamma(\nabla\Psi_{j,\ell})=\tilde{\gamma}(\nabla\Psi_{j,\ell})$ as an operator in $\mathcal{L}(H^{s_{0}}(\mathbb{T}),H^{s_{0}-1}(\mathbb{T}))^{2})$.
-In light of this inequality and (2.75), we obtain by interpolation that
-$\|\tilde{\gamma}(\nabla\Psi_{j,\ell})\|_{\mathcal{L}(H^{s}(\mathbb{T}),(H^{s-1}(\mathbb{T}))^{2})}\leq C_{2}^{\kappa}\big{(}A(\tfrac{1}{2})M(\tfrac{1}{2})\big{)}^{\kappa(j+\ell)}\big{(}N_{0}(s_{0})M_{0}(s_{0})\big{)}^{1-\kappa}\big{(}A(s_{0})M(s_{0})\big{)}^{(1-\kappa)(j+\ell)}$
-for any $s=\kappa\frac{1}{2}+(1-\kappa)s_{0}\in(\frac{1}{2},s_{0})$, $\kappa\in(0,1)$.
-Consequently the series $\sum_{j=0}^{\infty}\sum_{\ell=0}^{\infty}\varepsilon^{j}(\beta-\beta^{0})^{\ell}\tilde{\gamma}(\nabla\Psi_{j,\ell})$ converges in $(H^{s-1}(\mathbb{T}))^{2})$ provided $|\varepsilon|$ and $|\beta-\beta^{0}|$ are small enough.
-The sum of this series must equal $\tilde{\gamma}(\nabla\Psi(\varepsilon,\beta))$ thanks to (2.79).
-Therefore, the mapping (2.35) is analytic in a neighborhood of $\varepsilon=0$ and $\beta=\beta^{0}$ for any $s\geq\frac{1}{2}$.
-The proof of Theorem 2.6 is now complete.
+$\begin{cases}\left(\partial_{x}^{2}+\big(\frac{h}{h_{\varepsilon}}\big)^{2}\partial_{z}^{2}\right)\Psi-\beta\tilde{J}\Psi=0\quad\text{ in }\Omega_{0}\ ,\\ \Psi(x,0)=f(x),\quad\partial_{z}\Psi(x,-h)=0.\end{cases}$ (2.53)
+Use of the chain rule in the definition (2.41) of $\mathcal{G}_{\varepsilon,\beta}$ yields
+$\tfrac{h_{\varepsilon}}{h}\mathcal{G}_{\varepsilon,\beta}f=\partial_{z}\Psi(\varepsilon,\beta)|_{z=0}.$ (2.54)
+We rewrite (2.53) as
+$\begin{cases}\Delta_{x,z}\Psi-\beta_{0}\Psi=(\beta\tilde{J}-\beta_{0})\Psi+\Big[1-\big(\frac{h}{h_{\varepsilon}}\big)^{2}\Big]\partial_{z}^{2}\Psi\quad\text{ in }\Omega_{0}\ ,\\ \Psi(x,0)=f(x),\quad\partial_{z}\Psi(x,-h)=0.\end{cases}$ (2.55)
+Denoting $F_{0}=F_{0}(\Psi,\varepsilon,\beta)=(\beta\tilde{J}-\beta_{0})\Psi$ and
+$G=G(\Psi,\varepsilon)=\begin{bmatrix}0\\ \big[1-\big(\frac{h}{h_{\varepsilon}}\big)^{2}\big]\partial_{z}\Psi\end{bmatrix}$,
+we rewrite (2.55) in the form of (2.47):
+$\begin{cases}\Delta_{x,z}\Psi-\beta_{0}\Psi=F_{0}+\operatorname{div}G\quad\text{ in }\Omega_{0}\ ,\\ \Psi(x,0)=0,\quad\partial_{z}\Psi(x,-h)=G_{2},\end{cases}$ (2.56)
+where the Neumann condition $\partial_{z}\Psi(x,-h)=G_{2}=\big[1-\big(\frac{h}{h_{\varepsilon}}\big)^{2}\big]\partial_{z}\Psi$ is equivalent to $\partial_{z}\Psi(x,-h)=0$.
+Hence, in terms of the analytic operator $\mathcal{S}$ (2.51), we have
+$\Psi=\mathcal{S}(F_{0}(\Psi,\varepsilon,\beta),G(\Psi,\varepsilon),f).$
+For fixed $f\in H^{s}(\mathbb{T})$, we denote
+$\mathcal{T}(\Psi,\varepsilon,\beta)=\Psi-\mathcal{S}(F_{0}(\Psi,\varepsilon,\beta),G(\Psi,\varepsilon),f).$
+Both $F_{0}$ and $G$ are linear in $\Psi$ and analytic in $(\varepsilon,\beta)\in\mathcal{O}_{s}$, a neighborhood of $(0,\beta_{0})$.
+Together with the analyticity (2.52) of $\mathcal{S}$, we deduce that $\mathcal{T}:H^{s+\frac{1}{2}}(\Omega_{0})\times\mathcal{O}_{s}\to H^{s+\frac{1}{2}}(\Omega_{0})$ is analytic.
+Set $\Psi_{0}=\mathcal{S}(0,0,f)$.
+Since $F_{0}(\cdot,0,\beta_{0})\equiv 0$ and $G(\cdot,0)\equiv 0$, we have $\mathcal{T}(\Psi_{0},0,\beta_{0})=0$ and $D_{1}\mathcal{T}(\Psi_{0},0,\beta_{0})=\mathrm{I}$.
+Therefore, by virtue of the analytic implicit function theorem, we deduce that
+$\Psi=\Psi(\varepsilon,\beta):\mathcal{O}_{s}\ni(\varepsilon,\beta)\to H^{s+\frac{1}{2}}(\Omega_{0})\quad\text{is~analytic},$ (2.57)
+where $\mathcal{O}_{s}$ can be shrunk if necessary.
+Next, we take the trace $\partial_{z}\Psi(\varepsilon,\beta)|_{z=0}$ to make the connection to $\mathcal{G}_{\varepsilon,\beta}f=\frac{h}{h_{\varepsilon}}\partial_{z}\Psi(\varepsilon,\delta)|_{z=0}$.
+If $s>1$, then combining (2.57) (and a similar property for $\Psi_{2}$) with the continuity of the standard trace operator $H^{s-\frac{1}{2}}(\Omega_{0})\to H^{s-1}(\partial\Omega_{0})$, we deduce that
+$\partial_{z}\Psi(\varepsilon,\beta)|_{z=0}:(\varepsilon,\beta)\ni\mathcal{O}_{s}\to H^{s-1}(\mathbb{T})\quad\text{is~analytic}$ (2.58)
+for some neighborhood $\mathcal{O}_{s}$ of $(\varepsilon,\beta)=(0,\beta_{0})$. To treat the case $s=\frac{1}{2}$ we consider the vector field
+$u=\begin{bmatrix}\partial_{x}\Psi(\varepsilon,\beta)\\ \big(\frac{h}{h_{\varepsilon}}\big)^{2}\partial_{z}\Psi(\varepsilon,\beta)\end{bmatrix}$.
+Because $\operatorname{div}u=\beta\tilde{J}\Psi(\varepsilon,\beta)\in H^{1}(\Omega_{0})\subset L^{2}(\Omega_{0})$ in view of (2.53), we have that $u\in H_{\operatorname{div}}(\Omega_{0})$, defined by (2.37).
+Therefore, by virtue of Proposition 2.7, the trace $\gamma(u)=:\big(\frac{h}{h_{\varepsilon}}\big)^{2}\partial_{z}\Psi(\varepsilon,\beta)|_{z=0}$ is well-defined in $H^{-\frac{1}{2}}(\mathbb{T})$.
+Therefore, (2.58) is still valid for $s=\frac{1}{2}$.
+Viewing $\Psi(\varepsilon,\beta)$ as the bounded linear operator $f\mapsto\Psi$ solution of (2.53), we deduce with the aid of linear interpolation that the map
+$f\mapsto\partial_{z}\Psi(\varepsilon,\beta)|_{z=0}:(\varepsilon,\beta)\ni\mathcal{O}_{s}\to\mathcal{L}(H^{s}(\mathbb{T}),H^{s-\frac{1}{2}}(\mathbb{T}))$
+is analytic for all $s\geq\frac{1}{2}$. This concludes the proof of the analyticity (2.35) of $\mathcal{G}_{\varepsilon,\beta}$. $\square$
 2.4. The reformulated instability problem
@@
 For the spectral analysis, we seek solutions of the linearized system (2.61)–(2.62) of the form $w_{j}(x,t)=e^{\lambda t}u_{j}(x)$, thereby arriving at the eigenvalue problem
-$\lambda U=\mathcal{L}_{\varepsilon,\beta}U:=\begin{bmatrix}\partial_{x}(p(x)\cdot)&\mathcal{G}_{\varepsilon,\beta}\\ -\frac{1+q(x)}{\zeta^{\prime}(x)}&p(x)\partial_{x}\end{bmatrix}U,\quad U=\begin{bmatrix}u_{1}\\ u_{2}\end{bmatrix}.$ (2.85)
+$\lambda U=\mathcal{L}_{\varepsilon,\beta}U:=\begin{bmatrix}\partial_{x}(p(x)\cdot)&\mathcal{G}_{\varepsilon,\beta}\\ -\frac{1+q(x)}{\zeta^{\prime}(x)}&p(x)\partial_{x}\end{bmatrix}U,\quad U=\begin{bmatrix}u_{1}\\ u_{2}\end{bmatrix}$.
+(2.64)
 Of course $\mathcal{L}_{\varepsilon,\beta}$ depends also on the depth $h$ but we will drop this dependence to ease notation.
@@
 Proof.
-The proof of Proposition 2.10 is given in Appendix A.2 of [7], noting that their $f|_{\varepsilon}$, $\mathfrak{p}_{\varepsilon}$, $p_{\varepsilon}$, and $a_{\varepsilon}$ correspond to $(h_{\varepsilon}-h)$, $(\zeta-x)$, $(p-c_{0})$, and $[(1+q)/\zeta^{\prime}-1]$, respectively, in this work.
-∎
+[…]
 Combining Theorem 2.6 (with $s=1$) and Proposition 2.9, we conclude that the mapping
@@
 We learn from [34] that the resonance condition is a double eigenvalue $\lambda^{0}_{+}(-(m+1),\beta)=\lambda^{0}_{-}(m,\beta)$. That is,
-$-c_{0}(m+1)+((m+1)^{2}+\beta)^{\frac{1}{4}}\tanh^{\frac{1}{2}}\left(h\left((m+1)^{2}+\beta\right)^{\frac{1}{2}}\right)=c_{0}m-(m^{2}+\beta)^{\frac{1}{4}}\tanh^{\frac{1}{2}}\left(h\left(m^{2}+\beta\right)^{\frac{1}{2}}\right).$ (3.4)
-For our purposes, we choose the simplest resonance $m=1$. Then, we have the following proposition on the existence, uniqueness and asymptotics of the resonant value of $\beta$.
+$c_{0}(m+1)+((m+1)^{2}+\beta)^{\frac{1}{4}}\tanh^{\frac{1}{2}}\left(h\left((m+1)^{2}+\beta\right)^{\frac{1}{2}}\right)=c_{0}m-(m^{2}+\beta)^{\frac{1}{4}}\tanh^{\frac{1}{2}}\left(h\left(m^{2}+\beta\right)^{\frac{1}{2}}\right).$ (3.4)
+For our purposes, we choose the simplest resonance $m=1$, which leads to instabilities with the largest growth rates for small Stokes waves as suggested by numerical computations [34].
+Then, we have the following proposition on the existence, uniqueness and asymptotics of the resonant value of $\beta$.
 Proposition 3.1.
 For each $h>0$, there exists a unique $\beta_{*}=\beta_{*}(h)\in(0,\infty)$ such that (3.4) is satisfied and
-(i) $30<\beta_{*}(h)<$ 3,
+(i) $0<\beta_{*}(h)<3$,
 (ii) $\beta_{*}(h)$ is real analytic for $h>0$,
 (iii) $\beta_{*}(h)=\frac{4}{3}h^{2}+\mathscr{o}(h^{2})$ as $h\rightarrow 0^{+}$, and
-(iv) $\beta_{*}(h)=\beta_{*,\infty}-\frac{12e^{-2h}}{\left(1+\beta_{*,\infty}\right)^{-\frac{3}{4}}+\left(4+\beta_{*,\infty}\right)^{-\frac{3}{4}}}+\mathscr{o}\left(e^{-2h}\right)$ as $h\rightarrow\infty$, where $\beta_{*,\infty}=2.7275...$ is the root of (3.7).
+(iv) $\beta_{*}(h)=\beta_{*,\infty}-\frac{12e^{-2h}}{\left(1+\beta_{*,\infty}\right)^{-\frac{3}{4}}+\left(4+\beta_{*,\infty}\right)^{-\frac{3}{4}}}+\mathscr{o}\left(e^{-2h}\right)$ as $h\rightarrow\infty$, where $\beta_{*,\infty}=2.7275$...
+is the root of (3.7).
 Proof.
@@
 where the first inequality follows from the monotonicity of $\tanh^{\frac{1}{2}}(\xi)$ for $\xi>0$.
-Consequently, for each $h>0$, there exists $30<\beta_{*}<$ 3 such that $F(\beta_{*},h)=0$.
-Because $\partial_{\beta}F<0$, we deduce that $\beta_{*}(h)$ is the unique positive zero of $F(\cdot,h)$ This proves (i).
+Consequently, for each $h>0$, there exists $0<\beta_{*}<3$ such that $F(\beta_{*},h)=0$.
+Because $\partial_{\beta}F<0$, we deduce that $\beta_{*}(h)$ is the unique positive zero of $F(\cdot,h)$ .
+This proves (i).
 By the Analytic Implicit Function Theorem, $\beta_{*}(h)$ is real analytic for $h>0$, proving (ii).
@@
 $3-(1+\beta_{*}(h))^{\frac{1}{2}}+O(h^{2})$ $=(4+\beta_{*}(h))^{\frac{1}{2}},$
-and upon squaring both sides, while using the bound $30<\beta_{*}(h)<$ 3, we find
+and upon squaring both sides, while using the bound $0<\beta_{*}(h)<3$, we find
 $\left(1+\beta_{*}(h)\right)^{\frac{1}{2}}=1+O(h^{2})\quad\textrm{as}\quad h\rightarrow 0^{+}.$
@@
 $\quad-\frac{(4+\beta)^{\frac{3}{4}}}{3}\left(\frac{\textrm{sech}\left(h\left(4+\beta\right)^{\frac{1}{2}}\right)}{\textrm{sech}(h)}\right)^{2}\left(\frac{\tanh(h)}{\tanh\left(h\left(4+\beta\right)^{\frac{1}{2}}\right)}\right)^{\frac{1}{2}}.$
-From the monotonicity of $\tanh$, we have $\tanh(h)/\tanh(wh)<1$ for $h>0,\ w>1$. Thus
+From the monotonicity of $\tanh$, we have $\tanh(h)/\tanh(wh)<1$ for h>0,w>1h>0,\ w>1. Thus
 $g(\beta,h)>1-\frac{(1+\beta)^{\frac{3}{4}}}{3}\left(\frac{\textrm{sech}\left(h\left(1+\beta\right)^{\frac{1}{2}}\right)}{\textrm{sech}(h)}\right)^{2}-\frac{(4+\beta)^{\frac{3}{4}}}{3}\left(\frac{\textrm{sech}\left(h\left(4+\beta\right)^{\frac{1}{2}}\right)}{\textrm{sech}(h)}\right)^{2}.$
-For $30<\beta<$ 3, we have the bounds
+For $0<\beta<3$, we have the bounds
 $\frac{(1+\beta)^{\frac{3}{4}}}{3}\left(\frac{\textrm{sech}\left(h\left(1+\beta\right)^{\frac{1}{2}}\right)}{\textrm{sech}(h)}\right)^{2}<\frac{4^{\frac{3}{4}}}{3}\quad\textrm{and}\quad\frac{(4+\beta)^{\frac{3}{4}}}{3}\left(\frac{\textrm{sech}\left(h\left(4+\beta\right)^{\frac{1}{2}}\right)}{\textrm{sech}(h)}\right)^{2}<\frac{7^{\frac{3}{4}}}{3}\left(\frac{\textrm{sech}(2h)}{\textrm{sech}(h)}\right)^{2},$
@@
 $g(\beta,h)>1-\frac{4^{\frac{3}{4}}}{3}-\frac{7^{\frac{3}{4}}}{3}\left(\frac{\textrm{sech}(2h)}{\textrm{sech}(h)}\right)^{2}.$
-Now $01-4^{\frac{3}{4}}/3>$ 0 and $\textrm{sech}(2h)/\textrm{sech}(h)\rightarrow 0$ as $h\rightarrow\infty$, so that there exists $H_{0}>0$ such that
+Now $1-4^{\frac{3}{4}}/3>0$ and $\textrm{sech}(2h)/\textrm{sech}(h)\rightarrow 0$ as $h\rightarrow\infty$, so that there exists $H_{0}>0$ such that
 $1-\frac{4^{\frac{3}{4}}}{3}-\frac{7^{\frac{3}{4}}}{3}\left(\frac{\textrm{sech}(2h)}{\textrm{sech}(h)}\right)^{2}>0\quad\forall h>H_{0}.$
-Consequently, $\partial_{h}F(\beta,h)>0$ for all $30<\beta<$ 3 and $h>H_{0}$.
+Consequently, $\partial_{h}F(\beta,h)>0$ for all $0<\beta<3$ and $h>H_{0}$.
 In particular, $(\partial_{h}F)(\beta_{*}(h),h)>0$ for $h>H_{0}$.
@@
 Taking the limit of $F(\beta_{*}(h),h)=0$ as $h\rightarrow\infty$, we arrive at the following equation for $\beta_{*,\infty}$:
-$0\displaystyle 3-(1+\beta_{*,\infty})^{\frac{1}{4}}-(4+\beta_{*,\infty})^{\frac{1}{4}}=$ 0 . (3.7)
+$3-(1+\beta_{*,\infty})^{\frac{1}{4}}-(4+\beta_{*,\infty})^{\frac{1}{4}}=0.$ (3.7)
 The solution of (3.7) is
@@
 We denote the spectral projection associated to $\Gamma$ by
-[…]
+$P_{\varepsilon,\delta}=-\frac{1}{2\pi i}\int_{\Gamma}(\mathcal{L}_{\varepsilon,\beta_{*}+\delta}-\lambda)^{-1}d\lambda\in\mathcal{L}\big((L^{2}(\mathbb{T}))^{2},(H^{1}(\mathbb{T}))^{2}\big).$ (3.17)
 The analyticity (2.91) of $\mathcal{L}_{\varepsilon,\beta}$ implies that $P_{\varepsilon,\delta}$ is analytic in $(\varepsilon,\delta)$ near $(0,0)$.
@@
 Using our notation $\beta=\beta_{*}+\delta$, the $2\times 2$ matrix that represents the linear operator $\mathcal{L}_{\varepsilon,\beta}=J\mathcal{H}_{\varepsilon,\beta}:\mathcal{V}_{\varepsilon,\delta}\to\mathcal{V}_{\varepsilon,\delta}$ with respect to the basis $\{V_{1}^{\varepsilon,\delta},V_{2}^{\varepsilon,\delta}\}$ is
-$\textrm{L}_{\varepsilon,\delta}=\begin{bmatrix}-\frac{i}{4\pi}(\mathcal{H}_{\varepsilon,\beta}V_{1}^{\varepsilon,\delta},V_{1}^{\varepsilon,\delta})&\frac{i}{4\pi}(\mathcal{H}_{\varepsilon,\beta}V_{1}^{\varepsilon,\delta},V_{2}^{\varepsilon,\delta})\\ -\frac{i}{4\pi}(\mathcal{H}_{\varepsilon,\beta}V_{2}^{\varepsilon,\delta},V_{1}^{\varepsilon,\delta})&\frac{i}{4\pi}(\mathcal{H}_{\varepsilon,\beta}U_{2}^{\varepsilon,\delta},V_{2}^{\varepsilon,\delta})\end{bmatrix}.$ (3.23)
+$\textrm{L}_{\varepsilon,\delta}=\begin{bmatrix}-\frac{i}{4\pi}(\mathcal{H}_{\varepsilon,\beta}V_{1}^{\varepsilon,\delta},V_{1}^{\varepsilon,\delta})&\frac{i}{4\pi}(\mathcal{H}_{\varepsilon,\beta}V_{1}^{\varepsilon,\delta},V_{2}^{\varepsilon,\delta})\\ -\frac{i}{4\pi}(\mathcal{H}_{\varepsilon,\beta}V_{2}^{\varepsilon,\delta},V_{1}^{\varepsilon,\delta})&\frac{i}{4\pi}(\mathcal{H}_{\varepsilon,\beta}U_{2}^{\varepsilon,\delta},V_{2}^{\varepsilon,\delta})\end{bmatrix}$.
+(3.23)
 $\textrm{L}_{\varepsilon,\delta}$ is purely imaginary and
@@
 The coefficients $r_{1,1},\dots,p_{3,3}$ are given in Prop.
-2.10.
+2.9.
 The nonlocal operators $R_{j,\beta}$ are Fourier multipliers.
@@
 $\zeta=x+\varepsilon\zeta^{1}+\varepsilon^{2}\zeta^{2}+\varepsilon^{3}\zeta^{3}+O(\varepsilon^{4}),\quad\eta^{*}=\varepsilon\eta^{1}+\varepsilon^{2}\eta^{2}+\varepsilon^{3}\eta^{3}+O(\varepsilon^{4}),$
-where the coefficients are given in Prop. 2.10 and Prop. 2.2. A Taylor expansion leads to
+where the coefficients are given in Prop. 2.9 and Prop. 2.2. A Taylor expansion leads to
 $\eta^{*}(\zeta(x))$ $=\varepsilon\eta^{1}(x)+\varepsilon^{2}(\zeta^{1}\partial_{x}\eta^{1}(x)+\eta^{2}(x))$ (4.8)
@@
 $\begin{cases}\Delta_{x,z}\Theta^{1}-\beta\Theta^{1}=2\beta\frac{C_{h+z}}{C_{h}}\zeta_{1,1}(\cos x)\Theta^{0}\quad\text{in }\{(x,z):-h<z<0\},\\ \Theta^{1}(x,0)=0,\\ \partial_{z}\Theta^{1}(x,-h)=0,\end{cases}$ (4.12)
-$&\Delta_{x,z}\Theta^{2}-\beta\Theta^{2}=\beta\biggr{[}\frac{C_{2(h+z)}}{2C_{h}^{2}}\zeta_{1,1}^{2}+\Big{(}4\frac{C_{2(h+z)}}{C_{2h}}\zeta_{2,2}+\frac{1}{2C_{h}^{2}}\zeta_{1,1}^{2}\Big{)}\cos(2x)\biggr{]}\Theta^{0}\\ &\hskip 88.2037pt+2\beta\frac{C_{h+z}}{C_{h}}\zeta_{1,1}\cos x\Theta^{1}\quad\text{in }\{(x,z):-h<z<0\},\\ &\Theta^{2}(x,0)=0,\\ &\partial_{z}\Theta^{2}(x,-h)=h_{2}\partial_{z}^{2}\Theta^{0}(x,-h),\end{aligned}\end{cases}$ (4.13)
-$\Delta_{x,z}\Theta^{3}-\beta\Theta^{3}&=2\beta\biggr{[}\Big{(}\frac{d_{2}S_{z}}{C_{h}^{2}}\zeta_{1,1}+\frac{C_{3(h+z)}}{C_{h}C_{2h}}\zeta_{1,1}\zeta_{2,2}+\frac{C_{h+z}}{C_{h}}\zeta_{3,1}\Big{)}\cos x+\Big{(}\frac{C_{h+z}}{C_{h}C_{2h}}\zeta_{1,1}\zeta_{2,2}\\ &\hskip 24.18501pt+3\frac{C_{3(h+z)}}{C_{3h}}\zeta_{3,3}\Big{)}\cos(3x)\biggr{]}\Theta^{0}+\beta\biggr{[}\frac{C_{2(h+z)}}{2C_{h}^{2}}\zeta_{1,1}^{2}+\Big{(}4\frac{C_{2(h+z)}}{C_{2h}}\zeta_{2,2}\\ &\hskip 24.18501pt+\frac{1}{2C_{h}^{2}}\zeta_{1,1}^{2}\Big{)}\cos(2x)\biggr{]}\Theta^{1}+2\beta\frac{C_{h+z}}{C_{h}}\zeta_{1,1}\cos x\Theta^{2}\quad\text{in}\{(x,z):-h<z<0\},\end{aligned}\\ \Theta^{3}(x,0)=0,\\ \partial_{z}\Theta^{3}(x,-h)=h_{2}\partial_{z}^{2}\Theta^{1}(x,-h).\end{cases}$ (4.14)
+$\begin{cases}\begin{aligned} &\Delta_{x,z}\Theta^{2}-\beta\Theta^{2}=\beta\biggr[\frac{C_{2(h+z)}}{2C_{h}^{2}}\zeta_{1,1}^{2}+\Big(4\frac{C_{2(h+z)}}{C_{2h}}\zeta_{2,2}+\frac{1}{2C_{h}^{2}}\zeta_{1,1}^{2}\Big)\cos(2x)\biggr]\Theta^{0}\\ &\hskip 88.2037pt+2\beta\frac{C_{h+z}}{C_{h}}\zeta_{1,1}\cos x\Theta^{1}\quad\text{in }\{(x,z):-h<z<0\},\\ &\Theta^{2}(x,0)=0,\\ &\partial_{z}\Theta^{2}(x,-h)=h_{2}\partial_{z}^{2}\Theta^{0}(x,-h),\end{aligned}\end{cases}$ (4.13)
+$\begin{cases}\begin{aligned} \Delta_{x,z}\Theta^{3}-\beta\Theta^{3}&=2\beta\biggr[\Big(\frac{d_{2}S_{z}}{C_{h}^{2}}\zeta_{1,1}+\frac{C_{3(h+z)}}{C_{h}C_{2h}}\zeta_{1,1}\zeta_{2,2}+\frac{C_{h+z}}{C_{h}}\zeta_{3,1}\Big)\cos x+\Big(\frac{C_{h+z}}{C_{h}C_{2h}}\zeta_{1,1}\zeta_{2,2}\\ &\hskip 24.18501pt+3\frac{C_{3(h+z)}}{C_{3h}}\zeta_{3,3}\Big)\cos(3x)\biggr]\Theta^{0}+\beta\biggr[\frac{C_{2(h+z)}}{2C_{h}^{2}}\zeta_{1,1}^{2}+\Big(4\frac{C_{2(h+z)}}{C_{2h}}\zeta_{2,2}\\ &\hskip 24.18501pt+\frac{1}{2C_{h}^{2}}\zeta_{1,1}^{2}\Big)\cos(2x)\biggr]\Theta^{1}+2\beta\frac{C_{h+z}}{C_{h}}\zeta_{1,1}\cos x\Theta^{2}\quad\text{in }\{(x,z):-h<z<0\},\end{aligned}\\ \Theta^{3}(x,0)=0,\\ \partial_{z}\Theta^{3}(x,-h)=h_{2}\partial_{z}^{2}\Theta^{1}(x,-h).\end{cases}$ (4.14)
 Comparing with the expansion (4.1), we find that $R_{j,\beta}f=\partial_{z}\Theta^{j}(\cdot,0)$.
@@
 We recall the basis $\{U_{1},U_{2}\}$ from (3.13) and the projections $P_{\varepsilon,\delta}$ from (3.17). Since $P_{0,0}U_{j}=U_{j}$ for $j=1,2$, we have
-$U^{\varepsilon,\delta}_{j}$ =$=\{1-(P_{\varepsilon,\delta}-P_{0,0})^{2}\}^{-\frac{1}{2}}P_{\varepsilon,\delta}U_{j}.$
+$U^{\varepsilon,\delta}_{j}$ $=\{1-(P_{\varepsilon,\delta}-P_{0,0})^{2}\}^{-\frac{1}{2}}P_{\varepsilon,\delta}U_{j}.$
 Next, we record the following third-order expansions of $U^{\varepsilon,\delta}_{j}$ and $\textrm{L}_{\varepsilon,\delta}$. We refer to [18] for the derivation of these formulas. Denoting
@@
 $p(\xi)>0.$
-A similar argument shows that $q(\xi)>0$, and in [17] we show that $r(\xi)>0$ and $21<\xi<$ 2. It follows that
+A similar argument shows that $q(\xi)>0$, and in [18] we show that $r(\xi)>0$ and $1<\xi<2$. It follows that
 $b_{3,0,\infty}<0.$
@@
 $\mathcal{Z}$ is bounded because $\lim_{h\to\infty}b_{3,0}(h)<0$.
-Moreover, 00 cannot be an accumulation point of $\mathcal{Z}$ because $\lim_{h\to 0^{+}}b_{3,0}(h)=\infty$.
+Moreover, 0 cannot be an accumulation point of $\mathcal{Z}$ because $\lim_{h\to 0^{+}}b_{3,0}(h)=\infty$.
 Thus $\mathcal{Z}$ is finite.
-[…]
+Although by analytic reasoning we have not been able to determine the number of zeros of $b_{3,0}$, a numerical calculation indicates that there is exactly one.
+Indeed, the numerical plot of $b_{3,0}$ in Figure 3 reveals that exactly one such depth occurs at $h_{crit}=0.25065$....
+In order to obtain Figure 3, we use Newton’s method to compute $\beta_{*}(h)$ for a given $h$ from the resonance condition (3.4).
+This in turn allows us to evaluate $b_{3,0}$ numerically at depth $h$.
+Repeating this procedure for $0.1<h<4$ in step sizes of $\Delta h=0.01$ yields Figure 3.
+The critical depth $h_{crit}=0.25065$...
+is obtained by bisection using increasing smaller $\Delta h$ near the critical depth.
 Figure 3.
@@
 $\lambda_{\pm}=i\Big(\sigma+\frac{1}{2}(A+C)\Big)\pm\frac{1}{2}\sqrt{\Delta(\varepsilon,\delta)},$ (5.7)
-where $A+C$ is the trace of $\text{L}_{\varepsilon,\delta}-i\sigma I$ and so is real-valued.
+where we recall that $\sigma+\frac{1}{2}(A+C)$ is real.
 Since $A$, $B$, and $C$ are real analytic in $(\varepsilon,\delta)$ in a neighborhood of $(0,0)$, so are $A+C$ and $\Delta(\varepsilon,\delta)$.
@@
 Taking inspiration from our work in infinite depth [18], we map $\delta$ to a new variable $\theta$ as follows:
-$\delta=\kappa_{0}\varepsilon^{2}+\theta\varepsilon^{3},\quad\textrm{where}\quad\kappa_{0}:=\frac{a_{2,0}-c_{2,0}}{a_{0,1}-c_{0,1}}.$ (5.8)
+$\delta=\kappa_{0}\varepsilon^{2}+\theta\varepsilon^{3},\qquad\textrm{where}\qquad\kappa_{0}:=\frac{a_{2,0}-c_{2,0}}{a_{0,1}-c_{0,1}}.$ (5.8)
 This transformation is well-defined by Lemma 5.2. Upon substituting (5.8) into the series expansions of $A$, $B$, and $C$ (5.2), we obtain the following expansion for $\Delta$:
@@
 Choosing $\theta$ such that
-$|\theta|<\kappa_{1},\quad\textrm{where}\quad\kappa_{1}:=\frac{2|b_{3,0}|}{|a_{0,1}-c_{0,1}|},$ (5.10)
+$|\theta|<\kappa_{1},\qquad\textrm{where}\qquad\kappa_{1}:=\frac{2|b_{3,0}|}{|a_{0,1}-c_{0,1}|},$ (5.10)
 we appear to achieve $\Delta(\varepsilon,\delta)>0$ for all sufficiently small $\varepsilon>0$, as desired.
@@
 $\lambda_{\pm}\equiv\lambda_{\pm}(\varepsilon,\theta)$ $=i\left[\sigma+\frac{a_{0,1}c_{2,0}-a_{2,0}c_{0,1}}{(a_{0,1}-c_{0,1})}\varepsilon^{2}+\frac{a_{0,1}+c_{0,1}}{2}\theta\varepsilon^{3}\right]$ (5.11)
-±$\qquad\pm\frac{1}{2}\left[4b_{3,0}^{2}-(a_{0,1}-c_{0,1})^{2}\theta^{2}\right]^{\frac{1}{2}}|\varepsilon|^{3}+O(\varepsilon^{4}).$
+$\qquad\pm\frac{1}{2}\left[4b_{3,0}^{2}-(a_{0,1}-c_{0,1})^{2}\theta^{2}\right]^{\frac{1}{2}}|\varepsilon|^{3}+O(\varepsilon^{4}).$
 By eliminating $\theta$ in from the real and imaginary parts of $\lambda_{\pm}$, we obtain
@@
 in the complex plane. See Figure 1.
-Appendix A Normal trace of vector fields
-We recall the following classical trace operator.
-Proposition A.1 (Section 3.2, Chapter IV, [9]).
-Let $U\subset\mathbb{R}^{n}$ be Lipschitz domain with compact boundary and denote
-$H_{\operatorname{div}}(U)=\{u\in L^{2}(U;\mathbb{R}^{n}):\operatorname{div}u\in L^{2}(U)\}.$ (A.1)
-There exists a bounded linear operator $\gamma:H_{\operatorname{div}}(U)\to H^{\frac{1}{2}}(\partial U)$ such that $\gamma(u)=u\cdot\nu$ for $u\in C^{\infty}_{0}(\overline{U})$, where $\nu$ is the unit outward normal to $\partial U$.
-In particular, there exists a constant $C>0$ such that
-$\|\gamma(u)\|_{H^{-\frac{1}{2}}(\partial U)}\leq C\Big{(}\|u\|_{L^{2}(U)}+\|\operatorname{div}u\|_{L^{2}(U)}\Big{)}\quad\forall u\in H_{\operatorname{div}}(U).$ (A.2)
-Moreover, if $u\in H_{\operatorname{div}}(U)$ and $w\in H^{1}(U)$, then the Stokes formula
-$\langle\gamma(u),\gamma_{0}(w)\rangle_{H^{-\frac{1}{2}}(\partial U),H^{\frac{1}{2}}(\partial U)}=\int_{U}u\cdot\nabla w+\int_{U}w\operatorname{div}u$ (A.3)
-holds, where $\gamma_{0}(w)$ is the trace of $w$.
-The following lemma provides an extension of the preceding trace operator for a cylindrical shell that is sufficient for our purposes.
-Lemma A.2.
-Let $U=\mathbb{T}\times(a,b)$, where $-\infty<a<b<\infty$.
-Since $\partial U=(\mathbb{T}\times\{a\})\cup(\mathbb{T}\times\{b\})$, we identify any function $g$ on $\partial U$ with the pair $(g_{1},g_{2})\equiv(g(\cdot,a),g(\cdot,b)):\mathbb{T}\to\mathbb{R}^{2}$.
-We denote a point in $U$ by $(x,z)$ and set
-$\tilde{H}_{\operatorname{div}}(U)=\{u\in L^{2}(U;\mathbb{R}^{2}):\operatorname{div}u=\partial_{x}^{2}F\quad\text{for~{}some~{}}F\in H^{1}(U)\}.$ (A.4)
-(i) There exists a linear operator $\tilde{\gamma}:\tilde{H}_{\operatorname{div}}(U)\to(H^{-\frac{1}{2}}(\mathbb{T}))^{2}$ with the following properties
-  • (a)
-    $\tilde{\gamma}(u)=\gamma(u)$ if $u\in\tilde{H}_{\operatorname{div}}(U)\cap H_{\operatorname{div}}(U)$.
-  • (b)
-    For some constant $C=C(a,b)>0$, we have
-    $\|\tilde{\gamma}(u)\|_{(H^{-\frac{1}{2}}(\mathbb{T}))^{2}}\leq C(\|u\|_{L^{2}(U)}+\|\partial_{x}F\|_{L^{2}(U)}),$ (A.5)
-    where the right-hand side of (A.5) is independent of the choice of $F$.
-  • (c)
-    If $u=\begin{bmatrix}u_{1}\\ 0\end{bmatrix}\in\tilde{H}_{\operatorname{div}}(U)$, then $\tilde{\gamma}(u)=0$.
-(ii) We extend $\tilde{\gamma}$ to $\tilde{H}_{\operatorname{div}}(U)+H_{\operatorname{div}}(U)$ by defining
-$\tilde{\gamma}(u)=\tilde{\gamma}(\tilde{v})+\gamma(v)$ (A.6)
-if $u=\tilde{v}+v$, $\tilde{v}\in\tilde{H}_{\operatorname{div}}(U)$, $v\in H_{\operatorname{div}}(U)$.
-This definition is independent of the choice of the pair $(\tilde{v},v)$.
-As a consequence, if $u\in L^{2}(U;\mathbb{R}^{2})$ and $\operatorname{div}u=G+\partial_{x}^{2}F$ for some $G\in L^{2}(U)$ and $F\in H^{1}(U)$, then
-$\|\tilde{\gamma}(u)\|_{(H^{-\frac{1}{2}}(\mathbb{T}))^{2}}\leq C(\|u\|_{L^{2}(U)}+\|G\|_{L^{2}(U)}+\|\partial_{x}F\|_{L^{2}(U)}).$ (A.7)
-Proof.
-(i) For any $u\in\tilde{H}_{\operatorname{div}}(U)$ with $\operatorname{div}u=\partial_{x}^{2}F$, using a lifting operator $\ell_{U}:(H^{\frac{1}{2}}(\mathbb{T}))^{2}\to H^{1}(U)$, we define $\tilde{\gamma}(u)\in(H^{-\frac{1}{2}}(\mathbb{T}))^{2}$ by
-$\langle\tilde{\gamma}(u),g\rangle_{(H^{-\frac{1}{2}}(\mathbb{T}))^{2},(H^{\frac{1}{2}}(\mathbb{T}))^{2}}:=(u,\nabla\ell_{U}g)_{L^{2}(U)}-(\partial_{x}F,\partial_{x}\ell_{U}g)_{L^{2}(U)}\quad\forall g\in(H^{\frac{1}{2}}(\mathbb{T}))^{2}.$ (A.8)
-Clearly (A.8) defines $\tilde{\gamma}(u)$ as a bounded linear functional on $(H^{\frac{1}{2}}(\mathbb{T}))^{2}$ with norm bounded by $C(\|u\|_{L^{2}(U)}+\|\partial_{x}F\|_{L^{2}(U)})$.
-Thus (A.5) is valid.
-This norm bound is independent of the choice of $F$.
-Indeed, if $\operatorname{div}u=\partial_{x}^{2}F=\partial_{x}^{2}F_{1}$, then $F(x,z)-F_{1}(x,z)$ is independent of $x$ due to the periodicity in $x$, whence $\partial_{x}F=\partial_{x}F_{1}$.
-Now suppose that $u\in\tilde{H}_{\operatorname{div}}(U)\cap H_{\operatorname{div}}(U)$, so that $L^{2}(U)\ni\operatorname{div}u=\partial_{x}^{2}F$ for some $F\in H^{1}(U)$.
-The Stokes formula (A.3) implies
-$\langle\gamma(u),g\rangle_{H^{-\frac{1}{2}}(\partial U),H^{\frac{1}{2}}(\partial U)}$ $=(u,\nabla\ell_{U}g)_{L^{2}(U)}+(\partial^{2}_{x}F,\ell_{U}g)_{L^{2}(U)}.$
-Since $\partial_{x}^{2}F=\operatorname{div}u\in L^{2}(U)$, $F\in H^{1}(U)$, and $\ell_{U}g\in H^{1}(U)$, we have $(\partial^{2}_{x}F,\ell_{U}g)_{L^{2}(U)}=-(\partial_{x}F,\partial_{x}\ell_{U}g)_{L^{2}(U)}$ by the periodicity in $x$, and hence for all $g\in H^{\frac{1}{2}}(\mathbb{T}))^{2}$,
-$\langle\gamma(u),g\rangle_{H^{-\frac{1}{2}}(\partial U),H^{\frac{1}{2}}(\partial U)}=(u,\nabla\ell_{U}g)_{L^{2}(U)}-(\partial_{x}F,\partial_{x}\ell_{U}g)_{L^{2}(U)}=\langle\tilde{\gamma}(u),g\rangle_{(H^{-\frac{1}{2}}(\mathbb{T}))^{2},(H^{\frac{1}{2}}(\mathbb{T}))^{2}}.$
-Therefore, $\tilde{\gamma}=\gamma$ on $\tilde{H}_{\operatorname{div}}(U)\cap H_{\operatorname{div}}(U)$.
-Assume that $u=\begin{bmatrix}u_{1}\\ 0\end{bmatrix}\in\tilde{H}_{\operatorname{div}}(U)$.
-Then $\partial_{x}^{2}F=\operatorname{div}u=\partial_{x}u_{1}$ for some $F\in H^{1}(U)$, and hence $u_{1}(x,z)=\partial_{x}F(x,z)+f(z)$ for some $f\in L^{2}((a,b))$.
-The definition (A.8) implies that for all $g\in(H^{\frac{1}{2}}(\mathbb{T}))^{2}$ we have
-$\langle\tilde{\gamma}(u),g\rangle_{(H^{-\frac{1}{2}}(\mathbb{T}))^{2},(H^{\frac{1}{2}}(\mathbb{T}))^{2}}$ $=(u_{1},\partial_{x}\ell_{U}g)_{L^{2}(U)}-(\partial_{x}F,\partial_{x}\ell_{U}g)_{L^{2}(U)}$
-$=(f,\partial_{x}\ell_{U}g)_{L^{2}(U)}=\int_{a}^{b}f(z)\int_{\mathbb{T}}\partial_{x}\ell_{U}g(x,z)dxdz=0$
-due to the periodicity in $x$. Thus $\tilde{\gamma}(u)=0$.
-(ii) Next, we show that the definition (A.6) of the extended $\tilde{\gamma}$ is well defined.
-Suppose that $u=\tilde{v}+v=\tilde{w}+w$, where $\tilde{v}$, $\tilde{w}\in\tilde{H}_{\operatorname{div}}(U)$, and $v$, $w\in H_{\operatorname{div}}(U)$.
-Then $\tilde{v}-\tilde{w}=w-v\in\tilde{H}_{\operatorname{div}}(U)\cap H_{\operatorname{div}}(U)$, and thus $\tilde{\gamma}(\tilde{v}-\tilde{w})=\gamma(w-v)$.
-It follows that $\tilde{\gamma}(\tilde{v})+\gamma(v)=\tilde{\gamma}(\tilde{w})+\gamma(w)$.
-Suppose that $u\in L^{2}(U;\mathbb{R}^{2})$ and $\operatorname{div}u=G+\partial_{x}^{2}F$ for some $G\in L^{2}(U)$ and $F\in H^{1}(U)$.
-Let $\phi:U\to\mathbb{R}$ be the $H^{2}(U)\cap H^{1}_{0}(U)$ solution to the Poisson equation $\Delta\phi=G$, and set $v=\nabla\phi$, $\tilde{v}=u-v$.
-Then $v\in H_{\operatorname{div}}(U)$, $\tilde{v}\in\tilde{H}_{\operatorname{div}}(U)$ with $\operatorname{div}\tilde{v}=\partial_{x}^{2}F$, and $\|\phi\|_{H^{2}(U)}\leq C\|G\|_{L^{2}(U)}$.
-Therefore it follows from (A.6), (A.5) and (A.2) that
-$\|\tilde{\gamma}(u)\|_{(H^{-\frac{1}{2}}(\mathbb{T}))^{2}}$ $\leq\|\tilde{\gamma}(\tilde{v})\|_{(H^{-\frac{1}{2}}(\mathbb{T}))^{2}}+\|\gamma(v)\|_{(H^{-\frac{1}{2}}(\mathbb{T}))^{2}}$
-$\leq C\left(\|\tilde{v}\|_{L^{2}(U)}+\|\partial_{x}F\|_{L^{2}(U)}+\|v\|_{L^{2}(U)}+\|\operatorname{div}v\|_{L^{2}(U)}\right)$
-$\leq C\left(\|u\|_{L^{2}(U)}+\|v\|_{L^{2}(U)}+\|\partial_{x}F\|_{L^{2}(U)}+\|v\|_{L^{2}(U)}+\|\Delta\phi\|_{L^{2}(U)}\right),$
-which implies (A.7). ∎
 Acknowledgment. The work of HQN was partially supported by NSF grant DMS-2205710. The work of RPC was partially supported by NSF grant DMS-2402044.